### v1

# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

Abstract
We study online multiple testing with feedback, where decisions are made sequentially and the true state of the hypothesis is revealed after the decision has been made, either instantly or with a delay. We propose GAIF, a feedback-enhanced generalized alpha-investing framework that dynamically adjusts thresholds using revealed outcomes, ensuring finite-sample false discovery rate (FDR)/marginal FDR control. Extending GAIF to online conformal testing, we construct independent conformal $p$-values and introduce a feedback-driven model selection criterion to identify the best model/score, thereby improving statistical power. We demonstrate the effectiveness of our methods through numerical simulations and real-data applications.
**footnotetext: The first two authors equally contributed to this work.footnotetext: Correspondence to: Haojie Ren <haojieren@sjtu.edu.cn>
Keywords: Conformal prediction; Distributional shifts; Generalized alpha-investing procedure; Model selection; Online conformal $p$-value; Online FDR control
1 Introduction
Real-time decision making plays a critical role in a growing number of modern applications, such as online recruitment for job hiring [Faliagka et al., 2014], real-time alignment of large language models [Huang et al., 2025], and time-series anomaly detection [Rebjock et al., 2021], etc. These tasks can be naturally formulated as online multiple testing problems [Foster and Stine, 2008]. Consider a potentially infinite stream of null hypotheses $\{\mathbb{H}_{01},\mathbb{H}_{02},\dots,\mathbb{H}_{0t},\dots\}$ tested sequentially based on arriving test statistics $\{z_{1},z_{2},\dots\}$. At each time $t$, a real-time decision must be made upon observing $z_{t}$. Let $\theta_{t}$ denote the true state of the hypothesis at time $t$, where $\theta_{t}=0$ if $\mathbb{H}_{0t}$ is true and $\theta_{t}=1$ otherwise. The testing problem at time $t$ is
$\mathbb{H}_{0t}:\theta_{t}=0\quad\text{vs.}\quad\mathbb{H}_{1t}:\theta_{t}=1.$
Let $\delta_{t}\in\{0,1\}$ be the decision at time $t$, with $\delta_{t}=1$ indicating rejection of $\mathbb{H}_{0t}$. To ensure the reliability of the testing procedure, it is essential to control a well-defined error rate. Define the false discovery proportion (FDP) and false discovery rate (FDR) [Benjamini and Hochberg, 1995] at time $t$ by
$\operatorname{FDP}(t)=\frac{V(t)}{1\vee R(t)}\;:=\;\frac{\sum_{j=1}^{t}\delta_{j}\bigl{(}1-\theta_{j}\bigr{)}}{1\vee\sum_{j=1}^{t}\delta_{j}},\quad\operatorname{FDR}(t)=\mathbb{E}\left\{\operatorname{FDP}(t)\right\},$
where $V(t)$ and $R(t)$ represent the numbers of false rejections (discoveries) and rejections at time $t$, respectively. Given a target level $\alpha\in(0,1)$, classical work on online multiple testing [Ramdas et al., 2017, 2018, Tian and Ramdas, 2019] aims to guarantee $\sup_{t}\operatorname{FDR}(t)\leq\alpha$ or its weaker variant [Foster and Stine, 2008, Zrnic et al., 2021], the marginal FDR (mFDR), i.e., $\operatorname{mFDR}(t)={\mathbb{E}\{V(t)\}}/{\mathbb{E}\{1\vee R(t)\}}$.
Unlike the classical setting, we consider another realistic scenario in which the decision-maker receives real-time feedback on $\theta_{t}$ after issuing the decision $\delta_{t}$. Concretely, at each time $t$, the decision $\delta_{t}$ is based on the current statistic $z_{t}$, the history $\{\delta_{1},\dots,\delta_{t-1}\}$, and any feedback observed up to that point. The feedback $\theta_{t}$ may arrive immediately or with some delay (for example, as $\theta_{t-d}$ for $d\geq 1$) before moving on to time $t+1$. This feedback-available framework naturally arises in many practical applications. To demonstrate its relevance, we present three motivating examples: online conformal selection, real-time alignment of large language models, and time-series anomaly detection.
  • Online conformal selection. Conformal selection aims to identify valuable individuals whose unknown label $Y$ satisfies a pre-specified requirement by leveraging machine learning predictions [Jin and Candès, 2023a, Wu et al., 2024]. The online conformal selection setting naturally aligns with the feedback-available online multiple testing framework. For example, in diabetes risk prediction, a patient’s true condition may later be confirmed by an expert, offering feedback that could improve future decisions. Some works [Huo et al., 2024, Xu and Ramdas, 2024] considered online conformal testing problems but did not exploit feedback information.
  • Real-time LLM alignment. Large language models (LLMs) are increasingly used in high-stakes domains such as healthcare, finance, and law, where outputs must be reliable and factual. However, LLMs can hallucinate—producing plausible but incorrect content [Huang et al., 2025]. A remedy is to filter or certify LLM outputs [Gui et al., 2024, Bai and Jin, 2024]. Some alignment approaches based on conformal testing are proposed for this use but are generally designed for offline environments, whereas many applications demand immediate and trustworthy screening. In such applications, follow-up feedback is usually available, which, if incorporated, could continuously improve alignment.
  • Time series anomaly detection. Detecting anomalies in time series is crucial for industrial monitoring, fraud detection, and healthcare analytics. To ensure reliability, prior work [Gang et al., 2021, Rebjock et al., 2021, Krönert et al., 2023] addresses online FDR control but typically ignores real-time feedback. In practice, such feedback is often available: once an anomaly is flagged, a subsequent system failure, user verification of a fraudulent transaction, or expert annotation may confirm or refute the alarm.
These observations underscore that real-time feedback is a foundational element of adaptive decision-making, playing a pivotal role across a wide range of online applications. Despite its importance, current online multiple testing procedures seldom incorporate feedback in a systematic way. This naturally raises a fundamental and compelling question: Can real-time feedback be effectively incorporated into online multiple testing procedures in a way that allows us to enhance statistical power while still ensuring valid error rate control? To this end, we develop a new framework that systematically integrates feedback into the online testing process, achieving significant performance gains without compromising statistical validity. Specifically, our proposed strategies utilize feedback through three key mechanisms: (1) within the proposed GAIF framework—a feedback-enhanced extension of generalized alpha-investing (GAI) [Ramdas et al., 2017, Javanmard and Montanari, 2018]—we adaptively adjust testing thresholds based on past decisions and revealed outcomes together; (2) by updating an online calibration dataset—used to generate explicit, valid, and independent online $p$-values for conformal testing; and (3) by guiding model selection in online conformal testing. These enhancements collectively improve power over existing GAI methods to a large degree, while still ensuring valid online FDR control.
1.1 Our contributions
To the best of our knowledge, this is the first work to incorporate feedback information directly into the construction of testing thresholds for online FDR procedures and to extend this idea to the setting of online conformal testing. Our main contributions are twofold:
  • Generalized alpha-investing with feedback (GAIF). GAIF is both comprehensive and flexible: it can enhance almost any existing method within the traditional GAI family, and it naturally extends to more complex scenarios—such as delayed feedback and local dependence. We prove that GAIF-based procedures maintain valid FDR control under independence and valid mFDR control under local dependence.
  • Online conformal multiple testing. We extend GAIF to the setting of online conformal multiple testing, providing explicit construction of valid, independent $p$-values (for null hypotheses) by dynamically updating the calibration dataset in this context. By suitably modifying the GAIF rules, we obtain a finite-sample guarantee of mFDR control for our procedure, online conformal testing with feedback (OCTF). This extension bridges the gap between traditional online multiple testing and conformal inference, yielding distribution-free, model-agnostic tools for real-time decision-making. Moreover, we introduce an online model-selection criterion: predictive models and conformity scores are chosen adaptively based on feedback, further improving power, especially under distribution shifts among non-nulls.
We provide rigorous proofs for all proposed methods. Extensive simulations and real-data experiments demonstrate that our procedures substantially outperform existing approaches while effectively controlling the online FDR when feedback information is provided.
1.2 Related works
Our work is situated at the intersection of online multiple testing and conformal inference. We review key developments in each area and highlight gaps that motivate our contribution.
Online multiple testing under independence.
Early works on online multiple testing began with the alpha-investing strategy of Foster and Stine [2008], later generalized by Aharoni and Rosset [2014] and Javanmard and Montanari [2018] into the generalized alpha-investing (GAI) framework, which led to the LORD algorithm. Building on this line of work, Ramdas et al. [2017] introduced LORD++, an improved version of GAI tailored for online FDR control. Subsequent refinements include SAFFRON [Ramdas et al., 2018], which adapts to the proportion of non-nulls, and ADDIS [Tian and Ramdas, 2019], which adjusts for conservative null $p$-values. These methods, including LORD++, SAFFRON, and ADDIS, achieve online FDR control when null $p$-values are independent of all other $p$-values. Separately, Gang et al. [2021] developed structure-adaptive rules based on local FDR, which improve power but only ensure asymptotic FDR control under correct model specification. For a comprehensive review, see Robertson et al. [2023]. All the above methods determine thresholds solely from past rejections, without considering real-time feedback.
Online multiple testing under dependence.
In practice, hypotheses often exhibit dependence, and applying methods designed for independence can lead to inflated error rates. To address arbitrary dependence, Xu and Ramdas [2024] proposed e-LOND, an FDR-controlling procedure based on e-values. Zhang et al. [2025] extended this approach to e-GAI, achieving improved power by dynamically allocating the testing levels. Alternatively, research has focused on local dependence structures. Zrnic et al. [2021] introduced LORDdep and SAFFRONdep, establishing mFDR control under local dependence; Rebjock et al. [2021] later adapted these methods to time-series anomaly detection. Recently, Fisher [2024] showed that LORD++ with suitable local modifications can maintain FDR control under certain dependence, while Fischer et al. [2024] proposed an online procedure under PRDS dependence. Despite these advances, existing dependence-aware methods do not incorporate any real-time feedback.
Conformal inference and conformal multiple testing.
Conformal inference [Vovk et al., 2005] offers a model-agnostic way to quantify prediction uncertainty. In the multiple testing setting, early works constructed conformal $p$-values and applied the Benjamini–Hochberg (BH) procedure [Benjamini and Hochberg, 1995] to achieve finite-sample FDR control [Bates et al., 2023, Jin and Candès, 2023a]. Subsequent extensions addressed covariate shift [Jin and Candès, 2023b], constrained selection [Wu et al., 2024, Nair et al., 2025], and conditional testing [Wu et al., 2025], as well as model selection [Bai and Jin, 2024, Gui et al., 2025]. However, these contributions remain confined to offline settings. Few efforts to extend conformal multiple testing to the online domain [Huo et al., 2024, Xu and Ramdas, 2024] have so far overlooked the role of feedback information. Although related research on the construction of conformal prediction sets [Gibbs and Candès, 2021, 2024] has considered feedback, multiple testing problems with FDR control remain largely unexplored. One very recent work, by Humbert et al. [2025], establishes asymptotic online FDP control through an online learning strategy, while we achieve feedback-enhanced testing based on the GAI framework.
1.3 Organization of the paper
The remainder of this paper is organized as follows. Section 2 introduces the GAIF procedure and establishes FDR control under independence and mFDR control under local dependence. In Section 3, we construct explicit online conformal $p$-values and apply modified GAIF rules to online conformal testing, providing finite-sample theoretical guarantees. We also extend the proposed framework to achieve model selection and address distribution shifts. Simulation and real-data experiment results are presented in Sections 4 and 5, respectively. Finally, we conclude the paper in Section 6.
2 Generalized Alpha-Investing with Feedback (GAIF)
In this section, we first revisit the traditional generalized alpha-investing (GAI) framework and its extensions in Section 2.1. Then, in Section 2.2, we introduce a new framework named GAIF, that incorporates feedback information. We provide two concrete approaches of GAIF in Section 2.3. In Section 2.4, we establish the finite-sample online FDR control of GAIF procedures under independence. Additionally, we discuss the extensions to address local dependence in Section 2.5.
2.1 Recap: Generalized alpha-investing
The first approaches for controlling error rates in an online setting were based on so-called “alpha-investing” strategies [Foster and Stine, 2008], and subsequently generalized into the generalized alpha-investing (GAI) framework [Aharoni and Rosset, 2014, Javanmard and Montanari, 2018]. The key idea is to compare each incoming $p$-value $p_{t}$ against a dynamically chosen threshold $\alpha_{t}$, to accumulate additional $\alpha$-wealth” upon each rejection, and to make testing decisions according to $\delta_{t}=\mathbb{I}\{p_{t}\leq\alpha_{t}\}$. Based on GAI, Ramdas et al. [2017] proposed controlling the online FDR by ensuring that an estimate of the FDP remains below a pre-specified level $\alpha$. A specific example is LORD++ algorithm, where the estimated $\operatorname{FDP}$ at time $t$ is given by
$\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\overset{(i)}{\geq}\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\approx\operatorname{FDP}(t):=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1},$ (1)
where $\mathcal{H}_{0}$ is the index set of null hypotheses. To ensure that $\operatorname{FDR}(t)\leq\alpha$, it suffices to enforce $\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{R(t)\vee 1}\leq\alpha.$ Building on this, some adaptive versions of LORD++ were proposed subsequently [Ramdas et al., 2018, Tian and Ramdas, 2019]. Specifically, similar to Storey-BH [Storey et al., 2004], Ramdas et al. [2018] proposed SAFFRON with a user-specified parameter $\lambda\in[0,1]$ to estimate the null proportion $\pi_{0}$, giving the estimated FDP as
$\widehat{\operatorname{FDP}}_{\text{SAFFRON}}(t)=\frac{\sum_{j\leq t}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}}{\sum_{j\leq t}\delta_{j}\vee 1}.$
All of the above GAI-based procedures guarantee mFDR control under the conditional super-uniformity assumption for null $p$-values:
$\Pr\bigl{(}p_{t}\leq\alpha_{t}\,\bigm{|}\,\mathcal{G}_{t-1}\bigr{)}\;\leq\;\alpha_{t}\quad\text{for all }t\in\mathcal{H}_{0},$ (2)
where $\mathcal{G}_{t}=\sigma(\delta_{1},\dots,\delta_{t})$ and each threshold $\alpha_{t}=f_{t}(\delta_{1},\dots,\delta_{t-1})$ is measurable with respect to past decisions. Furthermore, if all null $p$-values are independent of other $p$-values, these procedures also ensure strict FDR control.
Note that the inequality (i) in (1) is tight if and only if $\theta_{j}=0$ for all $j\leq t$. This conservative step is one of major sources of gap between the realized FDR of LORD++ and the target level $\alpha$. In our context, since potential feedback becomes available during the testing process, it can serve as a natural remedy.
2.2 Boosting GAI via feedback: GAIF
In our problem setup, at each time step $t$, feedback reveals the true values of $\{\theta_{j}\}_{j=1}^{t-1}$. The key idea lies in leveraging this feedback to reduce statistical slack—enhancing both the accuracy of FDP estimation and the efficiency of alpha-wealth allocation. We formalize this in the following GAI with Feedback (GAIF) framework.
Definition 2.1 (GAIF).
The GAIF procedure refers to any rule for assigning test levels $\alpha_{t}$ such that
$\widehat{\operatorname{FDP}}_{\rm{GAIF}}(t):=\frac{\sum_{j=1}^{t-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})+\alpha_{t}\,\kappa(p_{t})}{1\vee\sum_{j=1}^{t}\delta_{j}}\leq\alpha,$
where $\kappa\colon[0,1]\to\mathbb{R}_{\geq 0}$ denotes a predetermined weight function satisfying
${\mathbb{E}}\{\widehat{\operatorname{FDP}}_{{\rm GAIF}}(t)\}\geq{\mathbb{E}}\!\Bigl{\{}\tfrac{\sum_{j\leq t}(1-\theta_{j})\,\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\Bigr{\}}.$ (3)
The corresponding testing procedures are summarized in Algorithm 1.
Algorithm 1 GAIF procedures
0:  Target $\operatorname{FDR}$ level $\alpha$, pre-specified parameters for constructing test levels
1:  for $t=1,2,\dots$ do
2:   Observe $p$-value $p_{t}$
3:   Update $\alpha_{t}=\alpha_{t}^{\text{GAIF}}$, constructed according to Definition (2.1)
4:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$
5:   Obtain the revealed feedback $\theta_{t}$
6:  end for
6:  Rejection set ${\mathcal{R}}=\{t:\delta_{t}=1\}$.
The GAIF framework improves statistical power through two mechanisms. The first one is leveraging feedback $\{\theta_{i}\}_{i<t}$ to improve the accuracy of FDP estimation. By incorporating feedback, GAIF can identify which previously tested hypotheses are null ($j\leq t-1$), enabling a precise contribution of past tests to the FDP and reducing conservativeness. Importantly, this refinement can be applied to almost all GAI procedures—except for LOND [Javanmard and Montanari, 2015]; see Appendix A.7.
Another mechanism is to efficiently allocate the alpha wealth by adopting the weight function $\kappa_{j}$. Unlike offline multiple testing—where a common threshold applies to all hypotheses—online procedures must determine each $\alpha_{t}$ individually. Therefore, power depends not only on accurate FDP estimation but also on how alpha-wealth is allocated over time. To improve allocation, GAIF introduces data-adaptive weight functions $\kappa(\cdot)$.
Formally, GAIF selects $\alpha_{t}$ to satisfy
$\alpha_{t}\;\leq\;\frac{1}{\kappa(p_{t})}\Biggl{\{}\alpha\sum_{j=1}^{t}\delta_{j}\;-\;\sum_{j=1}^{t-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})\Biggr{\}}$ (4)
which, together with (3), provides the guiding principle for alpha-wealth allocation.
Choice of $\kappa(\cdot)$. The weight function $\kappa(\cdot)$ should satisfy:
  1. 1.
    Nonnegativity: $\kappa(u)\geq 0$ for all $u\in[0,1]$.
  2. 2.
    Null unbiasedness: $\mathbb{E}\{\kappa(U)\}=1$, where $U\sim\mathrm{Uniform}(0,1)$.
  3. 3.
    Concentration on large $p$-values: $\kappa$ places more weight on high $p$-values so that, when nulls prevail, “unused” alphaalpha-wealth is reclaimed and reallocated to future tests.
By these requirements, more $\alpha$-wealth is reserved for tests with small pp–values, increasing power whenever non-null alternatives are present. Indeed, from inequality (4) we see that, when the non-null proportion is large, for small $p_{t}$, the weighted rule produces an $\alpha_{t}$ that strictly exceeds its unweighted counterpart, thereby further boosting power.
Remark 2.1.
In some applications, the feedback is delayed: instead of seeing $\theta_{t-1}$ at time $t$, it arrives at time $t+d$, where $d\geq 0$ denotes the delay. At time $t$, we have $\{\theta_{i}\}_{i=1}^{t-d-1}$. Our GAIF framework can be extended to this case naturally, say we need
$\widehat{\operatorname{FDP}}_{\text{delay}}(t):=\frac{\sum_{j=1}^{t-d-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})+\sum_{j=t-d}^{t}\alpha_{j}\,\kappa(p_{j})}{1\vee\sum_{j=1}^{t}\delta_{j}}\leq\alpha.$
2.3 Two concrete approaches of GAIF: LF & SF
The remaining question is how to choose the weight function $\kappa(\cdot)$ in practice. We address this by presenting two concrete examples: LF and SF.
LF: an initial approach to improve FDP estimation
The simplest choice is setting all $\kappa_{j}(p_{j})=1$ by noting that $\theta_{t}\leq 1$, resulting in a feedback-enhanced LORD++ estimator, $\widehat{\operatorname{FDP}}_{\text{LF}}$ (denoted as LF), given as
$\widehat{\operatorname{FDP}}_{\text{LF}}(t):=\frac{\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})+\alpha_{t}}{\sum_{j\leq t}\delta_{j}\vee 1}.$
Enforcing $\widehat{\operatorname{FDP}}_{\text{LF}}(t)\leq\alpha$ is equivalent to $\sum_{j=1}^{t}\alpha_{j}\leq\sum_{j=1}^{t}\alpha\delta_{j}+\sum_{j=1}^{t-1}\alpha_{j}\theta_{j}$, and therefore we construct the corresponding test levels by:
$\alpha_{t}^{\text{LF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j},$ (5)
where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to one, $\tau_{j}$ is the time of the $j$-th rejection, and $s_{0}>0$ is the pre-specified initial wealth. When $\theta_{j}=0$ for all $j$, this approach reduces exactly to the LORD++.
As the test level of LF $\alpha_{t}^{\text{LF}}$ is not smaller than that of LORD++, this LF procedure would be more powerful than LORD++ by fully incorporating available feedback information. Next, we show that by explicitly utilizing the observed $p$-value patterns in the design of $\kappa(\cdot)$, rather than simply setting it to the constant 1, more accurate FDP estimation and more efficient alpha-wealth allocation can be achieved.
SF: an approach to achieve adaptive $\alpha$-wealth allocation
To exploit patterns in the observed $p$-values, we propose the following concrete instantiation:
$\kappa(p_{j})=\frac{\mathbf{1}\{p_{j}>\lambda\}}{1-\lambda},$ (6)
where $\lambda$ is a user-chosen constant parameter in the interval $[0,1]$ for identifying large $p$-values.
Under a true null hypothesis $\mathbb{H}_{0j}$, $\mathbb{E}\{\kappa(p_{j})\}=1$, thereby satisfying $\operatorname{FDR}(t)\leq{\mathbb{E}}\{\widehat{\operatorname{FDP}}_{\text{GAIF}}(t)\}$. This choice ensures that any “unused” budget is returned for future allocation after screening $p_{j}>\lambda$ in the construction of the $\{\alpha_{t}\}$ sequence, while preserving valid FDR control. When $\theta_{j}=0$ for all $j$, this approach reduces exactly to the SAFFRON algorithm. We call this approach as SF.
The thresholds $\{\alpha_{t}\}$ for SF can be derived analogously to those in Ramdas et al. [2018], but with the tightened, feedback-aware constraint in Eq.(4). For $t=1$, $\alpha_{1}^{\text{SF}}=\min\{\gamma_{1}s_{0},\lambda\}$. For $t>1$,
$\alpha_{t}^{\text{SF}}=\min\left\{\lambda,s_{0}\gamma_{t-C_{0+}}+\{(1-\lambda)\alpha-s_{0}\}\gamma_{t-\tau_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-C_{j+}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j}\right\},$ (7)
where we define $C_{j+}=C_{j+}(t)=\sum_{i=\tau_{j}+1}^{t-1}C_{i}$ and $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$.
The weight function $\kappa(\cdot)$ in Eq.(6) is also used in SAFFRON [Ramdas et al., 2018], where it serves to estimate the non-null proportion and thereby improves FDP estimation. Our GAIF framework provides another intuitive perspective on the role of this function: $\kappa(\cdot)$ can also serve as a guiding mechanism for alpha-wealth allocation, leading to improved power compared to using constant weights. The indicator $\mathbb{I}\{p_{j}>\lambda\}$ ensures that alpha-wealth is spent only on the tests with $p_{j}\leq\lambda$: If $p_{j}\leq\lambda$, then $\kappa(p_{j})=0$. In this case, the FDP bound does not penalize that test, allowing the alpha-wealth to be “invested” there.
In principle, there may exist other concrete examples of $\kappa(\cdot)$ to optimize the alpha-wealth allocation. However, this online optimization remains an open problem and a direction for future work [Gang et al., 2021]. Meanwhile, the SF construction offers a theoretically grounded and practically appealing compromise, achieving robust empirical performance with minimal tuning burden.
Figure 1 depicts the testing thresholds $\{\alpha_{t}\}$ over time $t$ for various procedures applied to Gaussian observations. It is clear that our methods yield larger thresholds after improving the gap via feedback, with $\alpha_{t}^{\text{SF}}>\alpha_{t}^{\text{SAFFRON}}$ and $\alpha_{t}^{\text{LF}}>\alpha_{t}^{\text{LORD++}}$ in average. This illustrates that the GAIF framework leverages alpha-wealth more effectively through feedback, thereby achieving higher power than the traditional GAI framework.
Figure 1: Results for testing Gaussian observations. Line charts depict the average of $\alpha_{t}$ over time $t$ under various testing procedures, based on 500 replications with a non-null proportion of $\pi_{1}=0.5$. Details of the data generation process are provided in Subsection 4.1.
2.4 Theoretical guarantee under independence
Our first theoretical result states that both of LF and SF guarantees online mFDR control under the conditional super-uniformity of null $p$-values
$\Pr\bigl{(}p_{t}\leq\alpha_{t}\,\bigm{|}\,\mathcal{F}_{t-1}\bigr{)}\;\leq\;\alpha_{t}\quad\text{for all }t\in\mathcal{H}_{0},$ (8)
where $\mathcal{F}_{t}:=\sigma(\delta_{1},\dots,\delta_{t};\theta_{1},\dots,\theta_{t})$, and achieves online FDR control under stronger independence and monotonicity assumptions. We say that the test level sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ is a monotonic function of the past if, for all $t\in\mathbb{N}$, $\alpha_{t}$ is coordinatewise nondecreasing in the past decisions $\{\delta_{i}:i<t\}$ and feedback $\{\theta_{i}:i<t\}$ for the LF procedure, and, for the SF procedure, additionally coordinatewise nondecreasing in $\{C_{i}:i<t\}$, where $C_{i}=\mathbb{I}\{p_{i}\leq\lambda\}$.
Theorem 1 (Online mFDR and FDR Control for GAIF).
Let $\{\alpha_{t}\}_{t\in\mathbb{N}}$ be a sequence of test levels such that $\widehat{\operatorname{FDP}}_{\rm{GAIF}}(t)\leq\alpha$, and suppose the weight function $\kappa(\cdot)$ satisfies the condition in Definition 2.1. Then:
  • (a)
    If the null $p$-values are conditionally super-uniform (8), the procedure guarantees $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
  • (b)
    If the null $p$-values are independent of each other and of the non-nulls, and if the test level sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ is a monotonic function of the past for all $t$, then we additionally have $\operatorname{FDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
  • (c)
    In particular, the LF and SF procedures are concrete instances of GAIF and thus satisfy the mFDR and FDR control guarantees under the respective conditions in (a) and (b).
2.5 Addressing local dependence
This subsection focuses on extending our framework to settings with local dependence (as defined in Definition 2.2). For a strategy that addresses a more general dependence structure via e-values, the reader may refer to Appendix D.
Definition 2.2 (Local dependence; Zrnic et al. [2021]).
We say that $p$-values $p_{1},p_{2},\dots,p_{t},\dots$ are local dependent if
$\text{for all}\;t>0,\;\text{there exists}\;L_{t}\in\mathbb{N}\;\text{such that}\;p_{t}\perp p_{t-L_{t}-1},p_{t-L_{t}-2},\dots,p_{1},$ (9)
where $\{L_{t}\}_{t\in\mathbb{N}}$ is a fixed sequence of parameters which we refer to as lags.
Zrnic et al. [2021] pioneered a strategy to handle this local dependency in online multiple testing, leading to the development of the $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$ methods. The following procedures are the modified counterparts of LF and SF under the local dependence:
$\widehat{\operatorname{FDP}}_{\text{LF}_{\text{dep}}}(t)$ $=\frac{\sum_{j\leq t-1}(1-\theta_{j})\alpha_{j}+\alpha_{t}}{(\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j})\vee 1}\leq\alpha,$
$\widehat{\operatorname{FDP}}_{\text{SF}_{\text{dep}}}(t)$ $=\frac{\sum_{j<t-L_{t}}(1-\theta_{j})\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{1-\lambda}+\sum_{j=t-L_{t}}^{t-1}(1-\theta_{j})\frac{\alpha_{j}}{1-\lambda}+\frac{\alpha_{t}}{1-\lambda}}{(\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j})\vee 1}\leq\alpha.$
The corresponding testing levels are presented in Appendix C.1. The $\text{LF}_{\rm dep}$ and $\text{SF}_{\rm dep}$ methods control mFDR under local dependence.
Theorem 2 (Online mFDR control under local dependence).
Suppose that the null $p$-values are locally dependent, as defined in (9). Then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{LF}_{\rm{dep}}}(t)\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\operatorname{SF}_{\rm{dep}}}(t)\leq\alpha$, we have $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
3 Applications on Online Conformal Selection
The online conformal selection [Huo et al., 2024, Xu and Ramdas, 2024] provides a canonical and intuitive instantiation of the GAIF framework. Here, decisions must be made in real time about whether an incoming observation satisfies a pre-specified requirement, while the true label—i.e., the feedback—can generally be observed after the decision. Consider a data pair $({\bf X},Y)\in{\mathcal{X}}\times{\mathcal{Y}}$, with a historical calibration dataset $\mathcal{D}_{{\mathcal{C}}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$ of size $n$. Let $\mathcal{C}$} denote its index set. Test samples $({\bf X}_{t},Y_{t})$ arrive sequentially for $t=1,2,\dots$, where covariates ${\bf X}_{t}$ are observed but responses $Y_{t}$ remain hidden until a real-time decision $\delta_{t}\in\{0,1\}$ is made. The goal at each step is to determine whether $Y_{t}$ lies in a target region ${\mathcal{A}}\subseteq{\mathcal{Y}}$, (e.g., $[a,b]$ or $b,\infty)$ in regression tasks). This can be framed as an online testing problem with $\theta_{t}=\mathbb{I}\{Y_{t}\in{\mathcal{A}}\}$. We will specialize GAIF to this context: we first construct provably independent online conformal p-values, and then adapt GAIF rules to yield finite-sample mFDR control.
3.1 Construction of online conformal $p$-values
To construct $p$-values, we introduce the notion of non-conformity score function $V({\bf X})$, where larger values indicate a higher likelihood that $\theta_{i}=0$. Typically, $V({\bf X})$ is a monotone transformation of the prediction $\widehat{\mu}({\bf X})$, assumed pre-trained to to estimate $Y_{t}$. For example, if $\theta=\mathbb{I}\{Y_{t}\geq b\}$, then one can take $V(X)=b-\widehat{\mu}({\bf X})$. For simplicity, we write $V_{i}=V({\bf X}_{i})$. A natural approach to achieving online FDR control is to compute conformal $p$-values [Bates et al., 2023, Jin and Candès, 2023b] and then apply the GAIF procedure or existing GAI rules. However, this is not directly applicable in the present setting, as standard conformal $p$-values do not satisfy Eq.(8) or independence assumptions.
To circumvent these issues, we adopt online conformal $p$-values [Vovk et al., 2003, Vovk, 2021], which sequentially update the calibration set. Under exchangeability, this yields null $p$-values that are independent, thereby avoiding the complex dependence structures inherent in offline conformal methods, which rely on shared calibration sets and produce $p$-values with PRDS dependence [Bates et al., 2023]. We need the following assumption.
Assumption 1 (Exchangeability in conformal setting).
The null data $\big{\{}({\bf X}_{i},Y_{i}):i>-n,\theta_{i}=0\big{\}}$ are exchangeable conditional on the non-null data $\big{\{}({\bf X}_{i},Y_{i}):i>-n,\theta_{i}=1\big{\}}$.
This assumption is common in conformal inference [Marandon et al., 2024], which is weaker than requiring $\{({\bf X}_{i},Y_{i})\}_{i>-n}$ are independent and identically distributed (i.i.d.).
The strategy is that each test point is incorporated into the calibration set for the next time step. Specifically, at time $t$, the point $({\bf X}_{t},Y_{t})$ serves as the test point, but from time $t+1$ onward, it becomes part of the calibration data. Let ${\mathcal{C}}_{0t}$ denote the dynamically updated index set of calibration samples. Specifically, ${\mathcal{C}}_{0t}$ is given by
${\mathcal{C}}_{0t}=\{-n<i<t:\theta_{i}=0\}.$ (10)
For test data ${\bf X}_{t}$ at time $t$, the online conformal $p$-value $p_{t}$ is defined as
$p_{t}=\frac{\sum_{i\in{{\mathcal{C}}}_{0t}}{\mathbb{I}}\{V_{i}<V_{t}\}+\xi_{t}\cdot\left[1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\{V_{i}=V_{t}\}\right]}{1+|{\mathcal{C}}_{0t}|},$ (11)
where $\xi_{t}\overset{\text{i.i.d.}}{\sim}\operatorname{Unif}[0,1]$ is an independent random variable for tie-breaking.
Similar forms of online conformal $p$-values have been studied for testing exchangeability [Vovk et al., 2003, Vovk, 2021] and constructing online conformal prediction intervals [Angelopoulos et al., 2024]. We formally state the validity and independence under the null in Proposition 3.1.
Proposition 3.1 (Validity and Independence of Online Conformal $p$-values).
Suppose Assumption 1 holds. Then under the null, the online conformal $p$-value $p_{t}$ defined in (11) is uniformly distributed on $[0,1]$, and the null p-values $\{p_{t}:t\in\mathbb{N},\theta_{t}=0\}$ are mutually independent.
Moreover, we further allow the score function to vary over time and be data-dependent, provided it remains symmetric on null data; see Appendix A.3. This flexibility enables leveraging non-null feedback to improve score quality, facilitating model selection and adaptation to distribution shift; see Section 3.3.
3.2 Online conformal testing procedures
Having established the properties of online conformal $p$-values, we now introduce the corresponding testing procedures. Although applying GAIF directly to those $p$-values yields satisfactory FDR control empirically, theoretical guarantees remain lacking. The key challenge is that null online conformal $p$-values are still dependent with those under the alternative. To overcome this issue, we propose modified GAIF rules tailored for online conformal testing, offering rigorous, distribution-free, and model-agnostic mFDR guarantees.
The key idea is to construct the sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ based solely on the rejections of true null hypotheses, rather than on all rejections. While this strategy may be conservative as the discard of non-null decision information, the integration of the feedback provides great compensation. Even when the proportion of non-nulls is relatively large, empirical evidence shows that our modified procedure can still achieve higher power than benchmarks that ignore feedback information.
We revise the rules in LF and SF by replacing $\tau_{j}$ with $\tilde{\tau}_{j}$, resulting in LFS and SFS rules, respectively, where $\tau_{j}$ denotes the time of the $j$-th rejection and $\tilde{\tau}_{j}$ denotes the time of the $j$-th rejection under the null, defined as
$\tilde{\tau}_{j}=\inf\left\{t\in\mathbb{N}:\sum_{i\leq t}\delta_{i}(1-\theta_{i})\geq j\right\}.$
The last letter S in “LFS” and “SFS” indicates “safe” since it guarantees rigorous finite-sample mFDR control. By doing so, $\tilde{\tau}_{j}$ is fixed given past null p-values, as its construction depends only on the rejections of true null hypotheses.
The Online Conformal Testing with Feedback (OCTF) procedure, based on online conformal $p$-values, is detailed in Algorithm 2. Building on the construction in Section 3.1 and the analysis above, we obtain finite-sample online mFDR control for the OCTF procedures, as formalized in Theorem 3.
Theorem 3 (Finite-sample Online mFDR control for OCTF).
Suppose Assumption 1 holds. The outputs $\mathcal{R}$} of Algorithm 2 satisfy $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
With respect to FDR control, the additional challenge lies in the fact that a past non-null decision $\delta_{i}=\mathbb{I}\{p_{i}\leq\alpha_{i}\}$ with $\theta_{i}=1$ may be dependent on a null $p_{t}$. FDR control requires decoupling the dependence between decisions $\sum_{i\leq t}\delta_{i}$ in the denominator and $p_{t}$, which is a more difficult task and warrants further research.
Algorithm 2 Online conformal testing with feedback (OCTF)
0:  Initial calibration data ${\mathcal{D}}_{\mathcal{C}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, non-conformity score function $V(\cdot)$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}_{t\in\mathbb{N}}$, constant $\lambda\in(0,1)$.
1:  Initialize ${\mathcal{C}}_{01}=\{i\in{\mathcal{C}}:\theta_{i}=0\}$
2:  for $t=1,2,\ldots$ do
3:   Observe test data ${\bf X}_{t}$
4:   Compute conformity scores $\{V_{i}\}_{i\in{\mathcal{C}}_{0t}\cup\{t\}}$
5:   Compute online conformal $p$-value $p_{t}$ via (11)
6:   Update $\alpha_{t}={\alpha}_{t}^{\text{LFS}}$ in Eq. (25) (or $\alpha_{t}={\alpha}_{t}^{\text{SFS}}$ in Eq.(26))
7:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$
8:   Obtain the revealed feedback $Y_{t}$ (and thus obtain $\theta_{t}$)
9:   Update the calibration dataset ${\mathcal{C}}_{0t}$.
10:  end for
10:  Rejection set $\mathcal{R}=\{t:\delta_{t}=1\}$.
3.3 Optimized online conformal multiple testing with feedback
The performance of online conformal testing depends on the combined choice of the non-conformity scores $V({\bf X})$ and the prediction model (e.g., random forests, neural networks, support vector machines, or regularized linear models) [Bai and Jin, 2024, Gasparin and Ramdas, 2024]. We collectively refer to these choices as model selection. Moreover, in practice, data distributions in non-null may shift over time: the $\{({\bf X}_{t},Y_{t})\}_{t\in\mathbb{N}}$ remains exchangeable under the null but may change in distribution under the alternative [Gang et al., 2021, Huo et al., 2024]. Dynamically selecting models adapted to the current distribution can substantially alleviate the shifting effects. This motivates us to design a versatile framework that ensures error rate control in online conformal testing while allowing flexible data reuse for model optimization and ultimately enhancing power.
Suppose there are $K$ pre-trained candidate score functions $\{V(\cdot;k):{\mathcal{X}}\to\mathbb{R}\}_{k=1}^{K}$. At each time $t$, a score is chosen as
$\widehat{k}_{t}=\underset{k\in[K]}{\arg\min}\;{\mathcal{M}}(k,{\mathcal{D}}_{t}),$
where ${\mathcal{M}}(k,{\mathcal{D}}_{t})$ is a given criterion evaluated by the currently observed data ${\mathcal{D}}_{t}=\{({\bf X}_{i},Y_{i}):-n<i<t\}\cup\{{\bf X}_{t}\}$. And the $p$-value is generated by the optimized score function $V(\cdot;\widehat{k}_{t})$ for our online testing procedure. However, naively optimizing the predictive model can yield invalid online conformal $p$-values or introduce intricate dependencies, ultimately leading to FDR inflation [Zhang et al., 2022]. To ensure validity and independence of online conformal $p$-values after model-optimization, we propose a model evaluation criteria specifically designed for the online setting with feedback. In particular, our adaptive model selection approach accommodates distribution shifts in the non-null data by dynamically selecting the best model based on the recent alternative distribution.
Considering the availability of real-time feedback on past true signals, we propose using the exponentially weighted moving average (EWMA) of past auxiliary non-null $p$-values as the model selection criterion. At each time $t$, the optimal model is chosen by minimizing this EWMA value, which serves as a good proxy for the unknown current non-null $p$-value. The rationale of this criterion is that a better model should yield smaller non-null $p$-values, leading to more powerful procedures.
Specifically, at time $t$, for each score function $k$, we construct auxiliary non-null $p$-values as
$\tilde{p}^{k}_{j}=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|},\quad j\in{\mathcal{C}}_{1t},$ (12)
where ${\mathcal{C}}_{1t}=\{-n<i<t:\theta_{i}=1\}$ denotes the set of online non-null samples observed prior to time $t$. The construction in (12) is carefully designed so that the resulting auxiliary non-null $p$-values are invariant under permutations of ${\mathcal{C}}_{0t}\cup\{t\}$, which is crucial for ensuring valid inference after model selection. We then define the model evaluation criterion as
${\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t})=\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,\tilde{p}^{k}_{j}\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}},$
where $\rho\in(0,1)$ is the user-specified decay parameter that downweights past observations, enabling dynamic adaptation to the recent non-null distribution. The optimal model is accordingly selected as $\widehat{k}_{t}={\arg\min}_{k\in[K]}{\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t}).$
Finally, the online conformal $p$-value after model optimization, denoted as $p^{\rm opt}_{t}$, is computed by Eq.(11) with $V(\mathbf{X}_{t};\widehat{k}_{t})$ instead of $V_{t}$, and then the OCTF procedures in Section 3.2 can be implemented with these optimized online conformal $p$-values. The entire optimized online conformal testing procedure (Opt-OCTF) is summarized in Algorithm 3. Intuitively, this approach helps the $p$-value construction dynamically adapt as new data arrives, while ensuring that the past samples are still considered, but with diminishing influence. By continually learning the current non-null pattern, our procedure remains powerful against potential non-null distribution shifts.
Algorithm 3 Optimized Online Conformal Testing with Feedback (Opt-OCTF)
0:  Initial calibration data ${\mathcal{D}}_{\mathcal{C}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, $K$ pretrained models $\{\widehat{\mu}_{k}\}_{k=1}^{K}$ and the non-conformity score function $V(\cdot;k)$, evaluating criterion $\mathcal{M}={\mathcal{M}}^{\text{EWMA}}$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}_{t\in\mathbb{N}}$, constant parameters $\rho\in(0,1)$, $\lambda\in(0,1)$.
1:  Initialize ${\mathcal{C}}_{01}=\{i\in{\mathcal{C}}:\theta_{i}=0\}$
2:  for $t=1,2\ldots$ do
3:   Observe test data ${\bf X}_{t}$
4:   Decide the predictive model for $t$-th test sample by $\widehat{k}_{t}={\arg\min}_{k\in[K]}{\mathcal{M}}^{\text{EWMA}}(k,{\mathcal{D}}_{t})$
5:   Construct optimized online conformal p-value $p^{\rm opt}_{t}$
6:   Update $\alpha_{t}=\alpha_{t}^{\text{LFS}}$ in Eq.(25) (or $\alpha_{t}=\alpha_{t}^{\text{SFS}}$ in Eq.(26))
7:   Make a decision $\delta_{t}=\mathbb{I}\{p_{t}^{\text{opt}}\leq\alpha_{t}\}$
8:   Obtain the revealed feedback $Y_{t}$ (and thus obtain $\theta_{t}$)
9:   Update the calibration dataset ${\mathcal{C}}_{0t}$.
10:  end for
10:  Rejection set $\mathcal{R}_{\text{opt}}=\{t:\delta_{t}=1\}$.
Importantly, the optimized online conformal $p$-values still remain valid and mutually independent under the null hypothesis, which is an extension of Proposition 3.1 by allowing that the score function at each time to be dependent on the non-null data. As null $p$-values are excluded from the computation of EWMA, we can conclude the score function $V(\cdot;\hat{k}_{t})$ is symmetric to the null data up to time $t$. Consequently, Opt-OCTF still achieves finite-sample FDR control, as established in the following Corollary 3.1.
Corollary 3.1 (Finite-sample Online mFDR control for Opt-OCTF).
Under the same assumptions of Theorem 3, the Optimized Online Conformal Testing with Feedback by criterion ${\mathcal{M}}^{\rm EWMA}$ (Opt-OCTF) procedure in Algorithm 3 satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}.$
4 Numerical Simulations
In this section, we present extensive synthetic experiments to demonstrate the validity and efficiency of our proposed methods. First, Section 4.1 reports results for online multiple testing in the non-conformal setting. Second, Sections 4.2 and 4.3 present the results under the conformal setting for OCTF and Opt-OCTF, respectively. For the conformal setting, we focus on binary classification scenarios; additional results, including a regression scenario, are provided in Appendix F. Following the setup in Robertson et al. [2023], we fix the pre-specified parameters as $\lambda=0.5$, $\gamma_{j}\propto j^{-1.6}$ for all $j\in\mathbb{N}$, and $s_{0}=\alpha/2$, and use these settings consistently across all experiments.
4.1 Synthetic experiments under non-conformal settings
We begin with experiments in traditional online multiple testing problems, covering three distinct scenarios:
  • Scenario I (Testing with Gaussian observations) We simulate $T$ independent test statistics $Z_{t}\sim N(\mu_{t},1)$ with hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$. One-sided $p$-values are given by $p_{t}=\Phi(-Z_{t})$, where $\Phi$ is the standard Gaussian CDF. The signal strengths $\mu_{t}$ follow a mixture model:
    $\mu_{t}=\begin{cases}0\quad\text{with probability}\;1-\pi_{1}\\ F_{1}\quad\text{with probability}\;\pi_{1},\end{cases}$ (13)
    where the random variable $F_{1}\sim N(\mu_{c},1)$ and $\mu_{c}=2.5$.
  • Scenario II (Testing with Beta alternatives) We generate $T$ $p$-values according to the following model:
    $p_{t}\sim\begin{cases}\text{Unif}[0,1]\quad\text{with probability}\;1-\pi_{1}\\ \text{Beta}(0.5,4)\quad\text{with probability}\;\pi_{1}.\end{cases}$ (14)
  • Scenario III (Testing under local dependence) We simulate $T$ correlated test statistics $(Z_{1},\dots,Z_{T})\sim N(\bm{\mu},\Sigma)$ where $\bm{\mu}=(\mu_{1},\ldots,\mu_{T})$ with $\mu_{t}=2$ for randomly a fraction $\pi_{1}$ of indices and $\mu_{t}=0$ otherwise. The covariance matrix $\Sigma$ has a block-diagonal structure: coordinates are grouped into blocks of size $n_{\mathrm{block}}=10$, with within-block correlation $\rho=0.8$ and diagonal elements as 1. We test hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$ using one-sided $p$-values $p_{t}=\Phi(-Z_{t})$.
We set $T=1000$ and the FDR level at $\alpha=0.1$. We compare our methods SF and LF with state-of-the-art algorithms for online FDR control, namely LOND [Javanmard and Montanari, 2015], LORD++ [Ramdas et al., 2017], and SAFFRON [Ramdas et al., 2018], using their default parameters ($\lambda=0.5$ chosen for SAFFRON). In the locally dependence setting, we additionally include the proposed $\textbf{SF}_{\text{dep}}$ and $\textbf{LF}_{\text{dep}}$ along with the existing $\textbf{SAFFRON}_{\text{dep}}$ and $\textbf{LORD}_{\text{dep}}$ from Zrnic et al. [2021]. We evaluate the performance via empirical FDR and power averaged over 500 independent replications. Appendix F confirms that mFDR and FDR show similar trends.
Figure 2 shows results for varying non-null proportion $\pi_{1}\in[0.1,0.8]$ under Scenarios I and II. All benchmark methods ensure valid FDR control across different $\pi_{1}$ under both Scenarios. Our feedback-based SF and LF methods enhance detection power while maintaining valid online FDR control: SF significantly outperforms SAFFRON, while LF yields higher power than LORD++ and surpasses SAFFRON when $\pi_{1}$ is small. In contrast, both LORD++ and LOND remain conservative across different values of $\pi_{1}$.
[…]
Figure 2: Results for Scenario I and Scenario II. Line charts of FDR and Power at stopping time with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$ after $500$ replications; The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
Figure 3: Results for Scenario III: FDR and power at stopping time across 500 replications with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
4.2 Results for online conformal testing with feedback
We next evaluate OCTF with a fixed score in a real-time binary classification task.
  • Scenario IV: Data is generated as $\mathbf{X}\mid Y=0\sim\mathcal{N}_{4}\left(\bm{\mu}_{1},\mathbf{I}_{4}\right)$, and $\mathbf{X}\mid Y=1\sim\mathcal{N}_{4}\left(\bm{\mu}_{2},\mathbf{I}_{4}\right)$, where $\bm{\mu}_{1}=(1,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. The target region is $\mathcal{A}=\{1\}$.
We set the pre-specified stopping time to $T=1000$ and target $\operatorname{FDR}$ level to $\alpha=0.2$. The predictive model is a neural network trained on $n_{\text{tr}}=1000$, and the initial calibration set has size $n_{\text{cal}}=1000$. Appendix F presents additional results for varying $n_{\text{cal}}$ and alternative training algorithms $\widehat{\mu}$, along with the corresponding results for a regression task.
All methods are implemented within our proposed OCTF workflow; the only difference lies in how the sequence $\{\alpha_{t}\}$ is generated for each benchmark. For clarity, we refer to each method by its corresponding $\alpha_{t}$-generating algorithm and omit the “OCTF” prefix. We compare our proposed methods (SF, SFS, LF, and LFS) with existing approaches: SAFFRON, LORD++, and LOND. Performance is evaluated in terms of empirical FDR and power at $T$, averaged over 500 replications.
Figure 4 reports the online FDR and power at the stopping time $T$ under Scenario IV across varying non-null proportions $\pi_{1}\in[0.1,0.8]$. All methods control the FDR below the nominal level $\alpha$, with SF aligning most closely with the target level among all competitors. In terms of power, as expected, SF consistently achieves the highest power, while LF also performs competitively and attains higher power than SAFFRON across all $\pi_{1}$. The LOND algorithm exhibits the lowest power. Thanks to the incorporation of feedback—reflected in the term $\sum_{j\leq t}\gamma_{t-j}\alpha_{j}\theta_{j}$ within the testing levels—SFS and LFS outperform LORD++, despite discarding certain rejections under the non-null to ensure theoretical guarantees. This safe strategy mitigates the risk of FDR inflation in complex real-data scenarios while still leveraging feedback to enhance power.
Figure 4: Results for Scenario IV: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.2$. Shaded areas show $\pm 1$ standard error.
4.3 Results for optimized online conformal testing with feedback
We now demonstrate the effectiveness of the proposed model optimization strategy from Section 3.3 for selecting predictive models in online conformal testing with FDR control, under scenarios where the non-null distribution shifts over time. Additional results for scenarios without distribution shifts are presented in Appendix F. We consider a binary classification example with the pre-specified stopping time $T=1000$.
  • Scenario V: Data are generated as $\mathbf{X}\mid Y=0\sim\mathcal{N}_{4}\left(\bm{\mu}_{1},\mathbf{I}_{4}\right)$, and $\mathbf{X}\mid Y=1\sim\mathcal{N}_{4}\left(\bm{\mu}_{2},\mathbf{I}_{4}\right)$, where $\bm{\mu}_{1}=(2,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. The target region is $\mathcal{A}=\{1\}$.
For the historical labeled dataset, the non-null proportion is fixed as we fix $\pi_{t}=\operatorname{Pr}(Y_{t}=1)=0.2$. For the testing data, we consider a sine pattern shifts, where the non-null proportion varies as $\pi_{t}=\{\sin(8\pi t/T)+1\}/4$, oscillating between 0 and 0.5. We consider $K=3$ candidate models: random forest (RF), neural network (NN), and support vector machine (SVM). The model evaluation criterion is based on ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with exponential weight parameter $\rho=0.95$. The target FDR level is $\alpha=0.1$. To reduce computational cost, we restrict the auxiliary non-null $p$-value computation to a sliding window of fixed length $L=200$: for $t>L$, we define ${\mathcal{C}}_{1t}=\{i:t-L\leq i<t,\theta_{i}=1\}$; for $t\leq L$, we use the full set ${\mathcal{C}}_{1t}=\{i<t:\theta_{i}=1\}$.
We evaluate several methods that, at each time $t$, select a conformity score from $K$ candidates, compute the corresponding conformal $p$-values, and apply a specific multiple testing procedure. In particular, we consider four proposed procedures (SF, SFS, LF, and LFS), each in two variants: an optimized version (“Opt-”) using our model selection strategy, and a random version (“Ran-") with randomly selected scores. As benchmarks, we include SAFFRON and LORD++, both using random score selection.
The results for Scenario V are shown in Figure 5. All methods successfully control the empirical FDR around the nominal level $\alpha=0.1$. In terms of power, the optimized (Opt) variants consistently outperform their randomly selected (Ran) counterparts, while both Opt-LFS and Ran-LFS achieve higher power than LORD++. Both Opt-SFS and Ran-SFS procedures are more conservative, exhibiting lower power than SAFFRON due to discarding rejections under the non-null to ensure theoretical guarantees. Although this strategy reduces power in this synthetic setting, we will later observe that Opt-SF may suffer from FDR inflation on real data due to the lack of theoretical guarantees.
Overall, these results highlight the effectiveness of our model optimization strategy. Feedback-enhanced methods achieve significantly higher power than feedback-ignoring benchmarks. In contrast, random model selection results in a substantial loss of power, underscoring the importance of informed model choice particularly under distribution shifts.
Figure 5: Results for Scenario V (sine pattern shifts): the values of $\operatorname{FDR}(t)$ and $\operatorname{Power}(t)$ across different time $t$. The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
5 Real Data Applications
In this section, we evaluate our proposed methods on four real-world datasets, illustrating their practical benefits in diverse online decision-making tasks.
  • Task 1: Online Candidate Screening. The first task is real-time candidate screening for selecting the candidates who can get into the interview process. We use the recruitment dataset [Kaggle, 2020], which contains $N=45,372$ candidates with 11 attributes including education status, handicapped or not and gender. The target binary variable is whether the candidate passes the job interview.
  • Task 2: High-Risk Diabetes Identification. The second task focuses on health screening, where selecting individuals at high risk of diabetes is critical for early intervention. We use the diabetes health indicators dataset [Kaggle, 2021], consisting of $N=70,692$ samples and 22 covariates including demographic attributes (e.g., sex, age, BMI), lifestyle-related features, and several binary health indicators. The target binary variable is whether an individual has diabetes.
  • Task 3: High-Income Individual Selection. The third task involves using the 1994 Census Bureau dataset [Becker and Kohavi, 1996] to identify a subset of individuals with high incomes (i.e., income >50K>50K) for precision marketing purposes. This dataset contains $N=32,561$ records with 14 attributes including gender, race, marital status, education level, and more.
  • Task 4: Airfoil Noise Detection. The final task involves using the airfoil dataset [Brooks et al., 2014] from the UCI Machine Learning Repository to identify samples with high sound pressure. This dataset contains $N=1,503$ observations with five physical covariates (log frequency, angle of attack, chord length, free-stream velocity, and suction side log displacement thickness). The response variable $Y$ represents the scaled sound pressure, and we aim to test $\mathbb{H}_{0t}:Y_{t}\in[-\infty,c)$, where $c$ is the $(1-\pi_{1})$-quantile of $Y$, with $\pi_{1}=0.4$.
For Tasks 1-3, we randomly sample three subsets from the full dataset: $n_{\text{tr}}=1{,}000$ training, $n_{\text{cal}}=1{,}000$ calibration, and $n_{\text{te}}=1{,}000$ test samples. For Task 4, we set $n_{\text{tr}}=300$, $n_{\text{cal}}=300$, and $n_{\text{te}}=1000$. We compare four optimized procedures Opt-SF, Opt-SFS, Opt-LF and Opt-LFS with two benchmarks SAFFRON and LORD++. At each time step, $K=3$ candidate training algorithms (RF, NN, and SVM) were employed, and the model evaluation criterion is ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with $\rho=0.9$. For our methods, we use the optimized conformity score and for the benchmarks, we randomly choose a score at each time $t$. All results are averaged over 500 replications.
Figure 6 summarizes the results across time $t$ for all four tasks. Table 1 reports the empirical FDR and power at the stopping time $T=1,000$. In Tasks 1-3, all methods control the empirical FDR near or below the nominal level, and, Opt-SF consistently achieves the highest power for all cases as expected. However, in task 4, Opt-SF shows mild FDR inflation, whereas Opt-SFS and Opt-LFS maintain valid control at all time points, demonstrating the necessity of the safe feedback strategies in finite-sample error-rate control guarantees. Notably, SAFFRON and LORD++ deliver substantially lower power than our optimized methods in all tasks, underscoring the strength of the OCTF framework.
In sum, our strategies exploit feedback in three key ways: refining FDP estimation within the GAIF framework, updating online calibration sets adaptively, and optimizing model selection. Empirically, these feedback-aware designs consistently outperform existing GAI methods that ignore feedback, achieving substantially higher power across diverse real-data tasks while preserving rigorous online FDR control.
Figure 6: Results for real-data applications: the values of $\operatorname{FDR}(\bm{\delta}^{t})$ and $\operatorname{Power}(\bm{\delta}^{t})$ over time $t$ for six benchmarks. The black dashed lines indicate the FDR level $\alpha=0.3$. Shaded areas show $\pm 1$ standard error.
Table 1: Average $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ for different tasks across four datasets (Candidate, Diabetes, Income, Airfoil). The target FDR level is $\alpha=0.3$. Bold numbers represent the best results.
Method Task 1 Task 2 Task 3 Task 4
FDR Power FDR Power FDR Power FDR Power
Opt-SF 0.161 0.247 0.238 0.517 0.287 0.398 0.363 0.726
Opt-SFS 0.124 0.083 0.141 0.166 0.132 0.168 0.239 0.440
Opt-LF 0.148 0.221 0.188 0.341 0.254 0.336 0.242 0.581
Opt-LFS 0.127 0.103 0.132 0.150 0.117 0.171 0.060 0.170
SAFFRON 0.112 0.007 0.185 0.160 0.165 0.083 0.251 0.355
LORD++ 0.086 0.004 0.104 0.004 0.078 0.052 0.098 0.193
6 Concluding Remarks
We study online multiple testing with feedback, aiming to develop reliable machine learning methods for real-time decision-making with rigorous FDR and mFDR control. Our key contribution is GAIF, a feedback-enhanced Generalized Alpha-Investing framework that guarantees online FDR control under the standard assumption of independence between null and all other $p$-values. Building on GAIF, we extend the framework to the conformal setting by constructing independent online conformal $p$-values for null hypotheses, and achieve finite-sample mFDR control via our OCTF procedures. To further enhance performance and address distribution shifts among non-nulls, we propose a feedback-driven model selection criterion.
We highlight several potential directions for future work. First, a current limitation of our framework is its focus on distribution shifts in the alternative data. To address alpha-death and piggybacking under such shifts, one can control weighted FDR using forgetting factors or decaying memory mechanism [Ramdas et al., 2017]. Second, extending our framework to accommodate more general forms of distribution shift is another promising direction. One potential approach is to construct online weighted conformal $p$-values [Prinster et al., 2025] and apply them within the OCTF framework. Finally, to establish guarantees under weaker assumptions, one could relax the definition of valid online FDR control and develop new error-rate notions—analogous to the average-coverage criterion studied in online conformal inference [Gibbs and Candès, 2021, Humbert et al., 2025].
Supplementary Material for “Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection”
This supplementary material contains:
  • Proofs of all the theoretical results (Appendix A).
  • Preliminary terms for self-containment (Appendix B);
  • Additional details of our algorithms (Appendix C);
  • Extensions of GAIF based on $e$-values (Appendix D).
  • Applications on real-time LLM alignment (Appendix E).
  • Additional experiments results (Appendix F).
Appendix A Technical details
A.1 Proof of Theorem 1
Our proof of Theorem 1 relies on the following lemmas, which are modified versions of Lemma 1 in Ramdas et al. [2017] and Lemma 1 in Ramdas et al. [2018], respectively. The modifications arise because, in our setting, the feedback information $\{\theta_{j}\}_{j=1}^{t-1}$ is available at time $t$, and the test levels $\alpha_{t}^{\text{GAIF}}$ depend on both past rejections $\delta_{j}$ and feedback $\theta_{j}$. The proof of Lemma 1 is analogous to that of Lemma 1 in Ramdas et al. [2017] and is omitted; likewise, Lemma 2 follows Lemma 1 in Ramdas et al. [2018] and is also omitted.
Lemma 1.
Suppose that the $p$-values $\{p_{t}\}_{t=1}^{T}$ are independent and let $g:\{0,1\}^{T}\to\mathbb{R}$ be any coordinate-wise non-decreasing function. And $\mathcal{F}_{t-1}=\sigma(\delta_{1},\dots,\delta_{t-1};\theta_{1},\dots,\theta_{t-1})$. Then for any index $t\leq T$ such that $H_{t}\in\mathcal{H}_{0}$, we have
${\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1},\dots,\delta_{t-1};\theta_{1},\dots,\theta_{t-1})\}}{g(\delta_{1},\dots,\delta_{t};\theta_{1},\dots,\theta_{t})}\mid\mathcal{F}_{t-1}\right]\leq{\mathbb{E}}\left[\frac{f_{t}(\delta_{1},\dots,\delta_{t-1};\theta_{1},\dots,\theta_{t-1})}{g(\delta_{1},\dots,\delta_{t};\theta_{1},\dots,\theta_{t})}\right].$
Lemma 2.
Assume that the $p$-values $\{p_{t}\}_{t=1}^{T}$ are independent and let $g(0,1)^{T}\to\mathbb{R}$ be any coordinatewise non-decreasing function. Let $\delta_{1:t}:=(\delta_{1},\dots,\delta_{t})$, and $C_{1:t}:=(C_{1},\dots,C_{t})$, $\theta_{1:t}:=(\theta_{1},\dots,\theta_{t})$, where $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$. Then, for any index $t\leq T$ such that $\theta_{t}=0$, we have
${\mathbb{E}}\left[\frac{f_{t}(\delta_{1:t-1},C_{1:t-1},\theta_{1:t-1})\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)g(\delta_{1:t};\theta_{1:t})}\mid\mathcal{F}_{t-1}\right]$
$\geq{\mathbb{E}}\left[\frac{f_{t}(\delta_{1:t-1},C_{1:t-1},\theta_{1:t-1})}{g(\delta_{1:t};\theta_{1:t})}\mid\mathcal{F}_{t-1}\right]$
$\geq{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1},C_{1:t-1},\theta_{1:t-1})\}}{g(\delta_{1:t};\theta_{1:t})}\mid\mathcal{F}_{t-1}.\right]$
Proof of Theorem 1. We first prove the results for general GAIF procedures. Note that for any time $t\in[T]$, we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid{\mathcal{F}}_{t-1}\right]\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\kappa(p_{j})+\alpha_{t}\kappa(p_{t})\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where the first inequality follows from the law of iterated expectations by conditioning on $\mathcal{F}_{t-1}$ and then applying the conditional super-uniformity property, and the second inequality follows by the fact that $1-\theta_{t}\leq 1$ and the definition of $\kappa(\cdot)$, and the last inequality follows from the construction that $\widehat{\operatorname{FDP}}_{\text{GAIF}}\leq\alpha$. Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for GAIF.
Furthermore, under the independence and the monotonicity assumptions, we have
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[\frac{(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1}\mid{\mathcal{F}}_{t-1}\right]\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}(1-\theta_{j})\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$\leq$ $\frac{\sum_{j=1}^{t-1}(1-\theta_{j})\,\alpha_{j}\,\kappa(p_{j})+\alpha_{t}\,\kappa(p_{t})}{1\vee\sum_{j=1}^{t}\delta_{j}}$ (15)
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\operatorname{GAIF}}(t)\right]$
$\leq$ $\alpha,$
where the first inequality follows form the law of iterated expectations by conditioning on $\mathcal{F}_{t-1}$ and Lemma 1, the second inequality follows from definition of GAIF, and the last inequality follows from the construction of $\alpha_{t}^{\operatorname{GAIF}}$, which completes the proof of FDR control.
The results for LF can be proved accordingly by setting $\kappa_{j}(\cdot)=1$ for all $j\leq t$.
We then prove the results for SF in a similar way. The conditional super-uniformity can be rephrased as:
${\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}>\alpha_{t}\}}{1-\alpha_{t}}\mid{\mathcal{F}}_{t-1}\right]\geq 1\geq{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq\alpha_{t}\}}{\alpha_{t}}\mid{\mathcal{F}}_{t-1}\right].$ (16)
Note that for any time $t\in[T]$, we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathbb{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$ (17)
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid{\mathcal{F}}_{t-1}\right]\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}(1-\theta_{t})\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}.\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}.\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where the first inequality follows from the law of iterated expectations by conditioning on $\mathcal{F}^{j-1}$ and then applying the conditional super-uniformity property, and the second inequality follows by applying the conditional super-uniformity property and Lemma 2, and the third inequality holds since the fact that $1-\theta_{t}\leq 1$, and the last inequality follows from the construction such that $\widehat{\operatorname{FDP}}_{\text{SF}}\leq\alpha$. Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for SF.
Under the independence and the monotonicity assumptions, we have
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[\frac{(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1}\mid{\mathcal{F}}_{t-1}\right]\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}(1-\theta_{j})\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$\leq$ ${\mathbb{E}}\left[\frac{\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}(1-\theta_{t})\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$\leq$ ${\mathbb{E}}\left[\frac{\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}}{\sum_{j\leq t}\delta_{j}\vee 1}\right]$
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\operatorname{SF}}(t)\right]$
$\leq$ $\alpha,$
[…]
A.2 Proof of Theorem 2
Proof.
Define locally conditional super-uniformity as follows: if the null hypothesis $H_{t}$ is true, then for all $\alpha_{t}\in[0,1]$,
$\Pr\left(p_{t}\leq\alpha_{t}\mid{\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}\right)\leq\alpha_{t},$
where ${\mathcal{X}}_{\text{dep}}^{t}:=\{t-L_{t},\dots,t-1\}$, and ${\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}:=\sigma(\delta_{1},\dots,\delta_{t-L_{t}-1};\theta_{1},\dots,\theta_{t-L_{t}-1})$. This condition is immediately true by local dependence [Zrnic et al., 2021]. Note that for any time $t\in[T]$, we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathbb{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$ (18)
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid{\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}\right]\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}[(1-\theta_{j})\alpha_{j}]={\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})+\alpha_{t}\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\right],\$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where the first inequality follows from the law of iterated expectations by conditioning on $\mathcal{F}_{\text{dep}}^{-{\mathcal{X}}^{t}}$ and then applying the conditional super-uniformity property and by noticing that the measurability of $\alpha_{j}$ with respect to $\mathcal{F}_{\text{dep}}^{-{\mathcal{X}}^{t}}$, and the second inequality follows by the fact that $1-\theta_{t}\leq 1$, and the third inequality follows from the construction such that $\widehat{\operatorname{FDP}}_{\text{LF}_{\text{dep}}}\leq\alpha$. Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for $\text{LF}_{\text{dep}}$.
Similar results can be obtained for $\text{SF}_{\text{dep}}$ as follows:
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathbb{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$ (19)
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid{\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}\right]\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}[(1-\theta_{j})\alpha_{j}]={\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}(1-\theta_{t})\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j<t-L_{t}}(1-\theta_{j})\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{1-\lambda}+\sum_{j=t-L_{t}}^{t-1}(1-\theta_{j})\frac{\alpha_{j}}{1-\lambda}+\frac{\alpha_{t}}{1-\lambda}\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\right]\$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right].$
A.3 Proof of Proposition 3.1
We first restate Proposition 3.1 as follows and finish the proof. The proof of Proposition A.1 essentially follows the argument of Theorem 8.2 in Angelopoulos et al. [2024], with an extension to our online conformal testing setting.
Proposition A.1 (Validity and Mutual Independence of Online Conformal $p$-values under Exchangeability and Symmetric Scores).
Suppose at each time $t$, the score function $V(\cdot;{\mathcal{D}}_{t})$ is constructed through the current data ${\mathcal{D}}_{t}=\big{(}({\bf X}_{i},Y_{i}):-n\leq i\leq t\big{)}$. The p-value of each time $t$ is constructed as
$p_{t}=\frac{\sum_{i\in{{\mathcal{C}}}_{0t}}{\mathbb{I}}\{V({\bf X}_{i};{\mathcal{D}}_{t})<V({\bf X}_{t};{\mathcal{D}}_{t})\}+\xi_{t}\cdot(1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\{V({\bf X}_{i};{\mathcal{D}}_{t})=V({\bf X}_{t};{\mathcal{D}}_{t})\})}{1+|{\mathcal{C}}_{0t}|}.$ (20)
Suppose Assumption 1 holds and the score function $V(\cdot;{\mathcal{D}}_{t})$ is symmetric to $\{({\bf X}_{i},Y_{i}):-n\leq i\leq t,\theta_{i}=0\}$. Then under the null hypothesis,
  1. 1.
    Each $p_{t}$ is marginally uniformly distributed on $[0,1]$.
  2. 2.
    The sequence $\{p_{t}\}$ is mutually independent.
Proof.
We adapt the standard argument for the validity of conformal $p$-values to our online setting with exchangeable data. For each time $t$, define
$\Phi_{t}=\Biggl{(}\{({\bf X}_{i},Y_{i}):i\in{\mathcal{C}}_{0t}\cup\{t\}\},\,(\theta_{i}:i\leq t),\,\bigl{(}({\bf X}_{i},Y_{i}):i\in{\mathcal{C}}_{1t}\bigr{)}\Biggr{)},$
where $\Phi_{t}$ contains the unordered set of conformity scores for indices in ${\mathcal{C}}_{0t}\cup\{t\}$, the true state $\theta_{i}$, and the data corresponding to indices not used in the calibration set (denoted here by ${\mathcal{C}}_{1t}$). To prove the mutual independence, consider any time indices $t\leq T$ and let $x_{t},x_{t+1},\dots,x_{T}\in[0,1]$ be arbitrary. Then
$\Pr\bigl{(}p_{t}\leq x_{t},\,p_{t+1}\leq x_{t+1},\,\dots,\,p_{T}\leq x_{T}\bigr{)}$
$=$ $\;\mathbb{E}\Bigl{[}\mathbf{1}\{p_{t}\leq x_{t}\}\mathbf{1}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbf{1}\{p_{T}\leq x_{T}\}\Bigr{]}$
$=$ $\;\mathbb{E}\Biggl{[}\mathbb{E}\Bigl{[}\mathbf{1}\{p_{t}\leq x_{t}\}\,\Big{|}\,\Phi_{t}\Bigr{]}\cdot\mathbf{1}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbf{1}\{p_{T}\leq x_{T}\}\Biggr{]}.$
The key observation is that, by exchangeability, the conditional distribution of $p_{t}$ given $\Phi_{t}$ is uniform on $[0,1]$, so that
$\mathbb{E}\Bigl{[}\mathbf{1}\{p_{t}\leq x_{t}\}\,\Big{|}\,\Phi_{t}\Bigr{]}=x_{t}.$
Thus,
$\Pr\bigl{(}p_{t}\leq x_{t},\,p_{t+1}\leq x_{t+1},\,\dots,\,p_{T}\leq x_{T}\bigr{)}=x_{t}\,\mathbb{E}\Bigl{[}\mathbf{1}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbf{1}\{p_{T}\leq x_{T}\}\Bigr{]}.$
Next, we note the following two key facts:
(a) Uniformity of single p-value $p_{t}$:
Fix ant $t\in[T]$. Let $\Omega_{t}$ be the sets of all permutations of ${\mathcal{C}}_{0}\cup[T]$ that fixes indices outside of ${\mathcal{C}}_{0t}\cup\{t\}$. Note that given $\Phi_{t}$, the only randomness for $p_{t}$ is the order of $\{({\bf X}_{i},Y_{i}):i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ and $\xi_{t}$.
For any $\sigma\in\Omega_{t}$ and given $\{{\bf X}_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}=\{{\bf x}_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ as the realizations, define
$p_{t}(({\mathcal{D}}_{T})_{\sigma})$
$=$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{\sigma(i)};({\mathcal{D}}_{T})_{\sigma})<V({\bf x}_{\sigma(t)};({\mathcal{D}}_{T})_{\sigma})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{\sigma(i)};({\mathcal{D}}_{T})_{\sigma})=V({\bf x}_{\sigma(t)};({\mathcal{D}}_{T})_{\sigma})\}}{1+|{\mathcal{C}}_{0t}|}$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{\sigma(i)};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{\sigma(i)};{\mathcal{D}}_{t})=V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}$
$\stackrel{{\scriptstyle(ii)}}{{=}}$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})=V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}.$
Equality (i) holds since $V(\cdot;\sigma(D_{t}))=V(\cdot;D_{t})$ from the symmetry of score function $V$. Equality (ii) is true by the property that $\sigma$ only permutes indices in ${\mathcal{C}}_{0t}\cup\{t\}$ such that $\{V({\bf X}_{i};D_{t}):i\in{\mathcal{C}}_{0t}\}=\{V({\bf X}_{\sigma(i)};D_{t}):i\in{\mathcal{C}}_{0t}\}$.
Denote $Q_{\alpha}(S_{i}:i\in\mathcal{I})$ as the $\alpha$-th quantile of the set $\{S_{i}\}_{i\in{\mathcal{I}}}$. Let $\{X_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}=\{x_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ as a set of realizations and
$q=Q_{\alpha}(V(x_{i};{\mathcal{D}}_{t}):i\in{\mathcal{C}}_{0t}\cup\{t\}).$
For any $\sigma\in\Omega_{t}$ such that $V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})>q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma})\geq\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}{\geq}\frac{N_{=}+N_{-}}{1+|{\mathcal{C}}_{0t}|}\geq\alpha;$
Here
$N_{-}=\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbf{1}\{(V(x_{i};{\mathcal{D}}_{t})<q\},\quad N_{=}=\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbf{1}\{(V(x_{i};{\mathcal{D}}_{t})=q\}.$
The last inequality holds by the property of quantile function such that $N_{=}+N_{-}\geq 1+\alpha(|{\mathcal{C}}_{0t}+1|)$.
For any $\sigma\in\Omega_{t}$ such that $V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})<q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})\leq V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}\leq\frac{N_{-}}{1+|{\mathcal{C}}_{0t}|}\leq\alpha,$
which is from the property that $N_{-}\leq\alpha(|{\mathcal{C}}_{0t}+1|)$.
And for any $\sigma\in\Omega_{t}$ such that $V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma})=$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})=V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}$
$\leq$ $\frac{N_{-}+\xi_{t}N_{=}}{1+|{\mathcal{C}}_{0t}|}.$
Hence,
$\Pr(p_{t}\leq\alpha\mid\Phi_{t})$ $=\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}\sum_{\sigma\in\Omega_{t}}\Pr(p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq\alpha)$
$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}\sum_{\sigma\in\Omega_{t}}\left(\mathbf{1}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})<q\}+\mathbf{1}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q\}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}\right)$
$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{|{\mathcal{C}}_{0t}|+1}\sum_{i\in{\mathcal{C}}_{0t}\cup\{j\}}\left(\mathbf{1}\{V({\bf x}_{i};{\mathcal{D}}_{t})<q\}+\mathbf{1}\{V({\bf x}_{i};{\mathcal{D}}_{t})=q\}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}\right)$
$\stackrel{{\scriptstyle(iii)}}{{=}}\frac{1}{|{\mathcal{C}}_{0t}|+1}\left(N_{-}+N_{=}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}\right)=\alpha,$
where equality (i) holds since $\xi_{t}$ is uniformly distributed and independent of everything else, making
${\mathbb{E}}[\mathbf{1}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q,p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq\alpha\}\mid({\mathcal{D}}_{T})_{\sigma}]=\mathbf{1}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q\}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}.$
Equality (ii) is from the fact that $\sum_{\sigma\in\Omega_{t},\sigma(t)=i}\mathbf{1}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})<q\}=|{\mathcal{C}}_{0t}|\mathbf{1}\{V({\bf x}_{i};{\mathcal{D}}_{t})<q\}$ and $\sum_{\sigma\in\Omega_{t},\sigma(t)=i}\mathbf{1}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q\}=|{\mathcal{C}}_{0t}|\mathbf{1}\{V({\bf x}_{i};{\mathcal{D}}_{t})=q\}$. And equality (iii) is direct by the definition of $N_{-}$ and $N_{=}$. Marginalizing over the $\Phi_{t}$ implies $\Pr(p_{t}\leq\alpha)=\alpha$ for all $t\in[T]$.
(b) Independence of future $p$-values from $\Phi_{t}$:
Define the data set ${\mathcal{D}}_{T}=(({\bf X}_{-n},Y_{-n}),\dots,({\bf X}_{T},Y_{T}))$. By exchangeability of the data, it holds that ${\mathcal{D}}_{T}\overset{d}{=}({\mathcal{D}}_{T})_{\sigma}$, where $({\mathcal{D}}_{T})_{\sigma}$ is obtained from ${\mathcal{D}}_{T}$ by permuting the data points according to $\sigma$. A key observation is that for any $\sigma\in\Omega_{t}$, defining $\sigma$ as above we have
$p_{t^{\prime}}({\mathcal{D}}_{T})=p_{t^{\prime}}(({\mathcal{D}}_{T})_{\sigma})$
for all $t^{\prime}\in\{t+1,\dots,T\}$. Intuitively, this indicates that permuting the data according to $\sigma$ does not change $p$-values after time $t$.
This is because $p_{t^{\prime}}(({\mathcal{D}}_{T})_{\sigma})$ is
$\frac{\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbf{1}\{V({\bf X}_{\sigma(i)};\sigma({\mathcal{D}}_{t^{\prime}}))<V({\bf X}_{\sigma(t^{\prime})};\sigma({\mathcal{D}}_{t^{\prime}}))+\xi_{t^{\prime}}(1+\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbf{1}\{V({\bf X}_{\sigma(i)};\sigma({\mathcal{D}}_{t^{\prime}}))=V({\bf X}_{\sigma(t^{\prime})};\sigma({\mathcal{D}}_{t^{\prime}}))\})}{1+|{\mathcal{C}}_{0t^{\prime}}|}$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\frac{\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbf{1}\{V({\bf X}_{\sigma(i)};{\mathcal{D}}_{t^{\prime}})<V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\}+\xi_{t^{\prime}}(1+\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbf{1}\{V({\bf X}_{\sigma(i)};{\mathcal{D}}_{t^{\prime}})=V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\})}{1+|{\mathcal{C}}_{0t^{\prime}}|}\;$
$\stackrel{{\scriptstyle(ii)}}{{=}}$ $\frac{\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbf{1}\{V({\bf X}_{i};{\mathcal{D}}_{t^{\prime}})<V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\}+\xi_{t^{\prime}}(1+\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbf{1}\{V({\bf X}_{i};{\mathcal{D}}_{t^{\prime}})=V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\})}{1+|{\mathcal{C}}_{0t^{\prime}}|}\;$
$=$ $p_{t^{\prime}}({\mathcal{D}}_{T}),$
where equality (i) holds since $\sigma(t^{\prime})=t^{\prime}$ by definition and $V(\cdot;\sigma(D_{t^{\prime}}))=V(\cdot;D_{t^{\prime}})$ from the symmetry of score function $V$. And equality (ii) is true as $\{V({\bf X}_{i};D_{t^{\prime}}):i\in{\mathcal{C}}_{0t^{\prime}}\}=\{V({\bf X}_{\sigma(i)};D_{t^{\prime}}):i\in{\mathcal{C}}_{0t^{\prime}}\}$.
Therefore,
$\Pr(p_{t}\leq x_{t},\cdots,p_{T}\leq x_{t})$
$=$ $\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}{\mathbb{E}}\left[\sum_{\sigma\in\Omega_{t}}\mathbf{1}\{p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq x_{t},\cdots,p_{T}(({\mathcal{D}}_{T})_{\sigma})\leq x_{t}\}\right]$
$=$ $\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}{\mathbb{E}}\left[\sum_{\sigma\in\Omega_{t}}\mathbf{1}\{p_{t}({\mathcal{D}}_{T})\leq x_{t},\cdots,p_{T}({\mathcal{D}}_{T})\leq x_{t}\}\right]$
$=$ $\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}{\mathbb{E}}\left[\left(\sum_{\sigma\in\Omega_{t}}\mathbf{1}\{p_{t}({\mathcal{D}}_{T})\leq x_{t}\}\right)\mathbf{1}\{p_{t+1}({\mathcal{D}}_{T})\cdots,p_{T}({\mathcal{D}}_{T})\leq x_{t}\}\right]$
$=$ ${\mathbb{E}}\left[\Pr\left(p_{t}\leq x_{t}\mid\Phi_{t}\right)\mathbf{1}\{p_{t+1}({\mathcal{D}}_{T})\cdots,p_{T}({\mathcal{D}}_{T})\leq x_{t}\}\right]$
$=$ $x_{t}\Pr(p_{t+1}({\mathcal{D}}_{T})\cdots,p_{T}({\mathcal{D}}_{T})\leq x_{t}).$
We can iterate the above conditioning argument. That is, applying the same argument for $p_{t+1}$ conditional on $\Phi_{t+1}$ (which contains information up to time $t+1$), we obtain
$\Pr\bigl{(}p_{t}\leq x_{t},\,p_{t+1}\leq x_{t+1},\dots,p_{T}\leq x_{T}\bigr{)}=x_{t}\,x_{t+1}\,\cdots\,x_{T}.$
Since the joint cumulative distribution function factors as the product of the marginals, it follows that the sequence $\{p_{t},\dots,p_{T}\}$ is mutually independent, with each $p_{t}$ marginally distributed as $\mathrm{Uniform}(0,1)$.
Thus, we have demonstrated that under the exchangeability assumption and the online updating scheme, the online conformal $p$-values defined in (20) are mutually independent. ∎
A.4 Proof of Theorem 3: finite sample mFDR control
Proof.
mFDR control for Algorithm 2. We first prove the results for LFS.
Denote $\Psi_{t}=\Big{(}(p_{k}:k\in{\mathcal{C}}_{0t}),(\theta_{k}:k<t),(V_{k}:k\in{\mathcal{C}}_{1t})\Big{)}$. We need to verify two facts, for any time $t\in[T]$,
(i): $\alpha_{t}^{\text{LFS}}$ is fixed given $\Psi_{t}$. Note that $\alpha_{t}^{\text{LFS}}$ is fixed given all past null decisions $(\delta_{i}:1\leq i<t)$ and a null decision $\delta_{i}$ is decided by $((p_{k},\alpha_{k}):\theta_{k}=0,1\leq k\leq i)$. By iterated discussion, $\alpha_{t}$ is decided by null p-values $(p_{k}:k\in{\mathcal{C}}_{0t})$, thereby determined fully by $\Psi_{t}$.
(ii): $p_{t}$ is super-uniform given $\Psi_{t}$ and $\theta_{t}=0$. This is direct as long as $p_{t}$ is independent of past null p-values given non-null data, which is verified by Proposition 3.1.
Then we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t}(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$
${=}$ $\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j}){\mathbb{E}}\left[\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid\Psi_{j},\theta_{j}=0\right]\right]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})+\alpha_{t}\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right]=\alpha\cdot{\mathbb{E}}[R(t)],$
where the equality (i) follows from the above two facts, thereby $\Pr(p_{j}\leq\alpha_{j}\mid\Psi_{j},\theta_{j}=0)\leq\alpha_{j}$. And the inequality (ii) follows by the fact that $1-\theta_{t}\leq 1$. For inequality (iii) follows from the construction of $\alpha_{t}^{\text{LFS}}$, and the inequality (iv) holds since $1-\theta_{j}\leq 1$ for all $j\leq t$. Therefore, we conclude that $\operatorname{mFDR}(t)\leq\alpha$ for LFS.
Similar results for SFS can also be proved as follows:
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t}(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$
${=}$ $\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j}){\mathbb{E}}\left[\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid\Psi_{j},\theta_{j}=0\right]\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}(1-\theta_{t})\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}.\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}(1-\theta_{j})\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right]=\alpha\cdot{\mathbb{E}}[R(t)].$
Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for the proposed OCTF procedures in Algorithm 2 with $\alpha_{t}=\alpha_{t}^{\text{LFS}}$ or $\alpha_{t}=\alpha_{t}^{\text{SFS}}$. ∎
A.5 Proof of Corollary 3.1
Proof.
It suffices to verify that the optimized score function $V(\cdot;\hat{k}_{t})$ is symmetric with respect to ${\mathcal{C}}_{0t}\cup\{t\}$, so that Proposition A.1 and Theorem 3 can be directly applied to establish the validity of mFDR control at the target level for Algorithm Opt-OCTF.
For any permutation $\sigma\in\Omega_{t}$ that only permutes the indices in ${\mathcal{C}}_{0t}\cup\{t\}$, we have
$\hat{k}_{t}^{\sigma}=\underset{k\in[K]}{\arg\min}\;{\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,({\mathcal{D}}_{T})_{\sigma})=\hat{k}_{t}.$
To see why, the auxiliary p-value for $j\in{\mathcal{C}}_{1t}$ after permutation $\sigma$ is
$(\tilde{p}^{k}_{j})_{\sigma}$ $=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{\sigma(s)};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|}$
$=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|}=\tilde{p}^{k}_{j}.$
This means $\{\tilde{p}^{k}_{j}\}_{j\in{\mathcal{C}}_{1t}}$ is permutation invariant to $\sigma$. Applying this, we have
${\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,({\mathcal{D}}_{t})_{\sigma})=\frac{\sum_{j=1}^{t-1}\rho^{t-1-j}\,(\tilde{p}^{k}_{j})_{\sigma}\theta_{j}\,}{\sum_{j=1}^{t-1}\rho^{t-1-j}\cdot\theta_{j}}=\frac{\sum_{j=1}^{t-1}\rho^{t-1-j}\,\tilde{p}^{k}_{j}\theta_{j}\,}{\sum_{j=1}^{t-1}\rho^{t-1-j}\cdot\theta_{j}}={\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t}),$
which keeps invariant to the permutation $\sigma$. Combing together, we have $\hat{k}_{t}$ is symmetric to ${\mathcal{C}}_{0t}\cup\{t\}$. And it indicates that $V(\cdot;\hat{k}_{t})$ is symmetric to ${\mathcal{C}}_{0t}\cup\{t\}$. ∎
A.6 Proof of Theorem 4
Proof.
We first prove the results for $e$-LF. Since $\delta_{j}=\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}$, by definition, we have
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}(1-\theta_{j})}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\mathbb{I}\{1/e_{j}\leq\alpha_{j}\}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})e_{j}\alpha_{j}}{R({j-1})+1}\right]$
$=$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j}){\mathbb{E}}[e_{j}\mid{\mathcal{F}}_{j-1}]\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}+\frac{\alpha_{t}}{1+R({t-1})}\right]$
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\text{e-LF}}(t)\right]\leq\alpha,$ (21)
where the first inequality holds since $R({j-1})+1\leq(R(t)\vee 1)$ for every $j\in\{j\leq t:\delta_{j}=1\}$ by definition, the second inequality holds since $\mathbb{I}\{e_{j}\alpha_{j}\geq 1\}\leq e_{j}\alpha_{j}$, the third inequality uses the law of iterated expectations by conditioning on $\mathcal{F}_{j-1}$ and then applies the property of $e$-values, and the fourth inequality holds since $1-\theta_{t}\leq 1$, and the last inequality follows from the construction of $e$-LF, which completes the proof of FDR control for $e$-LF.
We now proceed to establish the FDR control guarantee for the $e$-SF procedure. Specifically, we show that:
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}(1-\theta_{j})}{1\vee\sum_{j\leq t}\delta_{j}}\right]$ (22)
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\mathbb{I}\{1/e_{j}\leq\alpha_{j}\}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})e_{j}\alpha_{j}}{R({j-1})+1}\right]$
$=$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j}){\mathbb{E}}[e_{j}\mid{\mathcal{F}}_{j-1}]\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\cdot\frac{{\mathbb{E}}\left[\mathbb{I}\{e_{j}<1/\lambda_{j}\}\mid{\mathcal{F}}_{j-1}\right]}{1-\lambda_{j}}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\cdot\frac{\mathbb{I}\{e_{j}<1/\lambda_{j}\}}{1-\lambda_{j}}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t-1}\frac{\alpha_{j}(1-\theta_{j})}{R({j-1})+1}\cdot\frac{\mathbb{I}\{e_{j}\leq 1/\lambda_{j}\}}{1-\lambda_{j}}+\frac{\alpha_{t}}{R({t-1})+1}\cdot\frac{\mathbb{I}\{e_{t}\leq 1/\lambda_{t}\}}{1-\lambda_{t}}\right]$
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\text{e-SF}}(t)\right]\leq\alpha,$
where the first inequality holds since $R({j-1})+1\leq(R(t)\vee 1)$ for every $j\in\{j\leq t:\delta_{j}=1\}$ by definition, the second inequality holds since $\mathbb{I}\{e_{j}\alpha_{j}\geq 1\}\leq e_{j}\alpha_{j}$, the third inequality uses the law of iterated expectations by conditioning on $\mathcal{F}_{j-1}$ and then applies the property of $e$-values, and the fourth inequality holds since ${\mathbb{E}}\left[\mathbb{I}\{e_{j}<1/\lambda_{j}\}\mid{\mathcal{F}}_{j-1}\right]\geq 1-\lambda_{j}$ by the property of $e$-values, the fifth inequality follows from $1-\theta_{t}\leq 1$, and the last inequality follows from the construction of $e$-SF, which completes the proof of FDR control for $e$-SF. ∎
A.7 Discussion about improving LOND/e-LOND with feedback
Recall the setup of the LOND or $e$-LOND algorithm. Given a non-negative sequence $\{\gamma_{j}\}_{j=1}^{\infty}$ such that $\sum_{j=1}^{\infty}\gamma_{j}=1$, the test levels are set as
$\alpha_{t}^{\text{LOND}}=\alpha\gamma_{t}\left(\sum_{j=1}^{t-1}\delta_{j}\vee 1\right)=\alpha\gamma_{t}\left({R}(t-1)\vee 1\right).$
Zrnic et al. [2021] proved that LOND controls the FDR under PRDS. Note that for any $t\in\mathbb{N}$ :
$\operatorname{FDR}(t)=\mathbb{E}\left[\frac{\sum_{i\leq t,i\in\mathcal{H}_{0}}\mathbf{1}\left\{P_{i}\leq\alpha_{i}\right\}}{{R}(t)\vee 1}\right]\leq\sum_{i\leq t,i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{1\left\{P_{i}\leq\alpha_{i}\right\}}{{R}(i-1)\vee 1}\right]=\sum_{i\leq t,i\in\mathcal{H}_{0}}\gamma_{i}\alpha\mathbb{E}\left[\frac{\mathbf{1}\left\{P_{i}\leq\alpha_{i}\right\}}{\alpha_{i}}\right]$
where the first equality follows by definition of FDR, the sole inequality follows because the number of rejections can only increase with time, and the second equality follows by definition of the LOND rule for $\alpha_{i}$. Lemma 1 from Ramdas et al. [2019] now asserts that the term in the expectation is bounded by one under PRDS. Hence, by also noting that $\sum_{i<t}\gamma_{i}\leq 1$ completes the proof.
Improve LOND or $e$-LOND with feedback? (negative result)
Here we prove that the feedback cannot be used to improve LOND and $e$-LOND.
Proof. Denote $a_{j}=\sum_{i=1}^{j-1}\delta_{i}\theta_{i}$ and $b_{j}=\sum_{i=1}^{j-1}\delta_{i}(1-\theta_{i})$, so that ${R}({j-1})=a_{j}+b_{j}$. For LOND or e-LOND, the significance level is given by
$\alpha_{j}=\alpha\gamma_{j}(a_{j}+b_{j}+1).$
If $\theta_{1},\dots,\theta_{j-1}$ are known, we consider the modified significance level
$\tilde{\alpha}_{j}=\alpha\gamma_{j}(2w_{j}a_{j}+2(1-w_{j})b_{j}+1),$
where $w_{j}\in[0,1]$ is a weight that differentiates between true and false discoveries. Setting $w_{j}=1$ recovers $\tilde{\alpha}_{j}=\alpha_{j}$. To control $\operatorname{FDR}(t)\leq\alpha$, we require
$\sum_{j\in\mathcal{H}_{0}\cap[t]}\frac{\tilde{\alpha}_{j}}{{R}({j-1})+1}\leq\alpha.$
This implies
$2w_{j}a_{j}+2(1-w_{j})b_{j}+1\leq a_{j}+b_{j}+1.$
To improve power, we require
$\tilde{\alpha}_{j}>\alpha_{j},$
which translates to
$2w_{j}a_{j}+2(1-w_{j})b_{j}+1>a_{j}+b_{j}+1.$
[…]
Appendix B Preliminary Terms for Self-Containment
Here, we list the preliminary terms we use in the paper for the sake of clarity and self-containment.
  • $\operatorname{FDR}$ [Benjamini and Hochberg, 1995], false discovery rate, a widely-adopted error rate notion in the field of multiple testing, is defined as the expected proportion of incorrectly rejected null hypotheses as follows:
    $\operatorname{FDP}(t)=\frac{V_{t}}{1\vee R_{t}}:=\frac{\sum_{j=1}^{t}\delta_{j}(1-\theta_{j})}{1\vee\sum_{j=1}^{t}\delta_{j}}\quad\text{and}\quad\operatorname{FDR}(t):={\mathbb{E}}[\operatorname{FDP}(t)],$
    where $R(t)$ represents the number of rejected null hypotheses until time $t$ and $V(t)$ is the number of false discoveries.
  • $\operatorname{mFDR}$, modified false discovery rate, or marginal false discovery rate, is defined as:
    $\operatorname{mFDR}(t):=\frac{{\mathbb{E}}[V(t)]}{{\mathbb{E}}[1\vee R(t)]}=\frac{{\mathbb{E}}[\sum_{j=1}^{t}\delta_{j}(1-\theta_{j})]}{{\mathbb{E}}[1\vee\sum_{j=1}^{t}\delta_{j}]}.$
  • Testing levels of existing online multiple testing methods. For the online methods, denote the decision rule as $\delta_{t}=\mathbb{\{}p_{t}\leq\alpha_{t}\}$, where $p_{t}$ is the corresponding conformal $p$-value at time $t$ for our problem. The test levels $\{\alpha_{t}\}$ for LOND [Javanmard and Montanari, 2015], LORD++ [Ramdas et al., 2017], SAFFRON [Ramdas et al., 2018] and ADDIS [Tian and Ramdas, 2019], $\text{LORD}_{\text{dep}}$ [Zrnic et al., 2021], $\text{SAFFRON}_{\text{dep}}$ [Zrnic et al., 2021] are listed as follows:
    • 1.
      LOND: $\alpha_{t}=\gamma_{t}(R(t-1)+1)$, where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to $\alpha$ and $R(n)=\sum_{t=1}^{n}R_{t}$ denotes the number of discoveries in the first $n$ hypotheses tested.
    • 2.
      LORD++:
      $\alpha_{t}^{\text{LORD++}}=\gamma_{t}W_{0}+(\alpha-W_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}},$
      where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to one; $\tau_{j}$ is the time of the $j$-th rejection.
    • 3.
      SAFFRON: At each time $t$, define $C_{j+}=C_{j+}(t)=\sum_{i=\tau_{j}+1}^{t-1}$, where $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$. For $t=1$, $\alpha_{1}=\min\{\gamma_{1}W_{0},\lambda\}$; For $t=2,3,\dots$, $\alpha_{t}:=\min\{\lambda,\tilde{\alpha}_{t}\}$, where
      $\tilde{\alpha}_{t}=W_{0}\gamma_{t-C_{0+}}+((1-\lambda)\alpha-W_{0})\gamma_{t-\tau_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-C_{j+}}.$
    • 4.
      ADDIS: The testing levels for ADDIS are given by $\alpha_{t}=\min\{\lambda,\hat{\alpha}_{t}\}$, where
      $\hat{\alpha}_{t}=(\eta-\lambda)[\omega_{0}\gamma_{S^{t}-C_{0+}}+(\alpha-\omega_{0})\gamma_{S^{t}-\tau_{1}^{*}-C_{1+}}+\alpha\sum_{j\geq 2}\gamma_{S^{t}-\tau_{j}^{*}-C_{j+}}]$
      and $S^{t}=\sum_{i<t}\mathbb{I}\{p_{i}\leq\eta\}$, $\tau_{j}^{*}=\sum_{i\leq\tau_{j}}\mathbb{I}\{p_{i}\leq\eta\}$.
    • 5.
      $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$: Define $r_{k}$ under local dependence as:
      $r_{k}=\min\{i\in[t]:\sum_{j=1}^{i-L_{i+1}}\delta_{j}\geq k\}.$
      The corresponding test levels for $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$ are as follows:
      $\alpha_{t}^{\text{LORD}_{\text{dep}}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-r_{1}}\mathbb{I}\{r_{1}<t\}+\alpha\sum_{j=2}^{\infty}\gamma_{t-r_{j}}.$
      $\alpha_{t}^{\text{SAFFRON}_{\text{dep}}}:=\min\left\{\lambda,(1-\lambda)\left(w_{0}\gamma_{t-C_{0+}}+(\alpha-w_{0})\gamma_{t-r_{1}-C_{1+}}+\alpha(\sum_{j\geq 2}\gamma_{t-r_{j}-C_{j+}})\right)\right\},$
      where $C_{j+}=\sum_{i=r_{j}+1}^{t-L_{t}+1}C_{i}$.
  • Conformal $p$-values. The notion of conformal $p$-values was originally introduced by Vovk et al. [2005] for constructing prediction intervals. A conformal $p$-value quantifies how well a new observation conforms to a reference set, based on a chosen nonconformity score function. More recently, several works have applied conformal $p$-values to sample selection from a multiple testing perspective [Jin and Candès, 2023a, Bates et al., 2023, Wang et al., 2024]. The conformal $p$-values are defined as
    $p_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0}}{\mathbb{I}}\{V_{i}\leq V_{t}\}}{1+|{\mathcal{C}}_{0}|},$ (23)
    where ${\mathcal{C}}_{0}$ is a hold-out calibration dataset and $V(\cdot)$ is a nonconformity score function. However, in these approaches, the conformal $p$-values are constructed using a fixed offline calibration dataset, which limits their flexibility in online or adaptive settings.
Appendix C Additional Details of Our Algorithms
In this Section, we provide additional details of our algorithms, including the concrete testing levels of proposed algorithms in Appendix C.1, Optimized OCTF by adaptively tuned online learning in Appendix C.2, and the additional explanation about utilizing feedback in SAFFRON-type algorithms is in Appendix C.3.
C.1 Testing levels of the proposed approaches
  • 1.
    Testing levels for $\text{LF}_{\text{dep}}$ and $\text{SF}_{\text{dep}}$:
    Define $r_{k}$ under local dependence as:
    $r_{k}=\min\{i\in[t]:\sum_{j=1}^{i-L_{i+1}}\delta_{j}\geq k\}.$
    The corresponding test levels for $\text{LF}_{\text{dep}}$ and $\text{SF}_{\text{dep}}$ are constructed as follows:
    $\alpha_{t}^{\text{LF}_{\text{dep}}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-r_{1}}\mathbb{I}\{r_{1}<t\}+\alpha\sum_{j=2}^{\infty}\gamma_{t-r_{j}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j},$
    $\alpha_{t}^{\text{SF}_{\text{dep}}}:=\min\left\{\lambda,(1-\lambda)\left(w_{0}\gamma_{t-C_{0+}}+(\alpha-w_{0})\gamma_{t-r_{1}-C_{1+}}+\alpha(\sum_{j\geq 2}\gamma_{t-r_{j}-C_{j+}})\right)+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j}\right\},$
    where $C_{j+}=\sum_{i=r_{j}+1}^{t-L_{t}}C_{i}$.
  • 2.
    Testing levels for LFS and SFS: Recall that the proposed LF update rule is given by:
    $\alpha_{t}^{\text{LF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j},$ (24)
    where $\tau_{j}$ denotes the time of the $j$-th rejection. We revise this rule to:
    ${\alpha}_{t}^{\text{LFS}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}}\mathbb{I}\{{\tilde{\tau}}_{1}<t\}+\alpha\sum_{j:\tilde{\tau}_{j}<t,\tilde{\tau}_{j}\neq\tilde{\tau}_{1}}\gamma_{t-\tilde{\tau}_{j}}+\sum_{j:j<t}\gamma_{t-j}{\alpha}_{j}\theta_{j},$ (25)
    where $\tilde{\tau}_{j}$ denotes the time of the $j$-th rejection under the null, defined as
    $\tilde{\tau}_{j}=\inf\left\{t\in\mathbb{N}:\sum_{i\leq t}\delta_{i}(1-\theta_{i})\geq j\right\}.$
    $\alpha_{t}^{\text{SFS}}:=\min\{\lambda,\tilde{\alpha}_{t}^{\text{SFS}}\},$ (26)
    where $\tilde{\alpha}_{t}^{\text{SFS}}=s_{0}\gamma_{t-C_{0+}}+((1-\lambda)\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tilde{\tau}_{j}-C_{j+}}+\sum_{j:j<t}\gamma_{t-j}\alpha_{j}\theta_{j}.$
C.2 Optimized OCTF by adaptively tuned online learning
Our framework also allows online learning for the predictive models and the model selection is specific to tuning the parameters of model with online updating. Suppose we have $K$ candidate hyper-parameters. We will implement $K$ corresponding models in parallel throughout our procedure. At each time $t$, the models will be updated by training on the current data $\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{t-1}\cup\{({\bf X}_{t},\tilde{Y}_{t})\}$. Here as $Y_{t}$ is unobserved currently, we will regard $\tilde{Y}_{t}$ as the null value. To select a best model for our testing procedure, we can use a similar criterion as ${\mathcal{M}}^{\mathrm{EWMA}}_{t}(k,{\mathcal{D}}_{t})$. And the conformal p-value is constructed by the model $\tilde{\mu}_{\hat{k}_{t}}^{(t)}$. The detailed pseudo-codes of the adaptively tuned online learning procedure is shown in Algorithm 4.
Algorithm 4 Optimized OCTF by adaptively tuned online learning
0:  Initial data ${\mathcal{D}}_{1}=\{({\bf X}_{i},Y_{i})\}_{i=-N}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, $K$ candidate hyper-parameters $\{\eta_{k}\}_{k=1}^{K}$, loss function $L(\beta,Y;\eta_{k})$ for the predictive model using hyperparameter $\eta_{k}$, evaluating criterion $\mathcal{M}$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}$, stopping time $T$.
1:  for $t=1,\dots,T$ do
2:   Observe test data ${\bf X}_{t}$
3:   Online update one-step-ahead models for $k=1,\cdots,K$
[…]
and the corresponding predictive model is $\tilde{\mu}_{k}^{(t)}$. The score $V(\cdot;\tilde{\mu}_{k}^{(t)})$ is determined by the predictive model $\tilde{\mu}_{k}^{(t)}$.
4:   Decide the predictive model for $t$-th test sample by
$\widehat{k}_{t}=\underset{k\in[K]}{\arg\max}\;{\mathcal{M}}^{\text{EWMA}}(k,{\mathcal{D}}_{t})$
as the score function $V(\cdot;\tilde{\mu}_{k}^{(t)})$ is decided by $\tilde{\mu}_{k}^{(t)}$.
5:   Construct optimized conformal p-value by
$p^{\text{opt}}_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\{V({\bf X}_{i};\tilde{\mu}_{\hat{k}_{t}}^{(t)})\leq V({\bf X}_{t};\tilde{\mu}_{\hat{k}_{t}}^{(t)})\}}{1+|{\mathcal{C}}_{0t}|}$
6:   Update $\alpha_{t}={\alpha}_{t}^{\text{LFS}}$ in Equation (25) or $\alpha_{t}={\alpha}_{t}^{\text{SFS}}$ in Equation (26)
7:   Make a decision $\delta_{t}=\mathbb{I}\{p_{t}^{\text{opt}}\leq\alpha_{t}\}$
8:   Obtain the revealed feedback $Y_{t}$ (and thus obtain $\theta_{t}$)
9:   Update the calibration dataset ${\mathcal{C}}_{0t}$.
10:  end for
10:  Rejection set $\mathcal{R}_{\text{opt}}=\{i:\delta_{i}=1,~\delta_{i}\in{\bm{\delta}}^{T}\}$.
C.3 Explanation about utilizing feedback in SAFFRON
From another perspective, one may propose using feedback to estimate the null proportion and operate in a manner similar to SAFFRON, but in a form different from ours in Section 2.2, where the construction of $\alpha_{t}$ is
${\alpha}^{\rm SF-variant}_{t}=s_{0}\gamma_{t-D_{0+}}+(\alpha-s_{0})\gamma_{t-\tau_{1}-D_{1+}}+\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-D_{j+}},$ (27)
here $D_{j+}=\sum_{i=\tau_{j}+1}^{t-1}{\mathbb{I}}\{\theta_{i}=1\}$. This approach constructs $\alpha_{t}$ using $\{\theta_{i}\}_{i=1}^{t-1}$ only, and can provide the same finite-sample online FDR guarantee.
Here, we explain why this form is not adopted in the proposed GAIF. In practice, we typically have $D_{j}^{+}<C_{j}^{+}$, since the set filtered by $p$-values is larger than the true alternative set. Consequently, due to the exponential decay design of $\gamma_{t}$, we have
$\gamma_{t-\tau_{j}-D_{j+}}\ll\gamma_{t-\tau_{j}-C_{j+}}.$
Although the original SAFFRON includes an adjustment factor $1-\lambda$ which reduces the $\alpha_{t}$, this term becomes negligible compared to the exponential decay of $\gamma_{t-\tau_{1}-C_{1+}}$. As a result, the original SAFFRON achieves higher power than this naive feedback-based variant, especially when $\lambda$ is large.
Our empirical results in Figure 7 further confirm this phenomenon. In the Gaussian setting described in Section 4.1 with $\mu=2$ and $\pi=0.3$, we observe that when $\lambda>0.3$, the original SAFFRON outperforms the variant in terms of power. Moreover, its highest power occurs at $\lambda=0.8$, which is relatively large. This highlights the trade-off between $\gamma_{t-\tau_{j}-C_{j+}}$ and $1-\lambda$, where where the rapidly decaying $\gamma_{t-\tau_{j}-C_{j+}}$ dominates, rendering $1-\lambda$ less influential.
In conclusion, directly using feedback to estimate the null proportion is not advisable, as it leads to a poor construction of $\alpha_{t}$. By contrast, our SF strategy leverages the $p$-values to achieve adaptive $\alpha$-wealth allocation and demonstrates superior performance.
Figure 7: The FDR and Power for SAFFRON at stopping time $600$ under different $\lambda$ value for target FDR level $\alpha=0.1$. The red lines denote the results for the variant of feedback method.
Appendix D Extensions of GAIF based on $e$-values
Although feedback cannot be directly used to improve $e$-LOND [Xu and Ramdas, 2024], it can enhance $e$-LORD and $e$-SAFFRON [Zhang et al., 2025] through a feedback-driven approach analogous to the extension from GAI to GAIF. Denote $\delta_{j}=\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}$. Denote
$\operatorname{FDP}^{*}(t)=\sum_{j\leq t,j\in\mathcal{H}_{0}}\frac{\alpha_{j}}{1+R({j-1})}\leq\sum_{j\leq t}\frac{\alpha_{j}}{1+R({j-1})}=\widehat{\operatorname{FDP}}_{\text{e-LORD}}.$
Similar to the GAIF framework, we propose the estimators $\widehat{\operatorname{FDP}}_{\text{e-LF}}$ and $\widehat{\operatorname{FDP}}_{\text{e-SF}}$ as follows:
$\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\leq\widehat{\operatorname{FDP}}_{\text{e-LF}}=\sum_{j\leq t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}+\frac{\alpha_{t}}{1+R({t-1})}.$
$\widehat{\operatorname{FDP}}(t)_{\text{e-SF}}=\sum_{j\leq t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}\frac{\mathbb{I}\{e_{j}\leq 1/\lambda_{j}\}}{1-\lambda_{j}}+\frac{\alpha_{t}}{1+R({t-1})}\frac{\mathbb{I}\{e_{t}\leq 1/\lambda_{t}\}}{1-\lambda_{t}}.$
Then we require $\widehat{\operatorname{FDP}}_{\text{e-LF}}\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\text{e-SF}}\leq\alpha$ when constructing the testing levels. The corresponding testing levels are as follows:
$\alpha_{t}^{\text{e-LF}}=\omega_{t}\left(\alpha-\sum_{j=1}^{t-1}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}\right)(R({t-1})+1),$ (28)
$\alpha_{t}^{\text{e-SF}}=\omega_{t}\left(\alpha(1-\lambda)-\sum_{j=1}^{t-1}\frac{\alpha_{j}(1-\theta_{j})\mathbb{I}\{e_{j}<1/\lambda\}}{1+R({j-1})}\right)(R({t-1})+1),$ (29)
where $\omega_{t}\in(0,1)$ is updated by
$\omega_{t+1}=\omega_{t}+\omega_{1}\varphi^{t-R(t)}(1-\delta_{t})-\omega_{1}\psi^{R(t)}\delta_{t}$ (30)
with a user-defined initial allocation coefficient $\omega_{1}\in(0,1)$, and user-defined parameters $\varphi>0,\psi>0$.
The Generalized Alpha-Investing with Feedback procedure based on $e$-values (e-GAIF) is summarized in Algorithm 5. The $e$-LF and $e$-SF can control online FDR validly if the null $e$-values satisfy conditionally valid.
Algorithm 5 $e$-GAIF ($e$-LF and $e$-SF)
0:  Target $\operatorname{FDR}$ level $\alpha$, parameters $\lambda,\varphi$ and $\psi\in(0,1)$, initial allocation coefficient $\omega_{1}\in(0,1)$.
1:  for $t=1,2,\dots$ do
2:   Observe $e$-value $e_{t}$
3:   Update $\alpha_{t}=\alpha_{t}^{\text{e-LF}}$ in Eq.(28) (or $\alpha_{t}=\alpha_{t}^{\text{e-SF}}$ in Eq.(29))
4:   if $e_{t}\geq 1/\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$
5:   Update $R(t)=R(t-1)+\delta_{t}$ and $\omega_{t+1}$ by (30)
6:   Obtain the revealed feedback $\theta_{t}$
7:  end for
7:  Rejection set $\{t:\delta_{t}=1\}$.
Theorem 4 (Online FDR control for e-GAIF).
If the null $e$-values are conditional valid, i.e.,
${\mathbb{E}}\left[e_{t}\mid\mathcal{F}_{t-1}\right]\leq 1\;\text{for all}\;t\in\mathcal{H}_{0},$ (31)
where $\mathcal{F}_{t-1}=\sigma(\delta_{1},\dots,\delta_{t-1};\theta_{1},\dots,\theta_{t-1})$ is the sigma field generated from past rejections and feedback, then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{e-LF}}(t)\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\operatorname{e-SF}}(t)\leq\alpha$ , then we have
$\operatorname{FDR}(t)\leq\alpha\quad\text{for all}\quad t\in\mathbb{N}.$
Appendix E Applications on Real-time LLM Alignment
In this section, we introduce the potential application of our proposed OCTF procedure on the task of real-time LLM alignment. For example, in medical report generation tasks, we may need to sequentially select radiology images $t\in\{1,\dots,T\}$ for which the generated reports align with expert standards. Similarly, in question-answering tasks, our goal is to identify the generated answer that best matches the true reference answer in an online fashion. Specifically, let $f:\mathcal{X}\to\mathcal{Y}$ be a pre-trained foundation model that maps a prompt to an output. A holdout set $\mathcal{D}=({\bf X}_{i},E_{i})_{i=-n}^{0}$ is available, where ${\bf X}_{i}\in\mathcal{X}$ represents an input prompt, and $E_{i}\in\mathcal{E}$ serves as a reference for assessing alignment. The alignment score function $\mathcal{A}:\mathcal{Y}\times\mathcal{E}\to\mathbb{R}$ maps the generated output $f({\bf X})$ and reference $E$ to an alignment score $A=\mathcal{A}(f({\bf X}),E)$. For example, $A$ may represent the similarity score between the machine-generated report $f({\bf X})$ and a human expert report $E$. The test data $\{{\bf X}_{t}\}_{t=1}^{T}$ arrive sequentially, forming the following online multiple hypothesis testing problem at time $t$:
$\mathbb{H}_{0t}:A_{t}\leq c\quad\text{versus}\quad\mathbb{H}_{1t}:A_{t}>c,$
where $c\in\mathbb{R}$ is a pre-specified threshold. After making a decision $\delta_{t}$ at time $t$, the corresponding human expert report is revealed either immediately or with a delay of $d$ time steps.
Our goal is to control the online $\operatorname{FDR}$:
$\operatorname{FDR}(t)={\mathbb{E}}\left[\frac{\sum_{j\leq t}\mathbb{I}\{A_{j}\leq c,\delta_{j}=1\}}{1\vee\sum_{j\leq t}\delta_{j}}\right]\leq\alpha.$ (32)
Following the conformal alignment framework Gui et al. [2024], we randomly split $\mathcal{D}$ into two subsets: a training set $\mathcal{D}_{\mathcal{T}}$ and a calibration set $\mathcal{D}_{\mathcal{C}}$. Using $\mathcal{D}_{\mathcal{T}}$, we train an alignment predictor $g({\bf X})$ to estimate the alignment score based on features of LLM outputs ${\bf X}_{t}$ and compute the predicted alignment scores $\widehat{A}_{t}=g({\bf X}_{t})$ for each $t\in[T]$. Then applying OCTF with $\mathcal{D}_{\mathcal{C}}$ to select new images whose generated reports are aligned with expert standards ensures finite-sample online FDR control according to Theorem 3. The conformal $p$-value for $t\in[T]$ is:
$p_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}{\mathbb{I}}\{\widehat{A}_{i}\geq\widehat{A}_{t}\}}{1+|{\mathcal{C}}_{0t}|},$ (33)
The real-time LLM alignment procedure is summarized in Algorithm 6, which, similar to Theorem 3, guarantees finite-sample mFDR control under the same assumptions.
Algorithm 6 Real-time LLM conformal Alignment with feedback
0: Pre-trained foundation model $f$; alignment score function $\mathcal{A}$; reference dataset $\mathcal{D}=({\bf X}_{i},E_{i})_{i=-n}^{0}$; ; algorithm for fitting alignment predictor $\mathcal{G}$; alignment level $c$; target FDR level $\alpha$.
1: Compute the alignment score $A_{i}=\mathcal{A}(f({\bf X}_{i}),E_{i})$, $\forall i\in\mathcal{D}$.
2: Randomly split $\mathcal{D}$ into two disjoint sets: the training set $\mathcal{D}_{{\mathcal{T}}}$ and the calibration set $\mathcal{D}_{{\mathcal{C}}}$.
3: Fit the alignment score predictor with $\mathcal{D}_{{\mathcal{T}}}$: $g\leftarrow\mathcal{G}(\mathcal{D}_{{\mathcal{T}}})$.
4: Initialize ${\mathcal{C}}_{0t}=\{i\in{\mathcal{C}}:Y_{i}\leq c\}$
5:for $t\in[T]$ do
6:  Observe test data ${\bf X}_{n+t}$
7:  Compute the predicted alignment score: $\hat{A}_{i}\leftarrow g({\bf X}_{i})$, $\forall i\in{\mathcal{C}}$ and $\hat{A}_{t}\leftarrow g({\bf X}_{t})$.
8:  Compute the conformal p-value $p_{t}$ according to Equation (33).
9:  Update $\alpha_{t}=\alpha_{t}^{\text{LFS}}$ in Equation (25) (or $\alpha_{t}=\alpha_{t}^{\text{SFS}}$ in (26))
10:  Obtain the revealed feedback $\theta_{t}$.
11:  Update the calibration dataset ${\mathcal{C}}_{0t}$.
12:end for
12: The selection set $\mathcal{R}=\{t:\delta_{t}=1,t\in[T]\}$.
Appendix F Additional Experiments Results
In this section, we provide additional experimental results to further demonstrate the superior performance of our proposed algorithms, with a focus on online conformal testing.
Specifically, we report additional results for the classification example (Scenarios IV) as well as for a regression example (Scenario VI). The corresponding data generation process for Scenario VI is detailed below:
  • Scenario VI (Regression example): $Y=-0.5X_{1}^{2}+\exp{X_{2}}+(X_{3}+X_{4})^{2}+\varepsilon,$ with ${\bf X}\sim\mathcal{N}_{4}({\bf 0},\mathbf{I}_{4})$ and $\varepsilon\sim\mathcal{N}(0,2)$. The target region is $\mathcal{A}=[c,\infty)$, where $c$ is the $1-\pi_{1}$ quantile of $Y$.
In terms of the non-conformity score function, denote $W_{t}=\widehat{\mu}_{t}({\bf X}_{t})$, in classification settings, we set $V(W_{t})=1-W_{t}$. In regression settings, if $\mathcal{A}=[b,+\infty)$, we can use $V(W_{j})=b-W_{j}$. If $\mathcal{A}=(-\infty,a]\cup[b,+\infty)$, then we can choose $V(W_{t})=\max\{W_{t}-a,b-W_{t}\}$.
Results.
The results for Scenario VI using a fixed training algorithm (random forest) are shown in Figure 8. The performance trends are similar to those observed in Scenario IV.
Figure 8: Results for Scenario VI: values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportions $\pi_{1}$. The black dashed line denotes the FDR level $\alpha=0.2$.
The results for Scenarios IV and VI under different training algorithms—RF, SVM, and NN—with varying initial calibration sizes are presented in Figure 9-10. Thanks to the online updating of the calibration dataset, even a small initial calibration size does not significantly impact performance. While the choice of predictive model $\widehat{\mu}$ does affect power, all methods maintain valid FDR control. Notably, our SF and LF methods consistently outperform the baselines across all models, benefiting from the distribution-free and model-agnostic nature of online conformal $p$-values. The variation in performance across different algorithms further underscores the importance of careful model selection in practice.
Figure 9: Results for Scenario IV: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ vs. initial calibration size $n$ ($\pi_{1}=0.5$, $\alpha=0.2$).
Figure 10: Results for Scenario VI: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ vs. initial calibration size $n$ ($\pi_{1}=0.5$, $\alpha=0.2$).
The results with model selection for Scenarios IV and VI are shown below Figure 11-Figure 12. In both settings, the performance gap between the Opt methods and their randomly selected counterparts is also pronounced, with Opt-SF once again achieving the highest power.
Figure 11: Results for Scenario IV: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.1$.
Figure 12: Results for Scenario VI: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.1$.
To illustrate the similarity between mFDR and FDR, we present results under various settings below in Figure 13-Figure 14. We estimate mFDR by computing the ratio of the average number of false discoveries and the average total number of discoveries. Empirical mFDR closely tracks empirical FDR, and both are well controlled by our proposed methods.
Figure 13: Results for Scenario III (local dependence): Line charts of mFDR and FDR at stopping time with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$. The black dashed lines denote the target FDR level $\alpha=0.1$.
Figure 14: Results for Scenario IV and Scenario VI : Line charts of mFDR and FDR at stopping time with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$ after $500$ replications; The black dashed lines denote the target FDR level $\alpha=0.2$.