### v2

# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

Abstract
This work studies online multiple testing with feedback, where decisions are made sequentially, and the true state of the hypothesis is revealed after decisions are made, either instantly or with a delay, and under either full or bandit feedback. We propose Generalized alpha-investing with feedback (GAIF) along with its adaptive variants, a feedback-enhanced framework that dynamically adjusts thresholds using revealed outcomes, ensuring finite-sample false discovery rate (FDR)/marginal FDR (mFDR) control. We further extend GAIF to online conformal testing by constructing valid conformal $p$-values and developing feedback-enhanced testing rules with finite-sample mFDR control. We also propose a feedback-driven score selection criterion to adaptively choose the candidate score that is most effective for the testing procedure, together with a theoretical analysis of its optimality. Numerical simulations and real-data applications demonstrate the effectiveness of our methods.
**footnotetext: The first two authors equally contributed to this work.footnotetext: Correspondence to: Haojie Ren <haojieren@sjtu.edu.cn>
Keywords: Conformal prediction, distribution shifts, generalized alpha-investing procedure, model selection, online conformal $p$-value, online FDR control
1 Introduction
Real-time decision making plays a critical role in a growing number of modern applications, such as online recruitment for job hiring [Faliagka et al., 2014], real-time alignment of large language models [Huang et al., 2025], and time-series anomaly detection [Rebjock et al., 2021]. These tasks can be naturally formulated as online multiple testing problems [Foster and Stine, 2008]. Consider a potentially infinite stream of null hypotheses $\{\mathbb{H}_{01},\mathbb{H}_{02},\dots,\mathbb{H}_{0t},\dots\}$ that are tested sequentially based on incoming test statistics such as $p$-values $\{p_{1},p_{2},\dots p_{t},\dots\}$. At each time $t$, a real-time decision must be made based on the current test statistic. Let $\theta_{t}$ denote the true state of the hypothesis at time $t$, where $\theta_{t}=0$ if $\mathbb{H}_{0t}$ is true and $\theta_{t}=1$ otherwise. The testing problem at time $t$ is
$\mathbb{H}_{0t}:\theta_{t}=0\quad\text{vs.}\quad\mathbb{H}_{1t}:\theta_{t}=1.$
Let $\delta_{t}\in\{0,1\}$ be the decision at time $t$, with $\delta_{t}=1$ indicating rejection of $\mathbb{H}_{0t}$. To ensure the reliability of the testing procedure, it is essential to control a well-defined error rate. Define the false discovery proportion (FDP) and false discovery rate (FDR) [Benjamini and Hochberg, 1995] at time $t$ by
$\operatorname{FDP}(t)=\frac{V(t)}{1\vee R(t)}\;:=\;\frac{\sum_{j=1}^{t}\delta_{j}\bigl(1-\theta_{j}\bigr)}{1\vee\sum_{j=1}^{t}\delta_{j}},\quad\operatorname{FDR}(t)=\mathbb{E}\left\{\operatorname{FDP}(t)\right\},$
where $V(t)$ and $R(t)$ represent the numbers of false rejections (discoveries) and rejections at time $t$, respectively. Given a target level $\alpha\in(0,1)$, classical work on online multiple testing [Ramdas et al., 2017, 2018, Tian and Ramdas, 2019] aims to guarantee $\sup_{t}\operatorname{FDR}(t)\leq\alpha$ or its weaker variant [Foster and Stine, 2008, Zrnic et al., 2021], the marginal FDR (mFDR), i.e., $\operatorname{mFDR}(t)={\mathbb{E}\{V(t)\}}/{\mathbb{E}\{1\vee R(t)\}}$.
Unlike the classical setting, we consider another realistic scenario in which feedback on past hypotheses is revealed after decisions are made. Specifically, after issuing $\delta_{t}$, the corresponding state $\theta_{t}$ may be observed either immediately or after a delay, and in some settings it may be observed only partially. For instance, in the bandit setting [Neu et al., 2010], feedback $\theta_{t}$ is observed only when the decision $\delta_{t}=1$ is made. Our work encompasses all these feedback-available settings, including full or partial feedback, as well as immediate or delayed feedback.
Such feedback-available settings naturally arise in many practical applications. To demonstrate its relevance, we present three motivating examples: online conformal selection, real-time alignment of large language models, and time-series anomaly detection.
  • Online conformal selection. Conformal selection aims to identify valuable individuals whose unknown label $Y$ satisfies a pre-specified requirement by leveraging machine learning predictions [Wu et al., 2024, Jin and Candès, 2025]. The online conformal selection setting naturally aligns with the feedback-available online multiple testing framework. For example, in diabetes risk prediction, a patient’s true condition may later be confirmed by an expert, offering feedback that could improve future decisions. Some works [Huo et al., 2024, Xu and Ramdas, 2024] considered online conformal testing problems but did not exploit feedback information.
  • Real-time LLM alignment. Large language models (LLMs) are increasingly used in high-stakes domains such as healthcare, finance, and law, where outputs must be reliable and factual. However, LLMs can hallucinate—producing plausible but incorrect content [Huang et al., 2025]. A remedy is to filter or certify LLM outputs [Gui et al., 2024, Bai and Jin, 2024]. Some alignment approaches based on conformal testing are proposed for this use but are generally designed for offline environments, whereas many applications demand immediate and trustworthy screening. In such applications, follow-up feedback is usually available, which, if incorporated, could continuously improve alignment.
  • Time series anomaly detection. Detecting anomalies in time series is crucial for industrial monitoring, fraud detection, and healthcare analytics. To ensure reliability, prior work [Rebjock et al., 2021, Gang et al., 2023, Krönert et al., 2023] addresses online FDR control but typically ignores real-time feedback. In practice, such feedback is often available: once an anomaly is flagged, a subsequent system failure, user verification of a fraudulent transaction, or expert annotation may confirm or refute the alarm.
These observations underscore that feedback is a foundational element of adaptive decision-making, playing a pivotal role across a wide range of online applications. Despite its importance, current online multiple testing procedures seldom incorporate feedback in a systematic way. This naturally raises a fundamental and compelling question: Can available feedback be effectively incorporated into online multiple testing procedures in a way that allows us to enhance statistical power while still ensuring valid error rate control?
To this end, we develop a feedback-enhanced framework that systematically integrates revealed feedback into the generalized alpha-investing (GAI) framework [Aharoni and Rosset, 2014, Javanmard and Montanari, 2018], achieving significant performance gains without compromising statistical validity.
1.1 Our contributions
To the best of our knowledge, this is the first work to incorporate feedback information directly into the construction of testing thresholds for online FDR procedures and to employ this idea to online conformal testing. Our main contributions are summarized as follows:
  • GAIF and its variants. We introduce Generalized Alpha-Investing with Feedback (GAIF), a feedback-enhanced extension of the GAI framework, together with its adaptive variants. The key idea is to use revealed feedback to reduce the slack in FDP estimation, and, in the adaptive version, to further improve $\alpha$-wealth allocation. The framework applies to a broad range of feedback settings, including full or partial feedback, as well as immediate or delayed feedback, and it also extends to locally dependent scenarios. We establish online FDR control under independence and online mFDR control under local dependence for GAIF and its variants.
  • Online conformal testing with feedback. We extend the GAIF to online conformal testing. We first provide an explicit construction of valid, independent $p$-values (for null hypotheses) by dynamically updating the calibration dataset. We then combine them with suitably modified GAIF rules to obtain online conformal testing with feedback (OCTF), which enjoys finite-sample online mFDR control. This extension bridges the gap between traditional online multiple testing and conformal inference, yielding distribution-free, model-agnostic tools for real-time decision-making.
  • Feedback-driven score selection. We further introduce a strategy to adaptively select among candidate predictive models or conformity scores. This is particularly important under non-stationary non-null distributions, where the optimal score may vary over time. By locally averaging recent non-null $p$-values to track the best-performing score, our selection criterion can accommodate such shifts while maintaining validity.
We provide rigorous proofs for all proposed methods. Extensive simulations and real‐data experiments demonstrate that our procedures substantially outperform existing approaches while effectively controlling the online FDR when feedback information is provided.
1.2 Related works
Our work is situated at the intersection of online multiple testing and conformal inference. We review key developments in each area and highlight gaps that motivate our contribution.
Online multiple testing under independence.
Early works on online multiple testing began with the alpha-investing strategy of Foster and Stine [2008], later generalized by Aharoni and Rosset [2014] and Javanmard and Montanari [2018] into the GAI framework, which led to the LORD algorithm. Building on this line of work, Ramdas et al. [2017] introduced LORD++, an improved version of GAI tailored for online FDR control. Subsequent refinements include SAFFRON [Ramdas et al., 2018], which adapts to the proportion of non-nulls, and ADDIS [Tian and Ramdas, 2019], which adjusts for conservative null $p$-values. These methods, including LORD++, SAFFRON, and ADDIS, achieve online FDR control when null $p$-values are independent of all other $p$-values. Separately, Gang et al. [2023] developed structure-adaptive rules based on local FDR, which improve power but only ensure asymptotic FDR control under correct model specification. For a comprehensive review, see Robertson et al. [2023]. All the above methods determine thresholds solely from past rejections, without considering real-time feedback.
Online multiple testing under dependence.
In practice, hypotheses often exhibit dependence, and applying methods designed for independence can lead to inflated error rates. To address arbitrary dependence, Xu and Ramdas [2024] proposed e-LOND, an FDR-controlling procedure based on e-values. Zhang et al. [2025] extended this approach to e-GAI, achieving improved power by dynamically allocating the testing levels. Alternatively, research has focused on local dependence structures. Zrnic et al. [2021] introduced LORDdep and SAFFRONdep, establishing mFDR control under local dependence; Rebjock et al. [2021] later adapted these methods to time-series anomaly detection. Recently, Fisher [2024] showed that LORD++ with suitable local modifications can maintain FDR control under certain dependence, while Fischer et al. [2024] proposed an online procedure for p-values satisfying positive regression dependence on a subset (PRDS). Despite these advances, existing dependence-aware methods do not incorporate any real-time feedback.
Conformal inference and conformal multiple testing.
Conformal inference [Vovk et al., 2005] offers a model-agnostic way to quantify prediction uncertainty. In the multiple testing setting, early works constructed conformal $p$-values and applied the Benjamini–Hochberg (BH) procedure [Benjamini and Hochberg, 1995] to achieve finite-sample FDR control [Bates et al., 2023, Jin and Candès, 2025]. Subsequent extensions addressed covariate shift [Jin and Candès, 2023], constrained selection [Wu et al., 2024, Nair et al., 2025], and conditional testing [Wu et al., 2025], as well as model selection [Bai and Jin, 2024, Gui et al., 2025]. However, these contributions remain confined to offline settings. Few efforts to extend conformal multiple testing to the online domain [Huo et al., 2024, Xu and Ramdas, 2024, Liu et al., 2026] did not take the feedback information into account. Although related research on the construction of online conformal prediction sets [Gibbs and Candès, 2021, 2024, Xi et al., 2025] has considered feedback, multiple testing problems with FDR control remain largely unexplored. One very recent work, by Humbert et al. [2025], establishes asymptotic online FDP control through an online learning strategy, while we achieve feedback-enhanced testing based on the GAI framework.
1.3 Organization of the paper
The remainder of this paper is organized as follows. Section 2 introduces the GAIF and establishes FDR control under independence. Section 3 extends the GAIF to achieve adaptive $\alpha$-wealth allocation and to address local dependence. In Section 4, we construct explicit online conformal $p$-values and apply modified GAIF rules to online conformal testing, providing finite-sample theoretical guarantees. We also extend the proposed framework to achieve score selection and address distribution shifts. Simulation and real-data experiment results are presented in Sections 5 and 6, respectively. Finally, we conclude the paper in Section 7.
2 Generalized Alpha-Investing with Feedback (GAIF)
In this section, we begin by revisiting the generalized alpha-investing (GAI) framework in Section 2.1. In Section 2.2, we introduce our feedback-enhanced framework, termed generalized alpha-investing with feedback (GAIF), which incorporates feedback information to improve FDP estimation. In Section 2.3, we establish finite-sample online FDR control for GAIF procedures.
2.1 Recap: Generalized alpha-investing
Early approaches to error rate control in an online setting were based on so‐called “alpha‐investing” strategies [Foster and Stine, 2008], which were subsequently generalized into the GAI framework [Aharoni and Rosset, 2014, Javanmard and Montanari, 2018]. The key idea is to compare each incoming pp‐value $p_{t}$ against a dynamically chosen threshold $\alpha_{t}$, to accumulate additional $\alpha$‐wealth” upon each rejection, and to make testing decisions according to $\delta_{t}=\mathbb{I}\{p_{t}\leq\alpha_{t}\}$.
Based on GAI, Ramdas et al. [2017] proposed controlling online FDR by ensuring that an estimator of the FDP remains below a pre‐specified level $\alpha$. A specific example is the LORD++ algorithm, where the estimated $\operatorname{FDP}$ at time $t$ is given by
$\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\overset{(i)}{\geq}\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\approx\operatorname{FDP}(t):=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}},$ (1)
where $\mathcal{H}_{0}$ is the index set of null hypotheses. To ensure that $\operatorname{FDR}(t)\leq\alpha$, it suffices to enforce $\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{R(t)\vee 1}\leq\alpha.$ The corresponding testing level of LORD++ is
$\alpha_{t}^{\text{LORD++}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}},$
where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to one, $\tau_{j}$ is the time of the $j$-th rejection, and $s_{0}>0$ is the pre-specified initial wealth.
Building on this, some adaptive versions of LORD++ were proposed subsequently, such as SAFFRON [Ramdas et al., 2018] and ADDIS [Tian and Ramdas, 2019]. Specifically, SAFFRON introduces a user-specified parameter $\lambda\in[0,1]$ to estimate the proportion of true nulls in the online testing stream, and obtains a less conservative estimated FDP as
$\widehat{\operatorname{FDP}}_{\text{SAFFRON}}(t)=\frac{\sum_{j\leq t}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}}{1\vee\sum_{j\leq t}\delta_{j}}.$ (2)
These GAI‐based procedures guarantee online mFDR control under the conditional super‐uniformity assumption for null pp‐values:
$\Pr\bigl(p_{t}\leq\alpha_{t}\,\bigm|\,\mathcal{G}_{t-1}\bigr)\;\leq\;\alpha_{t}\quad\text{for all }t\in\mathcal{H}_{0},$ (3)
where $\mathcal{G}_{t}=\sigma(\delta_{1:t})$ and each threshold $\alpha_{t}=f_{t}(\delta_{1:t-1})$ is measurable with respect to past decisions. Here $\delta_{1:t}:=(\delta_{1},\dots,\delta_{t})$. Furthermore, if all null pp‐values are independent of other pp‐values, these procedures also ensure strict FDR control.
Note that the inequality (i) in (1) is tight if and only if all past hypotheses are null. This conservative step is one of the major sources of the gap between the realized FDR of LORD++ and the target level $\alpha$. When feedback becomes available over time, it allows us to sharpen the FDP estimator and accordingly reduce the resulting slack.
2.2 Boosting GAI via feedback: GAIF
We now propose generalized alpha-investing with feedback (GAIF), which leverages revealed feedback information to improve FDP estimation. The key idea is that once some past hypotheses are known to be non-nulls, their contribution to the FDP upper bound can be reduced accordingly, leaving more $\alpha$-wealth available for future tests.
Formally, let
[…]
denote the index set of hypotheses whose labels are available by time $t$, and let $\bar{\mathcal{I}}_{t}=[t]\setminus\mathcal{I}_{t}$, where $t]=\{1,2,\dots,t\}$. Realistic scenarios can lead to diverse feedback structures. For example, feedback may be available only when a rejection is made or may arrive after a delay $d\geq 0$. Our framework encompasses a variety of practically relevant cases:
  • Full and instant feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-1\}$;
  • Bandit and instant feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-1,\delta_{j}=1\}$;
  • Full and delayed feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-d-1\}$, where $d$ is the delay;
  • Bandit and delayed feedback: $\mathcal{I}_{t}=\{j\in[t]:j\leq t-d-1,\delta_{j}=1\}$.
At time $t$, the available feedback reveals the true values of $\{\theta_{j}\}_{j\in\mathcal{I}_{t}}$. We leverage this information to refine the FDP estimator as follows.
Definition 2.1 (GAIF).
Define the feedback-enhanced FDP estimator as
$\widehat{\mathrm{FDP}}_{\mathrm{GAIF}}(t):=\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\,\alpha_{j}\;+\;\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}}{1\,\vee\,\sum_{j=1}^{t}\delta_{j}}.$
A procedure is said to be a GAI with Feedback (GAIF) procedure if it assigns test levels $\{\alpha_{t}\}$ such that, for every $t$,
$\widehat{\mathrm{FDP}}_{\mathrm{GAIF}}(t)\leq\alpha.$
The GAIF framework aligns with intuition. For $j\in\mathcal{I}_{t}$, feedback reveals whether the corresponding hypothesis is null or non-null, so their contribution to the FDP bound can be accounted for precisely through $(1-\theta_{j})\alpha_{j}$. For indices not yet revealed, i.e., $j\in\bar{\mathcal{I}}_{t}$, we keep the contribution $\alpha_{j}$. Thus, GAIF improves the FDP estimator relative to the no-feedback case.
Importantly, this refinement leads directly to a less conservative constraint on the test levels $\alpha_{t}$. Recall that LORD++ enforces $\alpha_{t}\leq\alpha(1\vee\sum_{j\leq t}\delta_{j})-\sum_{j\leq t-1}\alpha_{j}$. In contrast, a GAIF procedure selects $\alpha_{t}$ to satisfy
$\alpha_{t}\leq\alpha(1\vee\sum_{j=1}^{t}\delta_{j})-\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})-\sum_{j\in\bar{\mathcal{I}}_{t}\setminus t}\alpha_{j}.$ (4)
The presence of revealed feedback enlarges the admissible upper bound for $\alpha_{t}$. Indeed, since $1-\theta_{j}\leq 1$ for every $j\in\mathcal{I}_{t}$, (4) implies
$\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})+\sum_{j\in\bar{\mathcal{I}}_{t}\setminus t}\alpha_{j}\leq\sum_{j\in\mathcal{I}_{t}}\alpha_{j}+\sum_{j\in\bar{\mathcal{I}}_{t}\setminus t}\alpha_{j}=\sum_{j=1}^{t-1}\alpha_{j}.$
Therefore, the upper bound of $\alpha_{t}$ under GAIF is always at least as large as that under LORD++. In this sense, an appropriate procedure with feedback is expected to improve power by sharpening the FDP estimator and releasing more $\alpha$-wealth for future allocation.
We next present a concrete instance of GAIF as follows:
$\alpha_{t}^{\text{GAIF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j}.$ (5)
Note that if all feedback is ignored so that $\mathcal{I}_{t}=\emptyset$, the $\alpha_{t}^{\rm GAIF}$ reduces exactly to that of LORD++. This further implies that the test levels $\alpha_{t}^{\rm GAIF}$ with the additional feedback term are never smaller than those of LORD++ whenever feedback is available.
Algorithm 1 summarizes the implementation of this concrete GAIF procedure under a general feedback regime. The only difference across settings lies in how the feedback-available set $\mathcal{I}_{t}$ is updated. We denote the resulting procedures under full and instant feedback, bandit and instant feedback, full and delayed feedback, and bandit and delayed feedback by LF, LF-BI, LF-FD, and LF-BD, respectively.
Algorithm 1 GAIF procedures
0:  Target FDR level $\alpha$; pre-specified parameters for constructing test levels.
1:  for $t=1,2,\dots$ do
2:   Observe the $p$-value $p_{t}$.
3:   Update $\mathcal{I}_{t}$ according to the specific feedback regime.
4:   Construct $\alpha_{t}=\alpha_{t}^{\mathrm{GAIF}}$ using Definition 2.1 with the current $\mathcal{I}_{t}$.
5:   if $p_{t}\leq\alpha_{t}$ then set $\delta_{t}=1$; else set $\delta_{t}=0$.
6:   Reveal feedback by obtaining $\theta_{j}$ for all $j\in\mathcal{I}_{t+1}\setminus\mathcal{I}_{t}$.
7:  end for
7:  Rejection set $\mathcal{R}=\{t:\delta_{t}=1\}$.
2.3 Theoretical guarantee under independence
The following result states that GAIF procedures guarantee online mFDR control under conditional super-uniformity of null $p$-values
$\Pr\bigl(p_{t}\leq\alpha_{t}\,\bigm|\,\mathcal{F}_{t-1}\bigr)\;\leq\;\alpha_{t}\quad\text{for all }t\in\mathcal{H}_{0},$ (6)
where the enlarged filtration $\mathcal{F}_{t-1}:=\sigma(\delta_{1:t-1};\{\theta_{j}\}_{j\in{\mathcal{I}}_{t}})$. Under stronger independence and monotonicity assumptions, GAIF further achieves online FDR control. Here, the test level sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ is said to be a monotonic function of the past if, for all $t\in\mathbb{N}$, $\alpha_{t}$ is coordinate-wise non-decreasing in the past decisions $\{\delta_{j}:j<t\}$ and revealed feedback $\{\theta_{j}:j\in\mathcal{I}_{t}\}$ for the GAIF procedure.
Theorem 1 (Online mFDR and FDR control for GAIF).
Let $\{\alpha_{t}\}_{t\in\mathbb{N}}$ be a sequence of test levels satisfying $\widehat{\mathrm{FDP}}_{\rm GAIF}(t)\leq\alpha$. Then:
  • (a)
    If the null $p$-values are conditionally super-uniformly distributed as in (6), then the procedure guarantees
    $\operatorname{mFDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
  • (b)
    If the null $p$-values are mutually independent and also independent of the non-null $p$-values, and if the test levels $\{\alpha_{t}\}$ are a monotone sequence of functions of the past for all $t$, then the procedure satisfies
    $\operatorname{FDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
The test levels of GAIF in (5) inherit the monotonicity structure from LORD++, and therefore satisfy the monotonicity condition needed for Theorem 1 (b). The proof of Theorem 1 is provided in Appendix F.2.
3 Extensions of GAIF
In this section, we discuss two important extensions of GAIF. We first develop the Adaptive GAIF framework in Section 3.1, which further enhances GAIF by enabling adaptive $\alpha$-wealth allocation. We then extend our methods to the settings with local dependence in Section 3.2.
3.1 Adaptive GAIF: improve GAIF via adaptive $\alpha$-wealth allocation
Section 2.2 shows that feedback improves power by sharpening the FDP estimate. In online testing, power also depends on how the available $\alpha$-wealth is allocated over time, since the test levels $\alpha_{t}$ must be chosen sequentially. This motivates an improvement direction beyond feedback-enhanced FDP estimation, namely adaptive $\alpha$-wealth allocation.
To motivate this idea, we revisit the difference between LORD++ and SAFFRON. The SAFFRON uses a weighting function $\kappa(p):=\frac{\mathbb{I}\{p>\lambda\}}{1-\lambda}$ in its FDP estimator (2). This construction is typically interpreted as a way to estimate the null proportion in the online FDR literature, and has been studied in related contexts, notably in the offline multiple testing setting [Storey et al., 2004, Döhler and Meah, 2023].
Here we emphasize another complementary perspective: $\kappa(\cdot)$ also serves as a guiding mechanism for adaptive $\alpha$-wealth allocation. Indeed, under LORD++, the test levels satisfy $\alpha_{t}\leq\alpha\sum_{j=1}^{t}\delta_{j}-\sum_{j=1}^{t-1}\alpha_{j}$. By contrast, SAFFRON enforces
$\alpha_{t}\leq\alpha\,\frac{1-\lambda}{\mathbb{I}\{p_{t}>\lambda\}}\sum_{j=1}^{t}\delta_{j}-\frac{1}{\mathbb{I}\{p_{t}>\lambda\}}\sum_{j=1}^{t-1}\alpha_{j}\mathbb{I}\{p_{j}>\lambda\}.$ (7)
When $p_{t}\leq\lambda$, the denominator in (7) vanishes, and the upper bound is interpreted as $\infty$, allowing more aggressive investment at promising tests. When $p_{t}>\lambda$ and $p_{j}>\lambda$ for all $j\leq t-1$, the upper bound becomes $\alpha(1-\lambda)\sum_{j=1}^{t}\delta_{j}-\sum_{j=1}^{t-1}\alpha_{j}$, which is smaller than the LORD++ bound. Thus, the weighting function $\kappa(\cdot)$ effectively spends less $\alpha$-wealth on unpromising tests and reserves more for promising ones.
This analysis suggests that the GAIF framework in Section 2.2 can be further improved by adopting an adaptive $\alpha$-wealth allocation strategy. Next, we demonstrate that explicitly incorporating observed $p$-value patterns enables more efficient $\alpha$-wealth allocation. We formalize this idea below:
Definition 3.1 (Adaptive GAIF).
Define the adaptive feedback-enhanced FDP estimator as
$\widehat{\mathrm{FDP}}_{\mathrm{Ada\text{-}GAIF}}(t):=\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\,\alpha_{j}\kappa(p_{j})\;+\;\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}\kappa(p_{j})}{1\vee\sum_{j=1}^{t}\delta_{j}},$
where $\kappa(p)=\frac{\mathbb{I}\{p>\lambda\}}{1-\lambda}$ and $\lambda\in[0,1]$ is a user-chosen parameter for identifying large pp‑values. A procedure is called an Adaptive GAIF procedure if it assigns test levels $\{\alpha_{t}\}$ such that, for every $t$,
$\widehat{\mathrm{FDP}}_{\mathrm{Ada\text{-}GAIF}}(t)\leq\alpha.$ (8)
The key difference between GAIF and Adaptive GAIF is the introduction of the $\kappa(p)$ in the FDP estimate. This weight returns the unused $\alpha$-budget from tests with large $p$-values (i.e., $p_{j}>\lambda$) for future allocation, while feedback continues to reduce the slack in FDP estimation through the revealed $\theta_{j}$’s. If feedback is ignored so that $\mathcal{I}_{t}=\emptyset$, the FDP estimator of Adaptive GAIF reduces to that of SAFFRON.
Remark 3.1.
In principle, other constructions of $\kappa(\cdot)$ may further optimize $\alpha$-wealth allocation while ensuring online FDR control. This online optimization remains an open problem and a direction for future work [Gang et al., 2023]. Instead, we focus on the SAFFORN-type weight $\frac{\mathbb{I}\{p>\lambda\}}{1-\lambda}$, which offers a theoretically grounded and practically appealing compromise with robust empirical performance.
We next present concrete instances of Adaptive GAIF. Its thresholds $\{\alpha_{t}\}$ are in the same spirit as SAFFRON, but with the tightened, feedback-aware constraint in (8). For $t=1$, $\alpha_{1}^{\text{Ada-GAIF}}=\min\{\gamma_{1}s_{0},\lambda\}$. For $t>1$,
$\alpha_{t}^{\text{Ada-GAIF}}=\min\bigl\{\lambda,s_{0}\gamma_{t-C_{0+}}+\{(1-\lambda)\alpha-s_{0}\}\gamma_{t-\tau_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-C_{j+}}+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j}\bigr\},$ (9)
where $C_{j+}=C_{j+}(t)=\sum_{i=\tau_{j}+1}^{t-1}C_{i}$ and $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$.
The Adaptive GAIF procedures benefit simultaneously from feedback-enhanced FDP estimation and adaptive $\alpha$-wealth allocation, and therefore are expected to achieve higher power. We denote the concrete Adaptive GAIF procedures under full and instant feedback, bandit and instant feedback, full and delayed feedback, and bandit and delayed feedback by SF, SF-BI, SF-FD, and SF-BD, respectively.
Consider LF of GAIF and SF of Adaptive GAIF under the full and instant feedback scenario. Figure 1 depicts the resulting testing thresholds $\{\alpha_{t}\}$ over time $t$ for various procedures applied to Gaussian observations. It is clear that our methods yield larger thresholds after improving the gap via feedback, with $\alpha_{t}^{\text{Ada-GAIF}}>\alpha_{t}^{\text{SAFFRON}}$ and $\alpha_{t}^{\text{GAIF}}>\alpha_{t}^{\text{LORD++}}$ on average. This illustrates that the GAIF and Adaptive GAIF procedures leverage the available $\alpha$-wealth more effectively through the consideration of feedback.
Figure 1: Results for testing Gaussian observations. Line charts depict the average of $\alpha_{t}$ over time $t$ under various testing procedures, based on 500 replications with a non-null proportion of $\pi_{1}=0.5$. Details of the data generation process are provided in Section 5.1.
We now establish the theoretical results for Adaptive GAIF. Similar to (6), the conditional super-uniformity condition on the null $p$-values is
$\Pr\bigl(p_{t}\leq\alpha_{t}\,\bigm|\,\mathcal{J}_{t-1}\bigr)\;\leq\;\alpha_{t}\quad\text{for all }t\in\mathcal{H}_{0},$ (10)
where $\mathcal{J}_{t-1}:=\sigma(\delta_{1:t-1};C_{1:t-1};\{\theta_{j}\}_{j\in{\mathcal{I}}_{t}})$ is the enlarged filtration, where $C_{j}=\mathbb{I}\{p_{j}\leq\lambda\}$. And we extend the monotonicity definition in Theorem 1 by additionally requiring $\alpha_{t}$ to be coordinate-wise non-decreasing in the indicators $\{C_{j}:j<t\}$.
Theorem 2 (Online mFDR and FDR control for Adaptive GAIF).
Let $\{\alpha_{t}\}_{t\in\mathbb{N}}$ be a sequence of test levels satisfying $\widehat{\mathrm{FDP}}_{\rm Ada\text{-}GAIF}(t)\leq\alpha$. Then:
  • (a)
    If the null $p$-values are conditionally super-uniform (10), the procedure guarantees
    $\operatorname{mFDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
  • (b)
    If the null $p$-values are mutually independent and independent of the non-null $p$-values, and if the test levels $\{\alpha_{t}\}$ form a monotone sequence of functions of the past for all $t$, then the procedure satisfies
    $\operatorname{FDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
The test levels of Adaptive GAIF in (9) inherit the monotonicity structure from SAFFRON, and therefore satisfy the monotonicity condition needed for Theorem 2 (b). The proof of Theorem 2 is provided in Appendix F.3.
3.2 GAIF under local dependence
This subsection focuses on extending the GAIF and Adaptive GAIF framework to settings with local dependence (as defined in Definition 3.2). This extension is useful in applications where nearby tests may be dependent, while long-range dependence is absent or weak. For a strategy that addresses more general dependence structures via e-values, the reader may refer to Appendix B.2.
Definition 3.2 (Local dependence; Zrnic et al. [2021]).
We say that $p$-values $p_{1},p_{2},\dots,p_{t},\dots$ are locally dependent if
$\text{for all}\;t>0,\;\text{there exists}\;L_{t}\in\mathbb{N}\;\text{such that}\;p_{t}\perp p_{t-L_{t}-1},p_{t-L_{t}-2},\dots,p_{1},$ (11)
where $\{L_{t}\}_{t\in\mathbb{N}}$ is a fixed sequence of parameters which we refer to as lags.
Zrnic et al. [2021] pioneered a strategy to handle this local dependence in online multiple testing, leading to the development of the $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$ methods. We extend the idea to the feedback-enhanced setting. Under local dependence, the dependence–adjusted FDP estimates for GAIF and Adaptive GAIF are, respectively:
$\widehat{\operatorname{FDP}}_{\mathrm{GAIF}_{\mathrm{dep}}}(t)$ $=\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\alpha_{j}+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}}{\bigl(\sum_{j\leq t,\;j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\bigr)\vee 1},$
and
$\widehat{\operatorname{FDP}}_{\mathrm{Ada\text{-}GAIF}_{\mathrm{dep}}}(t)=\frac{\sum_{j<t-L_{t}}\frac{\alpha_{j}\mathbb{I}\{p_{j}>\lambda\}}{1-\lambda}\mathcal{W}_{j}+\sum_{t-L_{t}\leq j\leq t}\frac{\alpha_{j}}{1-\lambda}\mathcal{W}_{j}}{\bigl(\sum_{j\leq t,\;j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\bigr)\vee 1},$
where $\mathcal{W}_{j}=(1-\theta_{j})\mathbb{I}\{j\in\mathcal{I}_{t}\}+\mathbb{I}\{j\in\bar{\mathcal{I}}_{t}\}$. The corresponding testing levels are chosen so that $\widehat{\mathrm{FDP}}_{\mathrm{GAIF}_{\mathrm{dep}}}(t)\leq\alpha$ or $\widehat{\mathrm{FDP}}_{\mathrm{Ada\text{-}GAIF}_{\mathrm{dep}}}(t)\leq\alpha$. The resulting testing levels are provided in Appendix A.2. Theorem 3 shows that $\text{GAIF}_{\rm dep}$ and $\text{Ada-GAIF}_{\rm dep}$ methods control mFDR under local dependence.
Theorem 3 (Online mFDR control under local dependence).
Suppose that the null $p$-values are locally dependent as defined in (11). Then if the test levels $\{\alpha_{t}\}_{t\in\mathbb{N}}$ satisfy $\widehat{\operatorname{FDP}}_{\operatorname{GAIF}_{\rm{dep}}}(t)\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\operatorname{Ada-GAIF}_{\rm{dep}}}(t)\leq\alpha$, we have $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
4 Applications of GAIF on Online Conformal Selection
Online conformal selection [Huo et al., 2024, Xu and Ramdas, 2024] provides a canonical and intuitive instantiation of the GAIF framework. Here, decisions must be made in real time about whether an incoming observation satisfies a pre-specified requirement, while the true label is typically observed after the decision and thus serves as feedback. Throughout this section, we focus on the full and instant feedback setting.
Consider a data pair $({\bf X},Y)\in{\mathcal{X}}\times{\mathcal{Y}}$, with a historical calibration dataset $\mathcal{D}_{{\mathcal{C}}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$ of size $n$, with index set $\mathcal{C}$}. Test samples $({\bf X}_{t},Y_{t})$ arrive sequentially for $t=1,2,\dots$. At each time $t$, the covariate ${\bf X}_{t}$ is observed but responses $Y_{t}$ remain hidden until a real-time decision $\delta_{t}\in\{0,1\}$ is made. The goal is to determine whether $Y_{t}$ lies in a target region ${\mathcal{A}}\subseteq{\mathcal{Y}}$, for example, ${\mathcal{A}}=[a,b]$ or ${\mathcal{A}}=[b,\infty)$ in regression tasks. This can be framed as an online testing problem with $\theta_{t}=\mathbb{I}\{Y_{t}\in{\mathcal{A}}\}$.
We apply the GAIF framework to online conformal selection as follows. Since no off-the-shelf $p$-values are available, we first construct feedback-assisted online conformal $p$-values in Section 4.1. We then incorporate them into a modified GAIF (or Adaptive GAIF) rule to obtain finite-sample mFDR control in Section 4.2. We further use feedback to adaptively select the score model for conformal testing in Section 4.3, and analyze its optimality in Section 4.4.
4.1 Construction of online conformal $p$-values
We begin by constructing suitable online $p$-values for this online conformal selection setting. Let $V({\bf X})$ be a non-conformity score, where larger values indicate a higher likelihood that $\theta_{i}=0$. Typically, $V({\bf X})$ is a monotone transformation of the prediction $\widehat{\mu}({\bf X})$, assumed pre-trained to estimate $Y_{t}$. For example, if $\theta=\mathbb{I}\{Y_{t}\geq b\}$, one can take $V(X)=b-\widehat{\mu}({\bf X})$. For simplicity, we write $V_{i}=V({\bf X}_{i})$.
A natural approach to achieving online FDR control is to compute conformal $p$-values with the given calibration set $\mathcal{C}$} [Bates et al., 2023, Jin and Candès, 2023] and then apply GAIF or existing GAI rules. However, standard conformal $p$-values are not suitable here, since they do not satisfy the conditional super-uniformity in (6) and independence assumptions. To circumvent these issues, we adopt online conformal $p$-values, which are constructed by sequentially updating the calibration set. Specifically, at time $t$, the current point $({\bf X}_{t},Y_{t})$ serves as the test point, but from time $t+1$ onward, it is added to the calibration set. Let ${\mathcal{C}}_{0t}$ denote the dynamically updated index set of calibration samples with
${\mathcal{C}}_{0t}=\{-n<i<t:\theta_{i}=0\}.$
For test data ${\bf X}_{t}$ at time $t$, the online conformal $p$-value $p_{t}$ is defined as
$p_{t}=\frac{\sum_{i\in{{\mathcal{C}}}_{0t}}{\mathbb{I}}\{V_{i}<V_{t}\}+\xi_{t}\cdot\left[1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\{V_{i}=V_{t}\}\right]}{1+|{\mathcal{C}}_{0t}|},$ (12)
where $\xi_{t}\overset{\text{i.i.d.}}{\sim}\operatorname{Unif}[0,1]$ is an independent random variable for tie-breaking. Similar forms of online conformal $p$-values have been studied for testing exchangeability [Vovk et al., 2003, Vovk, 2021] and constructing online conformal prediction intervals [Angelopoulos et al., 2024].
Under exchangeability, this construction yields null $p$-values that are independent, thereby avoiding the complex dependence induced by structures inherent in offline conformal methods, which rely on shared calibration sets and produce $p$-values that are PRDS [Bates et al., 2023]. We impose the following assumption.
Assumption 1 (Exchangeability in conformal setting).
The null data $\big\{({\bf X}_{i},Y_{i}):i>-n,\theta_{i}=0\big\}$ are exchangeable conditional on the non-null data $\big\{({\bf X}_{i},Y_{i}):i>-n,\theta_{i}=1\big\}$.
This assumption is common in conformal inference [Marandon et al., 2024], and is weaker than requiring that the full data $\{({\bf X}_{i},Y_{i})\}_{i>-n}$ are independent and identically distributed (i.i.d.). We formally state the validity and independence under the null in Proposition 4.1.
Proposition 4.1 (Validity and independence of online conformal $p$-values).
Suppose Assumption 1 holds. Then under the null, the online conformal $p$-value $p_{t}$ defined in (12) is uniformly distributed on $[0,1]$, and the null p-values $\{p_{t}:t\in\mathbb{N},\theta_{t}=0\}$ are mutually independent.
The mutual independence above indeed relies on the full and instant feedback setting, since only in this case does ${\mathcal{C}}_{0t}$ contain the complete set of past null samples, allowing the sequential rank-based independence argument to hold. For delayed feedback, additional adjustments are needed; see the finite-sample analysis in Appendix C.1. We also allow the score function to vary over time and be data-dependent, provided that it remains symmetric on the null samples; see Appendix F.5. This flexibility enables leveraging non-null feedback to improve score quality, facilitating score selection and adaptation to distribution shift.
4.2 Online conformal testing with feedback
Having established the properties of online conformal $p$-values, we now introduce the corresponding testing procedures. Although applying GAIF directly to those $p$-values yields satisfactory FDR control empirically, a finite-sample theoretical guarantee remains lacking. The key challenge is that null online conformal $p$-values may still be dependent on non-null decisions through the adaptive testing levels.
To overcome this issue, we propose a modified GAIF rule tailored for online conformal testing, offering rigorous, distribution-free, and model-agnostic mFDR guarantees. The key idea is to construct the sequence $\{\alpha_{t}\}_{t\in\mathbb{N}}$ based solely on the rejections of true nulls, rather than on all rejections. While this strategy may be conservative as discarding non-null decision information, the integration of the feedback provides substantial compensation.
More specifically, let $\tau_{j}$ be the time of the $j$-th rejection, and denote $\tilde{\tau}_{j}$ as the time of the $j$-th rejection under the null, defined as
$\tilde{\tau}_{j}=\inf\left\{t\in\mathbb{N}:\sum_{i\leq t}\delta_{i}(1-\theta_{i})\geq j\right\}.$
We obtain the safe versions of LF and SF by replacing $\tau_{j}$ with $\tilde{\tau}_{j}$, leading to the LFS and SFS procedures, respectively. The letter “S" stands for “safe", since it guarantees rigorous finite-sample mFDR control. By doing so, $\tilde{\tau}_{j}$ is fixed given past null p-values, as its construction depends only on the rejections of true null hypotheses. The Online Conformal Testing with Feedback (OCTF) procedureis detailed in Algorithm 2.
Algorithm 2 Online conformal testing with feedback (OCTF)
0:  Initial calibration data ${\mathcal{D}}_{\mathcal{C}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, non-conformity score function $V(\cdot)$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}_{t\in\mathbb{N}}$, constant $\lambda\in(0,1)$.
1:  Initialize ${\mathcal{C}}_{01}=\{i\in{\mathcal{C}}:\theta_{i}=0\}$.
2:  for $t=1,2,\ldots$ do
3:   Observe test covariate ${\bf X}_{t}$.
4:   Compute conformity scores $\{V_{i}\}_{i\in{\mathcal{C}}_{0t}\cup\{t\}}$.
5:   Compute online conformal $p$-value $p_{t}$ via (12).
6:   Update $\alpha_{t}={\alpha}_{t}^{\text{LFS}}$ in (18) (or $\alpha_{t}={\alpha}_{t}^{\text{SFS}}$ in (19)).
7:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$.
8:   Obtain the revealed feedback $Y_{t}$, and thus obtain $\theta_{t}$.
9:   Update the calibration dataset ${\mathcal{C}}_{0t}$.
10:  end for
10:  Rejection set $\mathcal{R}=\{t:\delta_{t}=1\}$.
Theorem 4 (Finite-sample online mFDR control for OCTF).
Suppose Assumption 1 holds. The output $\mathcal{R}$} of Algorithm 2 satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$.
With respect to FDR control, the additional challenge lies in the fact that a past non-null decision $\delta_{i}=\mathbb{I}\{p_{i}\leq\alpha_{i}\}$ with $\theta_{i}=1$ may be dependent on a current null $p_{t}$. Decoupling the dependence between decisions $\sum_{i\leq t}\delta_{i}$ in the denominator and $p_{t}$ is more challenging and warrants further research.
4.3 Optimized online conformal testing with feedback
The performance of online conformal testing depends on the choice of the non-conformity scores $V({\bf X})$ and the prediction model, e.g., random forests, neural networks, support vector machines, or regularized linear models [Bai and Jin, 2024, Gasparin and Ramdas, 2024]. In practice, the most effective score may also change over time, especially when the non-null distribution drifts [Gang et al., 2023, Huo et al., 2024]. This motivates a feedback-driven score-selection step built on top of the OCTF framework.
Suppose there are $K$ pre-trained candidate score functions $\{V(\cdot;k):{\mathcal{X}}\to\mathbb{R}\}_{k=1}^{K}$. At each time $t$, a score is chosen as
$\widehat{k}_{t}=\underset{k\in[K]}{\arg\min}\;{\mathcal{M}}({\mathcal{D}}_{t};k),$
where ${\mathcal{M}}({\mathcal{D}}_{t};k)$ is a given criterion evaluated by the currently observed data ${\mathcal{D}}_{t}=\{({\bf X}_{i},Y_{i}):-n<i<t\}\cup\{{\bf X}_{t}\}$. Our goal is to choose the score that is most favorable for detecting non-nulls, which motivates the design of a reliable criterion ${\mathcal{M}}({\mathcal{D}}_{t};k)$. Since the testing threshold $\alpha_{t}$ depends recursively on the past rejection path, directly optimizing rejection probability is difficult. We therefore use the proxy
${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1],$
where $p_{k,t}$ is the $t$-th conformal $p$-value obtained from the score function $V(\cdot;k)$. Smaller non-null $p$-values generally correspond to higher rejection probability $\Pr(p_{k,t}\leq\alpha_{t}\mid\theta_{t}=1)$, so this provides a natural threshold-free criterion.
To estimate ${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1]$ under shifted non-null distribution while preserving symmetry, we employ an exponentially weighted moving average (EWMA) of past auxiliary non-null $p$-values. Specifically, at time $t$, for each score function $k$, we construct auxiliary non-null $p$-values as
$\tilde{p}_{k,j}=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|},\quad j\in{\mathcal{C}}_{1t},$ (13)
where ${\mathcal{C}}_{1t}=\{-n<i<t:\theta_{i}=1\}$ denotes the set of online non-null samples observed prior to time $t$. When the non-null distribution varies slowly over time, these auxiliary non-null $p$-values are approximately distributed like the current non-null $p$-value $p_{k,t}$. The construction in (13) is carefully designed so that the resulting auxiliary non-null $p$-values are invariant under permutations of ${\mathcal{C}}_{0t}\cup\{t\}$, which is crucial for ensuring valid inference after score selection. We then define the EWMA criterion as
${\mathcal{M}}^{\mathrm{EWMA}}_{t}({\mathcal{D}}_{t};k)=\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,\tilde{p}_{k,j}\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}},$
where $\rho\in(0,1)$ is the user-specified decay parameter that downweights past observations, enabling dynamic adaptation to the recent non‑null distribution and precise estimation of ${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1]$. The selected score function is accordingly as $\widehat{k}_{t}={\arg\min}_{k\in[K]}{\mathcal{M}}^{\mathrm{EWMA}}_{t}({\mathcal{D}}_{t};k).$
The online conformal $p$-value after score selection, denoted as $p^{\rm opt}_{t}$, is then computed by (12) using the selected score $V(\mathbf{X}_{t};\widehat{k}_{t})$. Then running the OCTF procedures in Section 4.2 with these optimized $p$-values yields the whole optimized procedure, which we refer to as Opt-OCTF algorithm.
Corollary 4.1 (Finite-sample online mFDR control for Opt-OCTF).
Under the assumptions in Theorem 4, the Opt-OCTF procedure by ${\mathcal{M}}^{\rm EWMA}_{t}$ satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}.$
Although Corollary 4.1 is stated for the EWMA criterion, the underlying framework for optimized online conformal testing can be extended more broadly. In particular, we allow more general selection criteria $\mathcal{M}$} by using a full permutation strategy over ${\mathcal{C}}_{0t}\cup\{t\}$; see Appendix C.2. In addition, data-driven score functions obtained by online updating are also compatible with the framework under a suitable symmetry-preserving training scheme. However, we focus on selection among $K$ pre-trained candidates in the main text for computational reasons; see Appendix C.3 for details.
Remark 4.1.
In (13), we include the current time point $t$ in the null calibration set to preserve symmetry over ${\mathcal{C}}_{0t}\cup\{t\}$, which is essential for the validity of the optimized conformal $p$-value after score selection. Although this may slightly affect calibration when $t$ is non-null, the effect is typically minor in practice. A simple truncated variant can further reduce this influence while preserving validity; see Appendix E.5 for details.
4.4 Optimality analysis of EWMA criterion
We now give a supporting analysis for the EWMA criterion by establishing certain optimality of the online score selection step. To facilitate analysis, we work under the standard two-group model, which is widely used to study the performance of online testing procedures [Gang et al., 2023, Humbert et al., 2025].
Assumption 2 (Two-group model).
Suppose the score functions follow the two-group model for each $k\in[K]$:
$V({\bf X}_{t};k)\overset{\rm indep}{\sim}\pi_{t}F_{k}^{(0)}+(1-\pi_{t})F^{(1)}_{k,t},\quad\theta_{t}\overset{\rm indep}{\sim}\text{Bernoulli}(1-\pi_{t}),$
where $F_{k}^{(0)}$ and $F^{(1)}_{k,t}$ are the null and (time-varying) non-null score distributions for $k$-th model.
Under this model, the oracle conformal $p$-value associated with score $V(\cdot;k)$ is $F_{k}^{(0)}(V({\bf X}_{t};k))$. Conditioned on $\theta_{t}=1$, smaller value of $F_{k}^{(0)}(V({\bf X}_{t};k))$ indicates higher detection power. And we define the optimal score function at time $t$ as
$k_{t}^{*}=\arg\min_{k\in[K]}{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1].$
Note that ${\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1]$ is the population analogue of the empirical proxy ${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1]$. To show that the EWMA criterion tracks this oracle target and selects ${k}^{*}_{t}$ reliably, we impose standard regularity conditions.
Assumption 3 (Slowly varying distribution).
For each $k\in[K]$ and $t\in\mathbb{N}$, $F^{(1)}_{k,t}$ has a slowly varying distribution $\|F^{(1)}_{k,t}-F^{(1)}_{k,t-1}\|_{\infty}\leq\gamma$ such that $\gamma\to 0$.
Assumption 4 (Non-degeneracy).
There exist constants $0<\underline{\pi}\leq\overline{\pi}<1$ such that null proportion satisfies $\pi_{t}\in[\underline{\pi},\overline{\pi}]$.
Assumption 5 (Margin).
For any $t$ and a constant $c>0$, $\min_{i\neq k_{t}^{*}}|{\mathbb{E}}[F_{i}^{(0)}(V({\bf X}_{t};i))\mid\theta_{t}=1]-{\mathbb{E}}[F_{k_{t}^{*}}^{(0)}(V({\bf X}_{t};k_{t}^{*}))\mid\theta_{t}=1]|\geq c$.
Assumption 3 allows the non-null distribution $F^{(1)}_{k,t}$ to drift over time. Assumption 4 ensures that neither nulls nor non-nulls become too rare over time. Assumption 5 guarantees the optimal score is identifiable.
Theorem 5 (Optimality of online score selection).
Suppose Assumptions 2-4 hold.
  1. (i)
    If $t+n\geq 1+1/(1-\rho)$, then for any $\varepsilon_{1}\in\left(0,\sqrt{\underline{\pi}/2}\right)$ and $\varepsilon_{2}\in(0,1-\rho)$,
    $\Pr\left(|{\mathcal{M}}^{\rm EWMA}(k,{\mathcal{D}}_{t})-{\mathbb{E}}[F^{k}_{0}(V({\bf X}_{t};k))\mid\theta_{t}=1]|\leq\varepsilon_{1}+\varepsilon_{2}+\frac{2\gamma}{(1-\overline{\pi})(1-\rho)}+\frac{2}{(t+n-1)\underline{\pi}}\right)$
    $\geq 1-3\exp\left\{-(t+n-1)\underline{\pi}\varepsilon_{1}^{2}\right\}-4\exp\left\{-\frac{\varepsilon_{2}^{2}({1-\overline{\pi}})^{2}}{6(1-\rho)}\right\}.$
    Further assume that Assumption 5 holds. Then the following results in (ii)-(iii) hold.
  2. (ii)
    If $t+n\geq 1+{16}/({c\underline{\pi}})$ and ${\gamma}/({1-\rho})\leq{(1-\overline{\pi})c}/{32}$, we have the finite-sample error bound for the score selection
    $\Pr(\hat{k}_{t}\neq k_{t}^{*}\mid\theta_{t}=1)\leq\;3K\exp\!\left\{-(t+n-1)\underline{\pi}\frac{c^{2}}{64}\right\}+4K\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}.$
  3. (iii)
    Moreover, the score selection consistency for all non-null points holds as
    $\Pr(\exists i\leq t:\theta_{i}=1,\ \hat{k}_{i}\neq k_{i}^{*})\leq\frac{3K(1-\underline{\pi})\exp\!\left\{-n\underline{\pi}{c^{2}}/{64}\right\}}{1-\exp\!\left\{-\underline{\pi}{c^{2}}/{64}\right\}}+4K(1-\underline{\pi})t\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}.$
    In particular, if $(1-\rho)\log(t)\to 0$, $\gamma/(1-\rho)\to 0$ and $n\to\infty$, $\Pr(\exists i\leq t:\theta_{i}=1,\ \hat{k}_{i}\neq k_{i}^{*})\to 0$.
Theorem 5 (i) shows that the EWMA criterion concentrates around the oracle power proxy. The bound decomposes naturally into: an estimation term scaling with $t+n$, a non-null effective size term in EWMA operation by $(1-\rho)^{-1}$ and a drift bias of order $\gamma/(1-\rho)$ reflecting how well the EWMA estimator can track a drifting non-null distribution. Parts (ii)–(iii) translate this concentration into optimal score selection guarantees. The rate conditions reflect the classic EWMA trade-off: $\rho$ must be sufficiently close to 1 to ensure an adequate effective sample size, yet small enough to allow the procedure to adapt to distribution drift.
5 Numerical Experiments
In this section, we present extensive synthetic experiments to demonstrate the validity and efficiency of our proposed methods. First, Section 5.1 studies GAIF and Adaptive GAIF in the non-conformal setting. Second, Sections 5.2 and 5.3 present the results for OCTF and Opt-OCTF under the conformal setting, respectively. Following the setup in Robertson et al. [2023], we fix $\lambda=0.5$, $\gamma_{j}\propto j^{-1.6}$ for all $j\in\mathbb{N}$, and $s_{0}=\alpha/2$ throughout.
5.1 Synthetic experiments under non-conformal settings
We begin by presenting experiments on online multiple testing problems under three scenarios. Throughout, let $\pi_{1}$ denote the non-null proportion parameter in the data-generating model; its value varies across experiments.
  • Scenario I (Testing with Gaussian observations) We simulate $T$ independent test statistics $Z_{t}\sim N(\mu_{t},1)$ and test hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$ using one-sided $p$-values $p_{t}=\Phi(-Z_{t})$, where $\Phi$ is the standard Gaussian CDF. The signal strengths $\mu_{t}$ follow a mixture model:
    $\mu_{t}=\begin{cases}0\quad\text{with probability}\;1-\pi_{1}\\ F_{1}\quad\text{with probability}\;\pi_{1},\end{cases}$ (14)
    where the random variable $F_{1}\sim N(2.5,1)$.
  • Scenario II (Testing with Beta alternatives) We generate $p$-values according to the following model:
    $p_{t}\sim\begin{cases}\text{Unif}[0,1]\quad\text{with probability}\;1-\pi_{1}\\ \text{Beta}(0.5,4)\quad\text{with probability}\;\pi_{1}.\end{cases}$ (15)
  • Scenario III (Testing under local dependence) We simulate correlated test statistics $(Z_{1},\dots,Z_{T})^{\top}\sim N(\bm{\mu},\Sigma)$ where $\bm{\mu}=(\mu_{1},\ldots,\mu_{T})^{\top}$ with $\mu_{t}=2$ for randomly a fraction $\pi_{1}$ of indices and $\mu_{t}=0$ otherwise. The covariance matrix $\Sigma$ has a block-diagonal structure: coordinates are grouped into blocks of size $n_{\mathrm{block}}=10$, with within-block correlation $\rho=0.8$ and diagonal elements as 1. We test hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$ using one-sided $p$-values $p_{t}=\Phi(-Z_{t})$.
We set the stopping time as $T=1000$ and the FDR level at $\alpha=0.1$. We evaluate the performance via empirical FDR and power averaged over 500 independent replications. Appendix E confirms that mFDR and FDR show similar trends.
Firstly, we consider the full and instant feedback setting, and compare our methods SF and LF with state-of-the-art algorithms for online FDR control, namely LOND [Javanmard and Montanari, 2015], LORD++, and SAFFRON, using their default parameters ($\lambda=0.5$ chosen for SAFFRON). Figure 2 shows results for varying non-null proportion $\pi_{1}\in[0.1,0.8]$ under Scenarios I and II. All benchmarks ensure valid FDR control across different $\pi_{1}$ under both Scenarios. Our feedback-based SF and LF methods enhance detection power while maintaining valid online FDR control: SF significantly outperforms SAFFRON, while LF yields higher power than LORD++ and surpasses SAFFRON when $\pi_{1}$ is small. In contrast, both LORD++ and LOND remain conservative across different values of $\pi_{1}$. The plots have almost invisible error bars, as the variation across the 500 runs is negligible.
Figure 2: Results for Scenario I and Scenario II under full and instant feedback: FDR and Power at stopping time with non-null proportion $\pi_{1}$ ranging from $0.1$ to $0.8$ across $500$ replications; The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
Figures 3 present results under a positive local dependence structure in Scenario III. The baseline methods—LOND, LORD++, and SAFFRON—are included for comparison. The proposed procedures, $\text{SF}_{\rm dep}$ and $\text{LF}_{\rm dep}$, correspond to $\text{Ada\text{-}GAIF}_{\rm dep}$ and $\rm GAIF_{\rm dep}$ under full and instant feedback, respectively. In addition, we include the existing methods $\text{SAFFRON}_{\rm dep}$ and $\text{LORD}_{\rm dep}$ from Zrnic et al. [2021]. We find that SF, LF, and SAFFRON fail to control the FDR under dependence. In contrast, the dependence-aware procedures $\mathrm{SF}_{\mathrm{dep}}$ and $\mathrm{LF}_{\mathrm{dep}}$ successfully achieve valid FDR control and attain substantially higher power than their feedback-ignoring counterparts, $\mathrm{SAFFRON}_{\mathrm{dep}}$ and $\mathrm{LORD}_{\mathrm{dep}}$.
Figure 3: Results for Scenario III under full and instant feedback: FDR and Power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
Next, we evaluate the performance of our proposed methods under other feedback settings: bandit and instant, and full and delayed. Taking the classical methods LOND, LORD++, and SAFFRON as benchmarks, we implement the proposed LF-BI and SF-BI under bandit and instant feedback setting. The results in Scenario I are shown in Figure 4. All methods maintain valid FDR control. Importantly, bandit feedback continues to enhance detection power, although the improvement is less pronounced than that under full and instant feedback.
Figure 4: Results for Scenario I under bandit and instant feedback: FDR and Power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
For the full and delayed feedback setting, we consider the delay $d\in\{10,100,200\}$ and compare the proposed SF-FD and LF-FD with the baselines without consideration of feedback. Figure 5 displays the results in Scenario I. As the delay $d$ increases, the detection power of our methods decreases slightly, yet they still outperform the baselines by a substantial margin while maintaining valid FDR control.
Figure 5: Results for Scenario I under full and delayed feedback: FDR and power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
5.2 Results for online conformal testing with feedback
We next evaluate OCTF with a fixed score in a real-time binary classification task.
  • Scenario IV: Data is from $\mathbf{X}\mid Y=0\sim\mathcal{N}_{4}\left(\bm{\mu}_{1},\mathbf{I}_{4}\right)$, and $\mathbf{X}\mid Y=1\sim\mathcal{N}_{4}\left(\bm{\mu}_{2},\mathbf{I}_{4}\right)$, where $\bm{\mu}_{1}=(1,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. The target region is $\mathcal{A}=\{1\}$.
We set the stopping time to $T=1000$ and the target $\operatorname{FDR}$ level to $\alpha=0.2$. The initial calibration size is $n_{\text{cal}}=1000$. We consider the score as $V(\mathbf{X})=1-\hat{\mu}(\mathbf{X})$, where the predictive model $\hat{\mu}(\mathbf{X})$ is a neural network classifier trained on $n_{\text{tr}}=1000$. Appendix E reports additional results for varying $n_{\mathrm{cal}}$, alternative training algorithms, and a regression setting.
We evaluate the performance of the proposed LFS and SFS and implement the benchmarks LF, SF, LOND, LORD++ and SAFFRON within the OCTF workflow with their own testing levels $\{\alpha_{t}\}$. Figure 6 reports the empirical FDR and power for varying non-null proportions $\pi_{1}\in[0.1,0.8]$. All methods control the FDR below the nominal level $\alpha$. As expected, the feedback-enhanced procedures achieve clear power gains. It is noted that SF consistently achieves the highest power, while LF also performs competitively and attains higher power than SAFFRON. SFS and LFS outperform LORD++ and LOND owing to the incorporation of feedback information. This confirms that the safe conformal construction remains effective in practice, despite discarding certain rejections under the non-null for finite-sample guarantee.
Figure 6: Results for Scenario IV: FDR and power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed lines denote the FDR level $\alpha=0.2$. Shaded areas show $\pm 1$ standard error.
5.3 Results for optimized online conformal testing with feedback
We now demonstrate the effectiveness of the proposed score selection strategy from Section 4.3. The goal here is to examine whether feedback can be used not only for valid testing, but also for adaptive score selection when non-null distributions shift.
We employ the binary classification example in Scenario IV with $\bm{\mu}_{1}=(2,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. For the historical labeled dataset, the non-null proportion is fixed as $0.2$. For the testing data, we consider a sine pattern shifts, where the non-null proportion varies as $\pi_{t}=\operatorname{Pr}(Y_{t}=1)=\{\sin(8\pi t/T)+1\}/4$, oscillating between 0 and 0.5.
For the implementation of the score selection, we consider $K=3$ candidate models: random forest (RF), neural network (NN), and support vector machine (SVM) and employ the EWMA criterion ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with exponential weight parameter $\rho=0.95$. To reduce computational cost, the auxiliary non-null $p$-value is computed by a sliding window of fixed length $L=200$: for $t>L$, we define ${\mathcal{C}}_{1t}=\{i:t-L\leq i<t,\theta_{i}=1\}$; for $t\leq L$, we use the full set ${\mathcal{C}}_{1t}=\{i<t:\theta_{i}=1\}$.
To evaluate the score-selection step, we compare the proposed optimized rule with random score selection under different testing procedures. Here, “Opt-” means that the score model is selected adaptively by the proposed EWMA criterion, while “Ran-” means that the score model is chosen uniformly at random from the candidate library. For example, Opt-LFS denotes the optimized version of LFS, whereas Ran-LFS uses the same testing rule with random score selection.
Figure 7 shows the results under distribution shift. All methods successfully control empirical FDR around the nominal level $\alpha=0.1$. In terms of power, the optimized (Opt) variants consistently outperform their randomly selected (Ran) counterparts, implying that optimized score selection is useful under distribution shift. In addition, the LFS-based and SFS-based procedures exhibit lower power than LF-based and SF-based, respectively. However, in the next section, we find that the SF-based method may suffer from FDR inflation on real data due to the lack of finite-sample theoretical guarantees. Meanwhile, additional results for scenarios without distribution shifts are presented in Appendix E, and also highlight the effectiveness of our optimized score selection strategy.
Figure 7: Results for Scenario V (sine pattern shifts): $\operatorname{FDR}(t)$ and $\operatorname{Power}(t)$ across different time $t$. The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
6 Real Data Applications
In this section, we evaluate our proposed methods on four real-world datasets to illustrate their practical benefits in diverse online decision-making tasks.
  • Task 1: Online Candidate Screening. The first task is real-time candidate screening for selecting the candidates who can get into the interview process. We use the recruitment dataset [Kaggle, 2020], which contains $N=45,372$ candidates with 11 attributes, including education status, handicapped or not, and gender. The target binary variable is whether the candidate passes the job interview.
  • Task 2: High-Risk Diabetes Identification. The second task focuses on health screening, where selecting individuals at high risk of diabetes is critical for early intervention. We use the diabetes health indicators dataset [Kaggle, 2021], which contains $N=70,692$ samples and 22 covariates, including demographic attributes (e.g., sex, age, BMI), lifestyle-related features, and several binary health indicators. The target binary variable is whether an individual has diabetes.
  • Task 3: High-Income Individual Selection. The third task involves using the 1994 Census Bureau dataset [Becker,Barry and Kohavi,Ronny, 1996] to identify a subset of individuals with high incomes (i.e., income >50K>50K) for precision marketing. This dataset contains $N=32,561$ records with 14 attributes, including gender, race, marital status, education level, and more.
  • Task 4: Airfoil Noise Detection. The final task involves using the airfoil dataset [Brooks,Thomas et al., 2014] from the UCI Machine Learning Repository to identify samples with high sound pressure. This dataset contains $N=1,503$ observations with five physical covariates (log frequency, angle of attack, chord length, free-stream velocity, and suction side log displacement thickness). The response variable $Y$ represents the scaled sound pressure, and we aim to test $\mathbb{H}_{0t}:Y_{t}\in[-\infty,c)$, where $c$ is the $(1-\pi_{1})$-quantile of $Y$, with $\pi_{1}=0.4$.
For Tasks 1-3, we randomly sample three subsets from the full dataset: $n_{\text{tr}}=1{,}000$ training, $n_{\text{cal}}=1{,}000$ calibration, and $n_{\text{te}}=1{,}000$ test samples. For Task 4, we set $n_{\text{tr}}=300$, $n_{\text{cal}}=300$, and $n_{\text{te}}=1{,}000$. We compare four optimized procedures Opt-SF, Opt-SFS, Opt-LF and Opt-LFS with two benchmarks SAFFRON and LORD++. At each time step, $K=3$ candidate training algorithms (RF, NN, and SVM) are employed, and the score selection criterion is ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with $\rho=0.9$. Our methods use the optimized conformity score, while the benchmarks use a randomly selected score at each time $t$. All results are averaged over 500 replications.
Figure 8 summarizes the empirical FDR and power over time $t$, and Table 1 reports the corresponding values at the stopping time $T=1,000$ at the target FDR level $\alpha=0.3$. For Tasks 1-3, all methods control the empirical FDR near or below the nominal level, and Opt‑SF consistently achieves the highest power as expected. However, in task 4, Opt-SF shows mild FDR inflation, whereas the safe procedures Opt-SFS and Opt-LFS maintain valid control throughout. This is consistent with our theory in Section 4.3: the safe variants are more conservative, but they provide the finite-sample guarantees needed in difficult settings. Notably, SAFFRON and LORD++ deliver substantially lower power than our optimized methods in all tasks, underscoring the strength of the OCTF framework.
Overall, the proposed feedback-enhanced methods achieve higher power than feedback-ignoring baselines, while the safe variants maintain reliable error control across all four tasks.
Figure 8: Results for real-data applications: the values of $\operatorname{FDR}(\bm{\delta}^{t})$ and $\operatorname{Power}(\bm{\delta}^{t})$ over time $t$ for six benchmarks. The black dashed lines indicate the FDR level $\alpha=0.3$. Shaded areas show $\pm 1$ standard error.
Table 1: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ for different tasks across four datasets (Candidate, Diabetes, Income, Airfoil). The target FDR level is $\alpha=0.3$. Bold numbers represent the best results.
Method Task 1 Task 2 Task 3 Task 4
FDR Power FDR Power FDR Power FDR Power
Opt-SF 0.161 0.247 0.238 0.517 0.287 0.398 0.363 0.726
Opt-SFS 0.124 0.083 0.141 0.166 0.132 0.168 0.239 0.440
Opt-LF 0.148 0.221 0.188 0.341 0.254 0.336 0.242 0.581
Opt-LFS 0.127 0.103 0.132 0.150 0.117 0.171 0.060 0.170
SAFFRON 0.112 0.007 0.185 0.160 0.165 0.083 0.251 0.355
LORD++ 0.086 0.004 0.104 0.004 0.078 0.052 0.098 0.193
7 Concluding Remarks
We study online multiple testing with feedback, aiming to develop reliable machine learning methods for real-time decision-making with rigorous FDR and mFDR control. Our key contribution is GAIF and its adaptive variants, a feedback-enhanced Generalized Alpha-Investing framework that guarantees online FDR control under the standard assumption of independence between null and all other $p$-values. Building on GAIF, we further develop OCTF for online conformal testing by constructing independent online conformal $p$-values for null hypotheses, which achieves finite-sample mFDR control. To further enhance performance and address distribution shifts among non-nulls, we propose a feedback-driven score selection strategy and provide theoretical analysis characterizing its optimality properties.
We highlight several potential directions for future work. First, the current framework mainly focuses on distribution shifts in the non-null data. To address alpha-death and piggybacking under such shifts, one can consider controlling weighted FDR using forgetting factors or decaying memory mechanism [Ramdas et al., 2017]. Second, extending our framework to accommodate more general forms of distribution shift is another promising direction. One potential approach is to incorporate adaptive strategies that exploit local information, in the spirit of the adaptive conformal inference [Gibbs and Candès, 2021]. Finally, to establish guarantees under weaker assumptions, one could relax the definition of valid online FDR control and develop new error-rate notions, analogous to the average-coverage criterion studied in online conformal inference [Gibbs and Candès, 2021, Humbert et al., 2025].
Acknowledgments
We would like to acknowledge the action editor and three anonymous referees for their valuable comments and suggestions, which have improved the manuscript greatly.
Supplementary Material for “Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection”
This supplementary material includes:
  • Notation and Algorithmic Details (Appendix A)
  • Additional Discussions on GAIF and Adaptive GAIF (Appendix B)
  • Additional Discussions on OCTF (Appendix C)
  • Applications on Real-time LLM Alignment (Appendix D)
  • Additional Experiments Results (Appendix E)
  • Proofs of all the theoretical results (Appendix F)
Appendix A Notation and Algorithmic Details
This appendix gathers the notation and additional algorithmic details used throughout the paper. For completeness, we first review several preliminary terms and benchmark procedures, and then provide the concrete forms of the proposed testing levels and related implementation details.
A.1 Preliminary terms for self-containment
Here, we list the preliminary terms we use in the paper for the sake of clarity and self-containment.
  • $\operatorname{FDR}$ [Benjamini and Hochberg, 1995], false discovery rate, a widely-adopted error rate notion in the field of multiple testing, is defined as the expected proportion of incorrectly rejected null hypotheses as follows:
    $\operatorname{FDP}(t)=\frac{V_{t}}{1\vee R_{t}}:=\frac{\sum_{j=1}^{t}\delta_{j}(1-\theta_{j})}{1\vee\sum_{j=1}^{t}\delta_{j}}\quad\text{and}\quad\operatorname{FDR}(t):={\mathbb{E}}[\operatorname{FDP}(t)],$
    where $R(t)$ represents the number of rejected null hypotheses until time $t$ and $V(t)$ is the number of false discoveries.
  • $\operatorname{mFDR}$, modified false discovery rate, or marginal false discovery rate, is defined as:
    $\operatorname{mFDR}(t):=\frac{{\mathbb{E}}[V(t)]}{{\mathbb{E}}[1\vee R(t)]}=\frac{{\mathbb{E}}[\sum_{j=1}^{t}\delta_{j}(1-\theta_{j})]}{{\mathbb{E}}[1\vee\sum_{j=1}^{t}\delta_{j}]}.$
  • Testing levels of existing online multiple testing methods. For the online methods, denote the decision rule as $\delta_{t}=\mathbb{\{}p_{t}\leq\alpha_{t}\}$, where $p_{t}$ is the corresponding conformal $p$-value at time $t$ for our problem. The test levels $\{\alpha_{t}\}$ for LOND [Javanmard and Montanari, 2015], LORD++ [Ramdas et al., 2017], SAFFRON [Ramdas et al., 2018] and ADDIS [Tian and Ramdas, 2019], $\text{LORD}_{\text{dep}}$ [Zrnic et al., 2021], $\text{SAFFRON}_{\text{dep}}$ [Zrnic et al., 2021] are listed as follows:
    • 1.
      LOND: $\alpha_{t}=\gamma_{t}(R(t-1)+1)$, where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to $\alpha$ and $R(n)=\sum_{t=1}^{n}R_{t}$ denotes the number of discoveries in the first $n$ hypotheses tested.
    • 2.
      LORD++:
      $\alpha_{t}^{\text{LORD++}}=\gamma_{t}W_{0}+(\alpha-W_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}},$
      where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to one; $\tau_{j}$ is the time of the $j$-th rejection.
    • 3.
      SAFFRON: At each time $t$, define $C_{j+}=C_{j+}(t)=\sum_{i=\tau_{j}+1}^{t-1}$, where $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$. For $t=1$, $\alpha_{1}=\min\{\gamma_{1}W_{0},\lambda\}$; For $t=2,3,\dots$, $\alpha_{t}:=\min\{\lambda,\tilde{\alpha}_{t}\}$, where
      $\tilde{\alpha}_{t}=W_{0}\gamma_{t-C_{0+}}+((1-\lambda)\alpha-W_{0})\gamma_{t-\tau_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-C_{j+}}.$
    • 4.
      ADDIS: The testing levels for ADDIS are given by $\alpha_{t}=\min\{\lambda,\hat{\alpha}_{t}\}$, where
      $\hat{\alpha}_{t}=(\eta-\lambda)[\omega_{0}\gamma_{S^{t}-C_{0+}}+(\alpha-\omega_{0})\gamma_{S^{t}-\tau_{1}^{*}-C_{1+}}+\alpha\sum_{j\geq 2}\gamma_{S^{t}-\tau_{j}^{*}-C_{j+}}]$
      and $S^{t}=\sum_{i<t}\mathbb{I}\{p_{i}\leq\eta\}$, $\tau_{j}^{*}=\sum_{i\leq\tau_{j}}\mathbb{I}\{p_{i}\leq\eta\}$.
    • 5.
      $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$: Define $r_{k}$ under local dependence as:
      $r_{k}=\min\{i\in[t]:\sum_{j=1}^{i-L_{i+1}}\delta_{j}\geq k\}.$
      The corresponding test levels for $\text{LORD}_{\text{dep}}$ and $\text{SAFFRON}_{\text{dep}}$ are as follows:
      $\alpha_{t}^{\text{LORD}_{\text{dep}}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-r_{1}}\mathbb{I}\{r_{1}<t\}+\alpha\sum_{j=2}^{\infty}\gamma_{t-r_{j}}.$
      $\alpha_{t}^{\text{SAFFRON}_{\text{dep}}}:=\min\left\{\lambda,(1-\lambda)\left(w_{0}\gamma_{t-C_{0+}}+(\alpha-w_{0})\gamma_{t-r_{1}-C_{1+}}+\alpha(\sum_{j\geq 2}\gamma_{t-r_{j}-C_{j+}})\right)\right\},$
      where $C_{j+}=\sum_{i=r_{j}+1}^{t-L_{t}+1}C_{i}$.
  • Conformal $p$-values. The notion of conformal $p$-values was originally introduced by Vovk et al. [2005] for constructing prediction intervals. A conformal $p$-value quantifies how well a new observation conforms to a reference set, based on a chosen nonconformity score function. More recently, several works have applied conformal $p$-values to sample selection from a multiple testing perspective [Jin and Candès, 2025, Bates et al., 2023, Wang et al., 2024]. The conformal $p$-values are defined as
    $p_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0}}{\mathbb{I}}\{V_{i}\leq V_{t}\}}{1+|{\mathcal{C}}_{0}|},$ (16)
    where ${\mathcal{C}}_{0}$ is a hold-out calibration dataset and $V(\cdot)$ is a nonconformity score function. However, in these approaches, the conformal $p$-values are constructed using a fixed offline calibration dataset, which limits their flexibility in online or adaptive settings.
A.2 Testing levels of the proposed approaches
  • 1.
    Testing levels for $\text{GAIF}_{\text{dep}}$ and $\text{Ada-GAIF}_{\text{dep}}$:
    Define $r_{k}$ under local dependence as:
    $r_{k}=\min\{i\in[t]:\sum_{j=1}^{i-L_{i+1}}\delta_{j}\geq k\}.$
    The corresponding test levels for $\text{LF}_{\text{dep}}$ and $\text{SF}_{\text{dep}}$ are constructed as follows:
    $\alpha_{t}^{\text{GAIF}_{\text{dep}}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-r_{1}}\mathbb{I}\{r_{1}<t\}+\alpha\sum_{j=2}^{\infty}\gamma_{t-r_{j}}+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j},$
    $\alpha_{t}^{\text{Ada-GAIF}_{\text{dep}}}:=\min\left\{\lambda,(1-\lambda)\left(w_{0}\gamma_{t-C_{0+}}+(\alpha-w_{0})\gamma_{t-r_{1}-C_{1+}}+\alpha(\sum_{j\geq 2}\gamma_{t-r_{j}-C_{j+}})\right)+\sum_{j:j\in\mathcal{I}_{t}}\gamma_{j}\alpha_{j}\theta_{j}\right\},$
    where $C_{j+}=\sum_{i=r_{j}+1}^{t-L_{t}}C_{i}$.
  • 2.
    Testing levels for LFS and SFS: Recall that the proposed test levels of GAIF under full and instant feedback, i.e., $\alpha_{t}^{\rm LF}$, are given by:
    $\alpha_{t}^{\text{LF}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}}+\sum_{j:j<t}\gamma_{j}\alpha_{j}\theta_{j},$ (17)
    where $\tau_{j}$ denotes the time of the $j$-th rejection. We revise this rule to:
    ${\alpha}_{t}^{\text{LFS}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}}\mathbb{I}\{{\tilde{\tau}}_{1}<t\}+\alpha\sum_{j:\tilde{\tau}_{j}<t,\tilde{\tau}_{j}\neq\tilde{\tau}_{1}}\gamma_{t-\tilde{\tau}_{j}}+\sum_{j:j<t}\gamma_{j}{\alpha}_{j}\theta_{j},$ (18)
    where $\tilde{\tau}_{j}$ denotes the time of the $j$-th rejection under the null, defined as
    $\tilde{\tau}_{j}=\inf\left\{t\in\mathbb{N}:\sum_{i\leq t}\delta_{i}(1-\theta_{i})\geq j\right\}.$
    And similarly, the test levels of SFS are:
    $\alpha_{t}^{\text{SFS}}:=\min\{\lambda,\tilde{\alpha}_{t}^{\text{SFS}}\},$ (19)
    where $\tilde{\alpha}_{t}^{\text{SFS}}=s_{0}\gamma_{t-C_{0+}}+((1-\lambda)\alpha-s_{0})\gamma_{t-\tilde{\tau}_{1}-C_{1+}}+(1-\lambda)\alpha\sum_{j\geq 2}\gamma_{t-\tilde{\tau}_{j}-C_{j+}}+\sum_{j:j<t}\gamma_{j}\alpha_{j}\theta_{j}.$
A.3 Choice of non-conformity score function $V$
In terms of the non-conformity score function, denote $W_{t}=\widehat{\mu}_{t}({\bf X}_{t})$, in classification settings, we set $V(W_{t})=1-W_{t}$. In regression settings, if $\mathcal{A}=[b,+\infty)$, we can use $V(W_{j})=b-W_{j}$. If $\mathcal{A}=(-\infty,a]\cup[b,+\infty)$, then we can choose $V(W_{t})=\max\{W_{t}-a,b-W_{t}\}$.
Appendix B Additional Discussions on GAIF and Adaptive GAIF
In this section, we provide additional details of our GAIF and Adaptive GAIF algorithms in Sections 2 and 3.
B.1 Explanation about utilizing feedback in SAFFRON
From another perspective, one may propose using feedback to estimate the null proportion and operate in a manner similar to SAFFRON, but in a form different from ours in Section 3.1, where the construction of $\alpha_{t}$ is
${\alpha}^{\rm SF-variant}_{t}=s_{0}\gamma_{t-D_{0+}}+(\alpha-s_{0})\gamma_{t-\tau_{1}-D_{1+}}+\alpha\sum_{j\geq 2}\gamma_{t-\tau_{j}-D_{j+}},$ (20)
here $D_{j+}=\sum_{i=\tau_{j}+1}^{t-1}{\mathbb{I}}\{\theta_{i}=1\}$. This approach constructs $\alpha_{t}$ using $\{\theta_{i}\}_{i=1}^{t-1}$ only, and can provide the same finite-sample online FDR guarantee.
Here, we explain why this form is not adopted in the proposed GAIF. In practice, we typically have $D_{j}^{+}<C_{j}^{+}$, since the set filtered by $p$-values is larger than the true non-null set. Consequently, due to the exponential decay design of $\gamma_{t}$, we have
$\gamma_{t-\tau_{j}-D_{j+}}\ll\gamma_{t-\tau_{j}-C_{j+}}.$
Although the original SAFFRON includes an adjustment factor $1-\lambda$ which reduces the $\alpha_{t}$, this term becomes negligible compared to the exponential decay of $\gamma_{t-\tau_{1}-C_{1+}}$. As a result, the original SAFFRON achieves higher power than this naive feedback-based variant, especially when $\lambda$ is large.
Our empirical results in Figure 9 further confirm this phenomenon. In the Gaussian setting described in Section 5.1 with $\mu=2$ and $\pi=0.3$, we observe that when $\lambda>0.3$, the original SAFFRON outperforms the variant in terms of power. Moreover, its highest power occurs at $\lambda=0.8$, which is relatively large. This highlights the trade-off between $\gamma_{t-\tau_{j}-C_{j+}}$ and $1-\lambda$, where where the rapidly decaying $\gamma_{t-\tau_{j}-C_{j+}}$ dominates, rendering $1-\lambda$ less influential.
In conclusion, directly using feedback to estimate the null proportion is not advisable, as it leads to a poor construction of $\alpha_{t}$. By contrast, our SF strategy leverages the $p$-values to achieve adaptive $\alpha$-wealth allocation and demonstrates superior performance.
Figure 9: The FDR and Power for SAFFRON at stopping time $600$ under different $\lambda$ value for target FDR level $\alpha=0.1$. The red lines denote the results for the variant of feedback method.
B.2 Extensions of GAIF based on $e$-values
Although feedback cannot be directly used to improve $e$-LOND [Xu and Ramdas, 2024], it can enhance $e$-LORD and $e$-SAFFRON [Zhang et al., 2025] through a feedback-driven approach analogous to the extension from GAI to GAIF. Denote $\delta_{j}=\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}$. Denote
$\operatorname{FDP}^{*}(t)=\sum_{j\leq t,j\in\mathcal{H}_{0}}\frac{\alpha_{j}}{1+R({j-1})}\leq\sum_{j\leq t}\frac{\alpha_{j}}{1+R({j-1})}=\widehat{\operatorname{FDP}}_{\text{e-LORD}}.$
Similar to the GAIF framework, we propose the estimates $\widehat{\operatorname{FDP}}_{\text{e-GAIF}}$ and $\widehat{\operatorname{FDP}}_{\text{e-Ada-GAIF}}$ as follows:
$\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\leq\widehat{\operatorname{FDP}}_{\text{e-GAIF}}=\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}+\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({t-1})}.$
$\widehat{\operatorname{FDP}}(t)_{\text{e-Ada-GAIF}}=\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}\frac{\mathbb{I}\{e_{j}\leq 1/\lambda\}}{1-\lambda}+\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({j-1})}\frac{\mathbb{I}\{e_{j}\leq 1/\lambda\}}{1-\lambda}.$
Then we require $\widehat{\operatorname{FDP}}_{\text{e-GAIF}}\leq\alpha$ or $\widehat{\operatorname{FDP}}_{\text{e-Ada-GAIF}}\leq\alpha$ when constructing the testing levels. The corresponding testing levels are as follows:
$\alpha_{t}^{\text{e-GAIF}}=\omega_{t}\left(\alpha-\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}-\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({j-1})}\right)(R({t-1})+1),$ (21)
$\alpha_{t}^{\text{e-Ada-GAIF}}=\omega_{t}\left(\alpha(1-\lambda)-\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})\mathbb{I}\{e_{j}<1/\lambda\}}{1+R({j-1})}-\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}\mathbb{I}\{e_{j}<1/\lambda\}}{1+R({j-1})}\right)(R({t-1})+1),$ (22)
where $\omega_{t}\in(0,1)$ is updated by
$\omega_{t+1}=\omega_{t}+\omega_{1}\varphi^{t-R(t)}(1-\delta_{t})-\omega_{1}\psi^{R(t)}\delta_{t}$ (23)
with a user-defined initial allocation coefficient $\omega_{1}\in(0,1)$, and user-defined parameters $\varphi>0,\psi>0$.
The Generalized Alpha-Investing with Feedback procedure based on $e$-values is summarized in Algorithm 3. The $e$-GAIF and $e$-Ada-GAIF can control online FDR validly if the null $e$-values satisfy conditionally valid.
Algorithm 3 $e$-GAIF ($e$-GAIF and $e$-Ada-GAIF)
0:  Target $\operatorname{FDR}$ level $\alpha$, parameters $\lambda,\varphi$ and $\psi\in(0,1)$, initial allocation coefficient $\omega_{1}\in(0,1)$.
1:  for $t=1,2,\dots$ do
2:   Observe $e$-value $e_{t}$
3:   Update $\alpha_{t}=\alpha_{t}^{\text{e-GAIF}}$ in Eq.(21) (or $\alpha_{t}=\alpha_{t}^{\text{e-Ada-GAIF}}$ in Eq.(22))
4:   if $e_{t}\geq 1/\alpha_{t}$ then $\delta_{t}=1$, else $\delta_{t}=0$
5:   Update $R(t)=R(t-1)+\delta_{t}$ and $\omega_{t+1}$ by (23)
6:   Obtain the revealed feedback $\theta_{t}$
7:  end for
7:  Rejection set $\{t:\delta_{t}=1\}$.
Theorem 6 (Online FDR control for e-GAIF and e-Ada-GAIF).
  • (a.)
    If the null $e$-values are conditional valid, i.e.,
    ${\mathbb{E}}\left[e_{t}\mid\mathcal{F}_{t-1}\right]\leq 1\;\text{for all}\;t\in\mathcal{H}_{0},$ (24)
    where $\mathcal{F}_{t-1}=\sigma(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t-1}})$ is the sigma field generated from past rejections and feedback, then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{e-GAIF}}(t)\leq\alpha$, then we have
    $\operatorname{FDR}(t)\leq\alpha\quad\text{for all}\quad t\in\mathbb{N}.$
  • (b.)
    Denote $C_{t}=\mathbb{I}\{e_{t}<1/\lambda\}.$ If the null $e$-values are conditional valid, i.e.,
    ${\mathbb{E}}\left[e_{t}\mid\mathcal{J}_{t-1}\right]\leq 1\;\text{for all}\;t\in\mathcal{H}_{0},$ (25)
    where $\mathcal{J}_{t-1}=\sigma(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t-1}};C_{1:t-1})$ is the enlarged filtration, then if the parameters $\{\alpha_{t}\}_{t\in\mathbb{N}}$ are selected such that $\widehat{\operatorname{FDP}}_{\operatorname{e-Ada-GAIF}}(t)\leq\alpha$ , then we have
    $\operatorname{FDR}(t)\leq\alpha\quad\text{for all}\quad t\in\mathbb{N}.$
B.3 Discussion about improving LOND/e-LOND with feedback
In this subsection, we discuss the feasibility of improving LOND or $e$-LOND through our feedback-enhanced framework.
Recall the setup of the LOND or $e$-LOND algorithm. Given a non-negative sequence $\{\gamma_{j}\}_{j=1}^{\infty}$ such that $\sum_{j=1}^{\infty}\gamma_{j}=1$, the test levels are set as
$\alpha_{t}^{\text{LOND}}=\alpha\gamma_{t}\left(\sum_{j=1}^{t-1}\delta_{j}\vee 1\right)=\alpha\gamma_{t}\left({R}(t-1)\vee 1\right).$
Improve LOND or $e$-LOND with feedback? (negative result)
Here we prove that the feedback cannot be used to improve LOND and $e$-LOND.
Denote $a_{j}=\sum_{i=1}^{j-1}\delta_{i}\theta_{i}$ and $b_{j}=\sum_{i=1}^{j-1}\delta_{i}(1-\theta_{i})$, so that ${R}({j-1})=a_{j}+b_{j}$. For LOND or e-LOND, the significance level is given by
$\alpha_{j}=\alpha\gamma_{j}(a_{j}+b_{j}+1).$
If $\theta_{1},\dots,\theta_{j-1}$ are known, we consider the modified significance level
$\tilde{\alpha}_{j}=\alpha\gamma_{j}(2w_{j}a_{j}+2(1-w_{j})b_{j}+1),$
where $w_{j}\in[0,1]$ is a weight that differentiates between true and false discoveries. Setting $w_{j}=1$ recovers $\tilde{\alpha}_{j}=\alpha_{j}$. To control $\operatorname{FDR}(t)\leq\alpha$, we require
$\sum_{j\in\mathcal{H}_{0}\cap[t]}\frac{\tilde{\alpha}_{j}}{{R}({j-1})+1}\leq\alpha.$
This implies
$2w_{j}a_{j}+2(1-w_{j})b_{j}+1\leq a_{j}+b_{j}+1.$
To improve power, we require
$\tilde{\alpha}_{j}>\alpha_{j},$
which translates to
$2w_{j}a_{j}+2(1-w_{j})b_{j}+1>a_{j}+b_{j}+1.$
However, this contradicts the FDR control condition, making the approach infeasible.
Appendix C Additional Discussions on OCTF
This section collects several supplementary discussions and extensions of OCTF that complement Section 4. These include delayed feedback (Appendix C.1), a general framework for score selection (Appendix C.2), extensions to online-updated score functions (Appendix C.3), and the discussion of OCTF under null distribution shift (Appendix C.4).
C.1 Online conformal selection with delayed Feedback
In this section, we study how to adapt OCTF to settings with delayed feedback. We focus on the case of a fixed delay $d$, where at time $t$ only the feedback $\{\theta_{j}\}_{j=1}^{t-d-1}$ is available. To maintain finite-sample mFDR control under delayed feedback, both the construction of the test levels $\alpha_{t}$ and the calibration data sets must be modified accordingly.
Specifically, note that the feedback associated with test $t$ is revealed at time $t+d+1$. To accommodate this delay, we partition the time indices into $(d+1)$ disjoint sub-streams:
$S_{1}=\{1,\,2+d,\,3+2d,\,4+3d,\ldots\},$
$S_{2}=\{2,\,3+d,\,4+2d,\,5+3d,\ldots\},$
$\vdots$
$S_{d+1}=\{1+d,\,2+2d,\,3+3d,\,4+4d,\ldots\}.$
Within each sub-stream, the feedback for a test is revealed at the next time point belonging to the same sub-stream. For each sub-stream $j\in\{1,2,\ldots,d+1\}$, we maintain an independent online multiple testing procedure based on safe GAIF, which is used to compute and update the test levels for that sub-stream. In parallel, we maintain a separate calibration data set
${\mathcal{C}}_{0t}^{j}=\{\,i<t:i\in S_{j},\ \theta_{i}=0\,\}.$
For a test observation ${\bf X}_{t}$ arriving at time $t\in S_{j}$, the test level $\alpha_{t}$ is determined by the current state of the GAIF procedure associated with sub-stream $j$. The corresponding online conformal $p$-value is defined as
$p_{t}=\frac{\sum_{i\in{\mathcal{C}}_{0t}^{j}}\mathbb{I}\{V_{i}<V_{t}\}+\xi_{t}\left(1+\sum_{i\in{\mathcal{C}}_{0t}^{j}}\mathbb{I}\{V_{i}=V_{t}\}\right)}{1+|{\mathcal{C}}_{0t}^{j}|},$ (26)
where $\xi_{t}\overset{\text{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$ is an independent random variable used for tie-breaking.
The resulting online conformal testing procedure with delayed feedback, referred to as OCTF-delay, is summarized in Algorithm 4. Finite-sample mFDR control for OCTF-delay is established in Corollary C.1.
Corollary C.1 (Finite-sample online mFDR control for OCTF-delay).
Suppose that Assumption 1 holds. Then the rejection set $\mathcal{R}$} produced by Algorithm 4 satisfies
$\operatorname{mFDR}(t)\leq\alpha\quad\text{for all }t\in\mathbb{N}.$
The proof relies on a sub-stream decomposition argument. Under fixed delay $d$, the time indices are partitioned into $(d+1)$ disjoint sub-streams. Within each sub-stream, feedback for a test is revealed before the next test in the same sub-stream, so the corresponding online conformal $p$-values are super-uniform and independent under the null conditional on the sub-stream filtration. Applying safe GAIF separately to each sub-stream therefore guarantees finite-sample mFDR control at level $\alpha$ within every sub-stream. Since false discoveries and rejections are additive across disjoint sub-streams, the overall mFDR is a weighted average of the sub-stream–wise mFDRs, and is thus controlled at level $\alpha$ at all times.
Algorithm 4 Online conformal testing with delayed feedback (OCTF-delay)
0:  Initial calibration data ${\mathcal{D}}_{\mathcal{C}}=\{({\bf X}_{i},Y_{i})\}_{i=-n+1}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, non-conformity score function $V(\cdot)$, initial wealth $s_{0}$, sequence $\{\gamma_{t}\}_{t\in\mathbb{N}}$, delay $d$.
1:  Partition $\mathbb{N}$ into $(d+1)$ sub-streams $\{S_{j}\}_{j=1}^{d+1}$.
2:  For each sub-stream $j$, initialize an independent GAIF procedure with initial wealth $s_{0}$.
3:  Initialize calibration sets ${\mathcal{C}}_{01}^{j}=\{i\leq 0:\theta_{i}=0\}$ for $j=1,\ldots,d+1$.
4:  for $t=1,2,\ldots$ do
5:   Observe test data ${\bf X}_{t}$
6:   Let $j$ be the index such that $t\in S_{j}$
7:   Compute conformity score $V_{t}$
8:   Compute conformal $p$-value $p_{t}$ using ${\mathcal{C}}_{0t}^{j}$ via (26)
9:   Compute test level $\alpha_{t}$ using the GAIF update rule of sub-stream $j$
10:   if $p_{t}\leq\alpha_{t}$ then $\delta_{t}=1$ else $\delta_{t}=0$
11:   Record decision $\delta_{t}$
12:   if feedback for some $i<t$ is revealed at time $t$:
13:      update $\theta_{i}$ and the GAIF state of the corresponding sub-stream
14:      if $\theta_{i}=0$, add $i$ to ${\mathcal{C}}_{0t}^{j}$
15:  end for
15:  Rejection set $\mathcal{R}=\{t:\delta_{t}=1\}$.
We evaluate the proposed procedure for Scenario IV with delayed feedback in Figure 10. We observe that the OCTF-delay procedures (SF-delay, LF-delay, SFS-delay, and LFS-delay) continue to achieve valid FDR control under delayed feedback, while the SF-delay and LF-delay attaining higher power compared to methods that ignore feedback.
Figure 10: Results for Scenario IV with delayed feedback: Line charts of FDR and Power with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$ after $500$ replications; The black dashed lines denote the target FDR level $\alpha=0.2$. Delay $d=5$.
C.2 A general validity framework and practical principles for optimized OCTF
We present a general validity-preserving framework for online conformal testing after score selection. The key idea is that, even when the score selection criterion is arbitrary, one can still restore the symmetry required for conformal validity by applying the same score selection rule under all permutations of the current null calibration set and test point.
At time $t$, let $\Omega_{t}$ denote the set of all permutations of ${\mathcal{C}}_{0t}\cup\{t\}$ that leave indices outside ${\mathcal{C}}_{0t}\cup\{t\}$ unchanged. For each $\sigma\in\Omega_{t}$, we apply the same score selection rule to the permuted data $({\mathcal{D}}_{t})_{\sigma}$ and obtain
$\hat{k}_{t}^{\sigma}=\arg\min_{k\in[K]}{\mathcal{M}}\bigl(({\mathcal{D}}_{t})_{\sigma};k\bigr).$
We then define the permutation-based conformal $p$-value
$p_{t}^{\rm opt}=\frac{1}{|\Omega_{t}|}\left[\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{V({\bf X}_{\sigma(t)};\hat{k}_{t}^{\sigma})<V({\bf X}_{t};\hat{k}_{t})\}+\xi_{t}\!\left(1+\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{V({\bf X}_{\sigma(t)};\hat{k}_{t}^{\sigma})=V({\bf X}_{t};\hat{k}_{t})\}\right)\right].$ (27)
Intuitively, under each permutation, the candidate test point $\sigma(t)$ is evaluated using the score function that would have been selected under that same permutation. In this way, the score selection step and the conformal ranking step are coupled symmetrically across all elements of ${\mathcal{C}}_{0t}\cup\{t\}$, restoring the exchangeability needed for conformal validity.
With this construction, arbitrary score selection criteria can be incorporated without sacrificing finite-sample error control.
Theorem 7 (Finite-sample online mFDR control based on full permutations).
Under the same assumptions as in Theorem 4, for any score selection criterion $\mathcal{M}$}, the procedure in Algorithm 2 with $\{p_{t}\}$ replaced by $\{p_{t}^{\rm opt}\}$ satisfies
$\operatorname{mFDR}(t)\leq\alpha,\qquad\forall t\in\mathbb{N}.$
Practical principles for score selection
Theorem 7 shows that online conformal testing with score selection is possible in full generality. However, this result is mainly a theoretical benchmark: the full-permutation correction can be computationally infeasible in online settings and may also degrade power when the selected score function varies substantially across permutations. This observation motivates the search for score selection criteria that are not only valid in principle, but also practically effective.
Guided by this perspective, we highlight three principles that a practically useful criterion ${\mathcal{M}}({\mathcal{D}}_{t};\cdot)$ should satisfy.
  • Power alignment. At time $t$, a natural oracle objective for power is to choose the score function $k$ maximizing
    $\Pr(p_{k,t}\leq\alpha_{t}\mid\theta_{t}=1).$
    A useful criterion should therefore be aligned with this target, in the sense that models with stronger non-null rejection ability receive better scores.
  • Permutation robustness. The criterion should not vary drastically under permutation of ${\mathcal{C}}_{0t}\cup\{t\}$. Otherwise, the selected score function $\hat{k}_{t}^{\sigma}$ may differ substantially across $\sigma$, so that $p_{t}^{\rm opt}$ compares scores computed under different score models. Such cross-model comparisons can inject extra variability, weaken the non-null signal, and reduce power. A sufficient condition avoiding this issue is that ${\mathcal{M}}({\mathcal{D}}_{t};\cdot)$ is symmetric with respect to ${\mathcal{C}}_{0t}\cup\{t\}$, in which case $\hat{k}_{t}^{\sigma}\equiv\hat{k}_{t}$ for all $\sigma$.
  • Computational efficiency. In principle, full permutation requires enumerating $|\Omega_{t}|=(|{\mathcal{C}}_{0t}|+1)!$, which is prohibitive in online applications. Criteria with sufficient symmetry, especially permutation-invariant ones, allow one to avoid explicit enumeration and reduce computation essentially to the standard conformal ranking over $|{\mathcal{C}}_{0t}|+1$ scores.
Alternative criteria
These principles clarify why some seemingly natural criteria are unattractive in practice.
First, threshold-dependent criteria involving $\alpha_{t}$ typically violate both permutation robustness and computational efficiency. Since $\alpha_{t}$ is generated by the online testing rule and depends on the past rejection trajectory, such criteria are generally not symmetric with respect to ${\mathcal{C}}_{0t}\cup\{t\}$. Moreover, a full-permutation correction would require recomputing the threshold process under each permutation, making the procedure both computationally burdensome and statistically unstable.
A second illustrative example is the criterion ${\mathcal{M}}({\mathcal{D}}_{t};k)=p_{k,t}$. At first glance, this choice appears attractive, since smaller $p$-values are more likely to lead to rejection. However, if used directly for score selection, it leads to severe validity issues: under the null, the procedure selects the minimum of $K$ candidate $p$-values, which is generally not super-uniform.
Although Theorem 7 implies that validity can still be restored through full permutation, this criterion remains practically problematic. Indeed, under this choice, for each permutation $\sigma\in\Omega_{t}$,
${\mathcal{M}}(({\mathcal{D}}_{t})_{\sigma};k)=\frac{\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf X}_{i};k)<V({\bf X}_{\sigma(t)};k)\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf X}_{i};k)=V({\bf X}_{\sigma(t)};k)\}}{1+|{\mathcal{C}}_{0t}|}.$
Hence, when $\sigma(t)\neq t$ and denote $i=\sigma(t)$, the quantity ${\mathcal{M}}(({\mathcal{D}}_{t})_{\sigma};k)$ is simply a null conformal $p$-value for $V({\bf X}_{i};k)$ based on $\{V({\bf X}_{s};k)\}_{s\in{\mathcal{C}}_{0t}\cup\{t\}}$. Score selection is then driven by null $p$-values, which contain little useful information for identifying a powerful model. As a result, $\hat{k}_{t}^{\sigma}$ may fluctuate nearly randomly across permutations, and the final $p_{t}^{\rm opt}$ in (27) is formed by ranking scores obtained under different selected score functions, possibly on different scales. This can substantially degrade power; empirically, we find that such instability may even make the procedure perform worse than random score selection.
In contrast, our EWMA criterion is designed to satisfy the above three principles: it is aligned with power through a threshold-free proxy, robust to the required permutation symmetry, and computationally efficient in online implementation. For this reason, although Theorem 7 provides a general validity result for arbitrary criteria, we recommend the EWMA criterion for practical use.
Our framework also allows other power-aligned surrogates. For example, one may consider oracle-aligned objectives of the form ${\mathbb{E}}[\phi(p_{k,t})\mid\theta_{t}=1]$, where $\phi(\cdot)$ is a deterministic nondecreasing function that emphasizes small $p$-values, such as $\phi(p)=\min(p,c)$ for some constant $c$. Such choices could move the objective closer to the tail probability $\Pr(p_{k,t}\leq\alpha_{t}\mid\theta_{t}=1)$. However, selecting $c$ appropriately, especially when $\alpha_{t}$ varies over time, would require additional tuning and theoretical analysis. In this paper, we therefore focus on $\mathcal{M}_{t}^{\mathrm{EWMA}}$ as a simple choice.
C.3 Optimized OCTF with online-updated score functions
Beyond selecting from a pool of pre-trained score functions, our framework also accommodates data-driven score functions that are updated online. The key requirement is the same as in Proposition F.1: at time $t$, the score function used to construct the conformal $p$-value must be symmetric with respect to the samples in ${\mathcal{C}}_{0t}\cup\{t\}$. Equivalently, when fitting the score model for the $t$-th test point, the current sample must be inserted as a provisional null sample and treated exchangeably with the historical null samples in ${\mathcal{C}}_{0t}$.
We focus on the binary classification setting with $Y\in\{0,1\}$. The regression case can be handled similarly by replacing the response with the binary label $\mathbb{I}\{Y\in\mathcal{A}\}$. At each time $t$, for each candidate setup $k\in[K]$, we fit a one-step-ahead predictive model $\hat{\mu}_{k}^{(t)}$ by distinguishing the non-null samples $\{({\bf X}_{i},Y_{i})\}_{i\in{\mathcal{C}}_{1t}}$ from the null samples $\{({\bf X}_{i},Y_{i})\}_{i\in{\mathcal{C}}_{0t}}\cup\{({\bf X}_{t},0)\}$. The associated score function is then given by $V(\cdot;\hat{\mu}_{k}^{(t)})$.
This construction is valid as long as the training algorithm is symmetric to its input data (or more precisely, symmetric to the treated null samples $\{({\bf X}_{i},Y_{i})\}_{i\in{\mathcal{C}}_{0t}}\cup\{({\bf X}_{t},0)\}$), which is satisfied by most batch learning procedures, such as random forests and neural networks. In contrast, order-dependent updating rules, such as standard online gradient descent, generally violate this symmetry requirement and therefore cannot be directly used here.
In this perspective, we can extend our framework of score selection. Consider the $K$ candidate setups may correspond to different score function classes or to the same score function with different hyper-parameters. When $K=1$, the procedure reduces to pure online model/score updating. When $K>1$, we update all $K$ score functions in parallel and then select the score function used for inference at time $t$. Specifically, we choose
$\hat{k}_{t}=\arg\min_{k\in[K]}{\mathcal{M}}_{t}^{\mathrm{EWMA}}({\mathcal{D}}_{t};k),$
where ${\mathcal{M}}_{t}^{\mathrm{EWMA}}({\mathcal{D}}_{t};k)$ is the same evaluation criterion as in Section 3.3. The optimized conformal $p$-value is then constructed using the selected score function based on $\hat{\mu}_{\hat{k}_{t}}^{(t)}$. Since each candidate score function is symmetric in ${\mathcal{C}}_{0t}\cup\{t\}$, and the selection criterion depends only on permutation-invariant quantities, the selected score function remains symmetric as well. Therefore, the same argument as in Corollary 4.1 yields finite-sample mFDR control.
The detailed pseudo-code is given in Algorithm 5.
Algorithm 5 Optimized OCTF with online-updated score functions
0:  Initial data ${\mathcal{D}}_{1}=\{({\bf X}_{i},Y_{i})\}_{i=-N}^{0}$, target region $\mathcal{A}$, FDR target level $\alpha\in(0,1)$, $K$ candidate model setups, evaluation criterion $\mathcal{M}$, parameter $s_{0}$, parameter sequence $\{\gamma_{t}\}$, stopping time $T$.
1:  for $t=1,\dots,T$ do
2:   Observe test covariate ${\bf X}_{t}$.
3:   for $k=1,\dots,K$ do
4:    Fit the one-step-ahead model $\hat{\mu}_{k}^{(t)}$ by distinguishing $\{({\bf X}_{i},Y_{i})\}_{i\in{\mathcal{C}}_{1t}}$ from $\{({\bf X}_{i},Y_{i})\}_{i\in{\mathcal{C}}_{0t}}\cup\{({\bf X}_{t},0)\}$.
5:    Construct the score function $V(\cdot;\hat{\mu}_{k}^{(t)})$.
6:   end for
7:   Select the predictive model by
$\hat{k}_{t}=\arg\min_{k\in[K]}{\mathcal{M}}_{t}^{\mathrm{EWMA}}({\mathcal{D}}_{t};k).$
8:   Construct the optimized conformal $p$-value
$p_{t}^{\mathrm{opt}}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\!\left\{V({\bf X}_{i};\hat{\mu}_{\hat{k}_{t}}^{(t)})\leq V({\bf X}_{t};\hat{\mu}_{\hat{k}_{t}}^{(t)})\right\}}{1+|{\mathcal{C}}_{0t}|}.$
9:   Update $\alpha_{t}=\alpha_{t}^{\mathrm{LFS}}$ in (18) or $\alpha_{t}=\alpha_{t}^{\mathrm{SFS}}$ in (19).
10:   Make a decision $\delta_{t}=\mathbb{I}\{p_{t}^{\mathrm{opt}}\leq\alpha_{t}\}$.
11:   Observe the revealed feedback $Y_{t}$ (and hence $\theta_{t}$).
12:   Update the calibration sets ${\mathcal{C}}_{0t}$ and ${\mathcal{C}}_{1t}$.
13:  end for
13:  Rejection set $\mathcal{R}_{\mathrm{opt}}=\{i:\delta_{i}=1,\ \delta_{i}\in{\bm{\delta}}^{T}\}$.
Although online updating of score functions improves data utilization and adaptivity, it is substantially more expensive than selecting from a fixed pool of pre-trained models. To preserve the required symmetry, each candidate model typically needs to be refit from scratch after adding the provisional null sample $({\bf X}_{t},0)$. For this reason, in the main text we focus on score selection among pre-trained score functions, and view online updating as a valid but computationally heavier extension.
Experimental Evaluations
We illustrate this through an additional experiment. In particular, we compare our approach with an alternative procedure that retrains the score function (predictive model) at every time step using all previously observed data. We adopt the same data-generating process as Scenario IV with sine pattern shift in Section 5.2. We set the target level $\alpha=0.2$.
Consider three strategies for constructing the predictive score function:
  • Fixed score function. A score function based on Random Forest is trained once using historical data and then kept fixed throughout the online testing process.
  • Score selection (ours). We pre-train $K$ candidate score functions (Neural network, SVM, and Random Forest) using historical data, which induce $K$ candidate score functions. During the online testing process, the algorithm dynamically selects one candidate at each time step according to the feedback-based EWMA criterion described in Section 4.3.
  • Online retraining/updating. At each time step $t$, the score function based on Random Forest is retrained using all previously observed data. To maintain the validity of conformal $p$-values, samples with feedback $\theta_{t}=0$ are incorporated into the null calibration set in the same way as other null observations.
For each strategy, we evaluate the empirical FDR and statistical power across the testing horizon. In addition, we record the total running time required to process a sequence of $T=300$ test samples, which measures the computational cost of the corresponding score construction method.
Table 2 reports the empirical FDR and power for the three strategies. All methods control the FDR at the target level. The online retraining (score updating) strategy achieves slightly higher power than the fixed model, but its improvement over the proposed score-selection approach is marginal. From a computational perspective, however, the difference is substantial. Online retraining requires fitting a predictive model at every time step, leading to a computational cost that grows linearly with the length of the testing sequence. In contrast, the proposed method only evaluates $K=3$ pre-trained candidate score functions and updates the EWMA statistics, which incurs negligible additional overhead.
Table 2: Comparison of score construction strategies for SF.
Strategy FDR Power Runtime (s)
Fixed score function 0.199 0.809 0.552
Score selection (ours) 0.198 0.846 1.680
Online retraining 0.208 0.850 150.000
C.4 OCTF under null distribution shifts
In this subsection, we examine how shifts in the null distribution affect OCTF and discuss possible remedies.
Performance evaluation under null distribution shift
To assess robustness under non-stationarity, we consider a binary classification setting similar to Scenario IV in Section 5.2, but introduce a time-varying drift in the null distribution. The initial training and calibration sets are sampled from the distribution at time $t=0$, while the test stream evolves over time:
  • Null ($Y_{t}=0$): $\mathbf{X}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\mathbf{I}_{4})$, where
    $\bm{\mu}_{t}=\Bigl(2-\Delta\frac{t}{T},0,0,0\Bigr)^{\top},$
    and $\Delta\geq 0$ controls the drift magnitude.
  • Alternative ($Y_{t}=1$): $\mathbf{X}_{t}\sim\mathcal{N}(\bm{\mu}_{1},\mathbf{I}_{4})$ with fixed mean
    $\bm{\mu}_{1}=(0,0,0,0)^{\top}.$
We set $T=1000$ and $\pi_{1}=0.5$. To adapt to distribution shift, a natural idea is to use a sliding calibration window. We therefore compare:
  • Growing window: ${\mathcal{C}}_{0t}=\{i<t:\theta_{i}=0\}$.
  • Sliding window: ${\mathcal{C}}_{0t}=\{i:t-W\leq i<t,\theta_{i}=0\}$, with $W=200$.
The growing window preserves the mutual independence of null online conformal $p$-values in Proposition 4.1, whereas the sliding window introduces local dependence among consecutive conformal $p$-values.
We consider two drift directions, $\Delta\in\{-2,2\}$. The results are shown in Figures 1112. Our main findings are:
Figure 11: Results under null distribution shift: Line charts of FDR and Power across time $t$; The black dashed lines denote the target FDR level $\alpha=0.2$. We set the drift magnitude $\Delta=-2$ so that the null shifts away from the non-null.
Figure 12: Results under null distribution shift: Line charts of FDR and Power across time $t$; The black dashed lines denote the target FDR level $\alpha=0.2$. We set the drift magnitude $\Delta=2$ so that the null drifts towards the alternative/non-null region.
  • Dual Effects of Drift: Null drift does not necessarily invalidate OCTF. When the null moves away from the alternative, the procedure remains valid but becomes conservative. When the null moves toward the alternative, FDR inflation may occur.
  • Sliding windows may hurt validity. Although sliding windows are intended to improve adaptation, in our experiments they exacerbate FDR inflation relative to the growing-window baseline.
    We attribute this to the loss of independence: the sliding window creates strong local dependence among nearby conformal $p$-values, and standard OCTF rules applied to such dependent $p$-values can inflate FDR. This suggests that simply discarding old data is not enough; the gain in adaptation may be outweighed by the loss of error-rate control. In contrast, the dependence-corrected $\mathrm{LF}_{\mathrm{dep}}$ variant restores control, albeit at a substantial power cost.
Online weighted conformal $p$-values for handling null covariate shift
These observations suggest that handling null distribution shift requires more principled corrections, such as online weighted conformal $p$-values. However, obtaining exactly valid online $p$-values becomes highly nontrivial when the null distribution changes over time.
To illustrate this difficulty, consider a stylized setting in which all observations up to time $t$ are null, but follow different distributions:
$V_{s}\sim F_{s},\quad s=1,\ldots,t,$
with $F_{s}\neq F_{s^{\prime}}$ in general. Let $\Omega_{t}$ denote the set of all permutations of $\{1,\ldots,t\}$. Given an unordered realization $\{V_{1},\ldots,V_{t}\}=\{v_{1},\ldots,v_{t}\}$, an oracle posterior over permutations can be defined as
$w(\{v_{i}\}_{i=1}^{t};\sigma)\;\propto\;\prod_{s=1}^{t}f_{s}\!\bigl(v_{\sigma(s)}\bigr),\qquad\sigma\in\Omega_{t},$
where $f_{s}$ is the null density of $F_{s}$ and the weights are normalized so that $\sum_{\sigma\in\Omega_{t}}w(\{v_{i}\}_{i=1}^{t};\sigma)=1$. Intuitively, $w(\{v_{i}\};\sigma)$ quantifies how plausible it is that the time labels $1,\ldots,t$ are assigned to the observed $\{v_{i}\}_{i=1}^{t}$ according to $\sigma$.
Using this oracle weighting, an online conformal $p$-value could be constructed by averaging the randomized rank over permutations:
$p_{t}^{w}=\frac{1}{|\Omega_{t}|}\sum_{\sigma\in\Omega_{t}}w(\{V_{i}\}_{i=1}^{t};\sigma)\left(\mathbb{I}\{V_{\sigma(t)}<V_{t}\}+\xi_{t}\mathbb{I}\{V_{\sigma(t)}=V_{t}\}\right).$
In principle, such oracle weighting can restore super-uniformity and mutual independence by integrating over the uncertainty induced by non-identical null distributions. In offline covariate-shift settings, this weighting may simplify substantially [Jin and Candès, 2025]; for example, if only the test point has a different null distribution while the calibration points remain i.i.d., the weight depends only on the identity of the test point. In the online setting, however, when all $F_{1},\dots,F_{t}$ may differ, the weight generally depends on the entire permutation, leading to factorial computational complexity. Moreover, estimating the full sequence of drifting null distributions is itself difficult, making practical implementations unstable.
An empirical investigation based on weighted conformal $p$-values
A practical workaround is to maintain a fixed calibration set drawn from a stable null distribution and compute weighted conformal $p$-values relative to this reference set, similar to the offline covariate-shift approach in Prinster et al. [2025]. Although a rigorous theoretical guarantee remains open. We conduct a real-data application to verify the performance of adopting online weighted conformal $p$-values, where oracle weights are used.
We use the airfoil dataset [Brooks,Thomas et al., 2014] from the UCI Machine Learning Repository. The task is to identify samples with high sound pressure. The dataset contains $n=1503$ observations, where the response $Y$ is the scaled sound pressure and the covariates ${\bf X}\in\mathbb{R}^{5}$ include log frequency, angle of attack, chord length, free-stream velocity, and suction-side log displacement thickness. We test
$\mathbb{H}_{0t}:Y_{t}\in(-\infty,c),$
where $c$ is the $(1-\pi_{1})$-quantile of $Y$, and $\pi_{1}\in\{0.2,\dots,0.8\}$ controls the non-null proportion.
To simulate covariate shift, we construct a shifted test set $D_{\mathrm{shift}}$ by resampling $25\%$ of the points from $D_{\mathrm{test}}$ with probabilities proportional to
$\omega(x)=\exp(x^{\top}\beta),\qquad\beta=(2,0,0,0,-3).$
Following Prinster et al. [2025], we use a fixed calibration set and compare weighted conformal $p$-values with their unweighted counterparts. Figure 13 reports the results for weighted methods (wSFS, wLFS, wSF, and wLF) and unweighted methods (SFS, LFS, SF, and LF). The weighted safe procedures wSFS and wLFS consistently maintain valid FDR control under covariate shift, whereas SF, LF, and SFS suffer from FDR inflation due to their reliance on unweighted conformal $p$-values.
Figure 13: Results for airfoil data: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ for all benchmarks. The training algorithm is random forest. The black dashed lines denote the FDR level $\alpha=0.3$. The non-null proportion is $\pi_{1}\in\{0.2,\dots,0.8\}$.
Appendix D Applications on Real-time LLM Alignment
In this section, we introduce the potential application of our proposed OCTF procedure on the task of real-time LLM alignment. For example, in medical report generation tasks, we may need to sequentially select radiology images $t\in\{1,\dots,T\}$ for which the generated reports align with expert standards. Similarly, in question-answering tasks, our goal is to identify the generated answer that best matches the true reference answer in an online fashion. Specifically, let $f:\mathcal{X}\to\mathcal{Y}$ be a pre-trained foundation model that maps a prompt to an output. A holdout set $\mathcal{D}=({\bf X}_{i},E_{i})_{i=-n}^{0}$ is available, where ${\bf X}_{i}\in\mathcal{X}$ represents an input prompt, and $E_{i}\in\mathcal{E}$ serves as a reference for assessing alignment. The alignment score function $\mathcal{A}:\mathcal{Y}\times\mathcal{E}\to\mathbb{R}$ maps the generated output $f({\bf X})$ and reference $E$ to an alignment score $A=\mathcal{A}(f({\bf X}),E)$. For example, $A$ may represent the similarity score between the machine-generated report $f({\bf X})$ and a human expert report $E$. The test data $\{{\bf X}_{t}\}_{t=1}^{T}$ arrive sequentially, forming the following online multiple hypothesis testing problem at time $t$:
$\mathbb{H}_{0t}:A_{t}\leq c\quad\text{versus}\quad\mathbb{H}_{1t}:A_{t}>c,$
where $c\in\mathbb{R}$ is a pre-specified threshold. After making a decision $\delta_{t}$ at time $t$, the corresponding human expert report is revealed either immediately or with a delay of $d$ time steps.
Our goal is to control the online $\operatorname{FDR}$:
$\operatorname{FDR}(t)={\mathbb{E}}\left[\frac{\sum_{j\leq t}\mathbb{I}\{A_{j}\leq c,\delta_{j}=1\}}{1\vee\sum_{j\leq t}\delta_{j}}\right]\leq\alpha.$ (28)
Following the conformal alignment framework Gui et al. [2024], we randomly split $\mathcal{D}$ into two subsets: a training set $\mathcal{D}_{\mathcal{T}}$ and an initial calibration set $\mathcal{D}_{{\mathcal{C}}}$. Using $\mathcal{D}_{\mathcal{T}}$, we train an alignment predictor $g({\bf X})$ to estimate the alignment score based on features of LLM outputs ${\bf X}_{t}$ and compute the predicted alignment scores $\widehat{A}_{t}=g({\bf X}_{t})$ for each $t\in[T]$. Then applying OCTF with $\mathcal{D}_{{\mathcal{C}}_{0t}}$ to select new images whose generated reports are aligned with expert standards ensures finite-sample online FDR control according to Theorem 4, where the calibration set ${\mathcal{C}}_{0t}$ is updated online as the feedback revealed. The conformal $p$-value for $t\in[T]$ is:
$p_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}{\mathbb{I}}\{\widehat{A}_{i}\geq\widehat{A}_{t}\}}{1+|{\mathcal{C}}_{0t}|}.$ (29)
The real-time LLM alignment procedure is summarized in Algorithm 6, which, similar to Theorem 4, guarantees finite-sample mFDR control under the same assumptions.
Algorithm 6 Real-time LLM conformal Alignment with feedback
0: Pre-trained foundation model $f$; alignment score function $\mathcal{A}$; reference dataset $\mathcal{D}=({\bf X}_{i},E_{i})_{i=-n}^{0}$; ; algorithm for fitting alignment predictor $\mathcal{G}$; alignment level $c$; target FDR level $\alpha$.
1: Compute the alignment score $A_{i}=\mathcal{A}(f({\bf X}_{i}),E_{i})$, $\forall i\in\mathcal{D}$.
2: Randomly split $\mathcal{D}$ into two disjoint sets: the training set $\mathcal{D}_{{\mathcal{T}}}$ and the calibration set $\mathcal{D}_{{\mathcal{C}}}$.
3: Fit the alignment score predictor with $\mathcal{D}_{{\mathcal{T}}}$: $g\leftarrow\mathcal{G}(\mathcal{D}_{{\mathcal{T}}})$.
4: Initialize ${\mathcal{C}}_{0t}=\{i\in{\mathcal{C}}:A_{i}\leq c\}$
5:for $t\in[T]$ do
6:  Observe test data ${\bf X}_{n+t}$
7:  Compute the predicted alignment score: $\hat{A}_{i}\leftarrow g({\bf X}_{i})$, $\forall i\in{\mathcal{C}}$ and $\hat{A}_{t}\leftarrow g({\bf X}_{t})$.
8:  Compute the conformal p-value $p_{t}$ according to Equation (29).
9:  Update $\alpha_{t}=\alpha_{t}^{\text{LFS}}$ in Equation (18) (or $\alpha_{t}=\alpha_{t}^{\text{SFS}}$ in (19))
10:  Obtain the revealed feedback $\theta_{t}$.
11:  Update the calibration dataset ${\mathcal{C}}_{0t}$.
12:end for
12: The selection set $\mathcal{R}=\{t:\delta_{t}=1,t\in[T]\}$.
Appendix E Additional Experiments Results
In this section, we provide additional experimental results to further demonstrate the superior performance of our proposed algorithms, with a focus on online conformal testing.
Specifically, Appendix E.1 reports additional results for a regression example (Scenario V); Appendix E.2 reports results under different training algorithms for $\hat{\mu}$; Appendix E.3 reports additional results after score selection; Appendix E.4 reports empirical mFDR results; and Appendix E.5 reports results on auxiliary calibration for score selection.
E.1 Results on synthetic data for a regression task
The corresponding data generation process for the regression task is detailed below:
  • Scenario V (Regression example): $Y=-0.5X_{1}^{2}+\exp{X_{2}}+(X_{3}+X_{4})^{2}+\varepsilon,$ with ${\bf X}\sim\mathcal{N}_{4}({\bf 0},\mathbf{I}_{4})$ and $\varepsilon\sim\mathcal{N}(0,2)$. The target region is $\mathcal{A}=[c,\infty)$, where $c$ is the $1-\pi_{1}$ quantile of $Y$.
In terms of the non-conformity score function, denote $W_{t}=\widehat{\mu}_{t}({\bf X}_{t})$, in classification settings, we set $V(W_{t})=1-W_{t}$. In regression settings, if $\mathcal{A}=[b,+\infty)$, we can use $V(W_{j})=b-W_{j}$. If $\mathcal{A}=(-\infty,a]\cup[b,+\infty)$, then we can choose $V(W_{t})=\max\{W_{t}-a,b-W_{t}\}$.
The results for the regression example (Scenario V) using a fixed training algorithm (random forest) are shown in Figure 14. The performance trends are similar to those observed in the classification case (Scenario IV).
Figure 14: Results for Scenario V: values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportions $\pi_{1}$. The black dashed line denotes the FDR level $\alpha=0.2$.
E.2 Additional experiments for different training algorithms
The results for Scenarios IV and V under different training algorithms—RF, SVM, and NN—with varying initial calibration sizes are presented in Figure 15-16. Thanks to the online updating of the calibration dataset, even a small initial calibration size does not significantly impact performance. While the choice of predictive model $\widehat{\mu}$ does affect power, all methods maintain valid FDR control. Notably, our SF and LF methods consistently outperform the baselines across all score functions, benefiting from the distribution-free and model-agnostic nature of online conformal $p$-values. The variation in performance across different algorithms further underscores the importance of careful score selection in practice.
Figure 15: Results for Scenario IV: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ vs. initial calibration size $n$ ($\pi_{1}=0.5$, $\alpha=0.2$).
Figure 16: Results for Scenario V: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ vs. initial calibration size $n$ ($\pi_{1}=0.5$, $\alpha=0.2$).
E.3 Additional experiments on score selection
The results with score selection for Scenarios IV and V are shown below Figure 17-Figure 18. In both settings, the performance gap between the Opt methods and their randomly selected counterparts is also pronounced, with Opt-SF once again achieving the highest power.
Figure 17: Results for Scenario IV: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.1$.
Figure 18: Results for Scenario V: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.1$.
E.4 Additional experiments results on empirical mFDR
To illustrate the similarity between mFDR and FDR, we present results under various settings below in Figure 19-Figure 20. We estimate mFDR by computing the ratio of the average number of false discoveries and the average total number of discoveries. Empirical mFDR closely tracks empirical FDR, and both are well controlled by our proposed methods.
Figure 19: Results for Scenario III (local dependence): Line charts of mFDR and FDR at stopping time with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$. The black dashed lines denote the target FDR level $\alpha=0.1$.
Figure 20: Results for Scenario IV and Scenario V : Line charts of mFDR and FDR at stopping time with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$ after $500$ replications; The black dashed lines denote the target FDR level $\alpha=0.2$.
E.5 Additional experiments on auxiliary calibration for score selection
In our EWMA-based model-selection procedure, we construct auxiliary non-null p-values in (13) based on ${\mathcal{C}}_{0t}\cup\{t\}$, where the current index $t$ is included. The main purpose of this design is to preserve finite-sample validity under data-adaptive score selection. In particular, including $t$ ensures the permutation symmetry of ${\mathcal{C}}_{0t}\cup\{t\}$ required by our validity argument.
From a power perspective, one may worry that, when $t$ is non-null and $t$ is small, including the current point could slightly perturb the auxiliary calibration distribution. In practice, however, this effect is limited in our setting. Our procedure starts from a historical calibration dataset $\mathcal{C}$} of size $n$, so even at early times the inclusion of a single point has only a minor impact, since the effective null calibration size is already $n+t$.
To further reduce the possible adverse effect when $t$ is non-null, we also consider a simple robustified version based on a truncated calibration set:
$\tilde{p}_{k,j}=\frac{\sum_{s\in{\mathcal{C}}_{t}^{\mathrm{trun},k}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{|{\mathcal{C}}_{t}^{\mathrm{trun},k}|},\quad j\in{\mathcal{C}}_{1t},$ (30)
where
${\mathcal{C}}_{t}^{\mathrm{trun},k}=\big({\mathcal{C}}_{0t}\cup\{t\}\big)\setminus\left\{\arg\min_{i\in{\mathcal{C}}_{0t}\cup\{t\}}V(\mathbf{X}_{i};k)\right\}.$
This operation remains permutation-invariant on ${\mathcal{C}}_{0t}\cup\{t\}$, so the symmetry needed for validity is preserved. Intuitively, when $t$ is non-null, its score $V(\mathbf{X}_{t};k)$ is more likely to be unusually small and hence removed by the truncation, reducing its impact.
To investigate the practical effect of including $t$, we compare the following four variants:
  • Opt-SFS: Our proposed feedback-enhanced approach using EWMA of past non-null auxiliary $p$-values based on calibration set ${\mathcal{C}}_{0t}\cup\{t\}$
  • OptEx-SFS: Similar to Opt-SFS but excluding $t$ from the calibration data, i.e. constructing auxiliary $p$-values based on ${\mathcal{C}}_{0t}$.
  • OptTr-SFS: Optimized selection using the truncated calibration set ${\mathcal{C}}_{t}^{\mathrm{trun},k}$ defined in Eq. (30).
  • Ran-SFS: A baseline that randomly selects a model at each time step to quantify the gain from feedback-driven selection.
We generated data streams of length $T=1000$ characterized by smoothly varying distribution shifts according to Scenario IV in Section 5.3. Conformal $p$-values were computed online using the above proposed EWMA-based selection strategies, and rejections were determined via standard SFS procedures in the main text. For each method, we tracked the empirical online FDR and empirical power across all time points across 200 replications.
Figure 21: Online FDR and power for various Opt-SFS variants versus the Ran-SFS baseline over time.
As shown in Figure 21, all three optimized methods (Opt-SFS, OptEx-SFS, and OptTr-SFS) maintain FDR control less than the nominal level ($\alpha=0.1$) while consistently achieving higher power than the random baseline.
A comparison between Opt-SFS and OptEx-SFS reveals that at early time points ($t$ small), differences in power among these strategies are relatively minor. As more observations accumulate, the differences among the three Opt methods diminish. These results empirically verify that the inclusion of $t$ has negligible influence on the power performance.
Appendix F Technical Details
F.1 Auxiliary lemmas and proof of lemmas
Our proof of Theorem 1 and Theorem 2 relies on the following lemmas, which are modified versions of Lemma 1 in Ramdas et al. [2017] and Lemma 1 in Ramdas et al. [2018], respectively. The modifications arise because, in our setting, the feedback information $\{\theta_{j}\}_{j\in\mathcal{I}_{t}}$ is available at time $t$, and the test levels $\alpha_{t}$ in GAIF and Adaptive GAIF procedures depend on both past rejections $\delta_{j}$ and feedback $\{\theta_{j}\}_{j\in\mathcal{I}_{t}}$.
For simplicity, denote $\delta_{1:t}=(\delta_{1},\dots,\delta_{t})$. Given a sequence $p_{1},p_{2},\dots$ of independent $p$-values (i.e., the null $p$-values are independent of all other $p$-values), we define a filtration via the sigma-field of decisions and feedback $\mathcal{F}_{t-1}=\sigma(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})$, where $\delta_{t}=\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}$ for some coordinate-wise non-decreasing function $f_{t}:\{0,1\}^{t-1+|\mathcal{I}_{t}|}\to\mathbb{R}$. With this set-up, we have the following guarantees in Lemma 1:
Lemma 1.
Let $g:\{0,1\}^{T}\to\mathbb{R}$ be any coordinate-wise non-decreasing function such that $g(\bm{x})>0$ for any vector $\bm{x}\neq(0,\dots,0)$. Then for any index $t\leq T$ such that $H_{t}\in\mathcal{H}_{0}$, we have
${\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}}{g(\delta_{1:T})}\mid\mathcal{F}_{t-1}\right]\leq{\mathbb{E}}\left[\frac{f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})}{g(\delta_{1:T})}\mid{\mathcal{F}}_{t-1}\right].$
Proof.
Denote ${p}_{1:T}=(p_{1},\dots,p_{T})$, and $\tilde{{p}}_{1:T}^{-t}=(\tilde{p}_{1},\dots,\tilde{p}_{T})$, where
$\tilde{p}_{i}=\begin{cases}0\quad\text{if}\;i=t\\ p_{i}\quad\text{if}\;i\neq t,\end{cases}$
i.e., the $t$-th component of $\tilde{{p}}_{1:T}^{-t}$ is zero and the other components of $\tilde{{p}}_{1:T}^{-t}$ equals to that of ${p}_{1:T}$. For all $i$, define $\tilde{\delta}_{i}=\mathbb{I}\{\tilde{{p}}_{i}\leq f_{i}(\tilde{\delta}_{1:i-1},\{\theta_{j}\}_{j\in\mathcal{I}_{i}})\}$. And let the decision vectors using ${p}_{1:T}$ and $\tilde{{p}}_{1:T}^{-t}$ be ${\delta}_{1:T}=(\delta_{1},\dots,\delta_{T})$ and $\tilde{{\delta}}_{1:T}^{-t}=(\tilde{\delta}_{1},\dots,\tilde{\delta}_{T})$. By construction, we have
[…]
from which we obtain $f_{i}(\delta_{1:i-1};\{\theta_{j}\}_{j\in\mathcal{I}_{i}})=f_{i}(\tilde{\delta}_{1:i-1};\{\theta_{j}\}_{j\in\mathcal{I}_{i}})$ for all $i\leq t$. Also, by noting that $\tilde{\delta}_{t}=1$ by construction and by definition of $g(\cdot)$, we have $g(\tilde{{\delta}}_{1:T}^{-t})>0$. Hence, on the event $\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}$, we have $\delta_{t}=\tilde{\delta}_{t}=1$ and hence also ${\delta}_{1:T}=\tilde{{\delta}}_{1:T}^{-t}$, which allows us to conclude that
$\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}}{g({\delta}_{1:T})}=\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}}{g(\tilde{{\delta}}_{1:T}^{-t})}.$
Since $\tilde{{\delta}}_{1:T}^{-t}$ is independent of $p_{t}$, we take conditional expectations to obtain
${\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}}{g({\delta}_{1:T})}\mid{\mathcal{F}}_{t-1}\right]$ $={\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}}{g(\tilde{{\delta}}_{1:T}^{-t})}\mid{\mathcal{F}}_{t-1}\right]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}{\mathbb{E}}\left[\frac{f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})}{g(\tilde{{\delta}}_{1:T}^{-t})}\mid{\mathcal{F}}_{t-1}\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}{\mathbb{E}}\left[\frac{f_{t}(\delta_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})}{g({{\delta}}_{1:T})}\mid{\mathcal{F}}_{t-1}\right],$
where inequality (i) follows by taking expectation only with respect to $p_{t}$ by the conditional super-uniformity property (6); and inequality (ii) follows because $g({\delta}_{1:T})\leq g(\tilde{{\delta}}_{1:T}^{-t})$ since $\delta_{i}\leq\tilde{\delta}_{i}$ for all $i$ by monotonicity of the online FDR rule. This concludes the proof of the Lemma 1.
Denote $C_{1:t}:=(C_{1},\dots,C_{t})$, where $C_{t}=\mathbb{I}\{p_{t}\leq\lambda\}$, with a fixed constant $\lambda\in(0,1)$. Furthermore, assume $\alpha_{t}=h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})$, for some coordinate-wise non-decreasing function $h_{t}:\{0,1\}^{2(t-1)+|\mathcal{I}_{t}|}\to[0,\lambda]$. Define the sigma-fields $\mathcal{J}_{t-1}=\sigma(\delta_{1:t-1};C_{1:t-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t}})$. For independent $p$-values, we have the following guarantees.
Lemma 2.
Let $g:\{0,1\}^{T}\to\mathbb{R}$ be any coordinate-wise non-decreasing function. Then, for any index $t\leq T$ such that $H_{t}\in\mathcal{H}_{0}$, we have
${\mathbb{E}}\left[\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)g(\delta_{1:T})}\mid\mathcal{J}_{t-1}\right]$
$\geq{\mathbb{E}}\left[\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})}{g(\delta_{1:T})}\mid\mathcal{J}_{t-1}\right]$
$\geq{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\}}{g(\delta_{1:T})}\mid\mathcal{J}_{t-1}\right]$
Proof.
The second inequality is a consequence of Lemma 1, so we only prove the first inequality. Denote ${p}_{1:T}=(p_{1},\dots,p_{T})$, and $\bar{{p}}^{t\to 1}_{1:T}=(\bar{p}_{1},\dots,\bar{p}_{T})$, where
$\bar{p}_{i}=\begin{cases}1\quad\text{if}\;i=t\\ p_{i}\quad\text{if}\;i\neq t,\end{cases}$
i.e., the $t$-th component of $\bar{{p}}^{t\to 1}_{1:T}$ is set to one and the other components of $\bar{{p}}^{t\to 1}_{1:T}$ equals to that of ${p}_{1:T}$. Define $\bar{C}_{i}=\mathbb{I}\{\bar{p}_{i}\leq\lambda\}$ and $\bar{\delta}_{i}=\mathbb{I}\{\bar{p}_{i}\leq h_{i}(\bar{\delta}_{1:i-1},\bar{C}_{1:i-1},\{\theta_{j}\}_{j\in\mathcal{I}_{i}})\}$ respectively. Let $\delta_{1:T}=(\delta_{1},\dots,\delta_{T})$ and $\bar{\delta}^{t\to 1}_{1:T}=(\bar{\delta}_{1},\dots,\bar{\delta}_{T})$ denote the decision vectors using $p_{1:T}$ and $\bar{p}_{1:T}$, respectively. Similarly, $C_{1:T}=(C_{1},\dots,C_{T})$ and $\bar{C}_{1:T}=(\bar{C}_{1},\dots,\bar{C}_{T})$. By construction, we have
  1. 1.
    $\bar{\delta}_{i}=\delta_{i}$ and $\bar{C}_{i}=C_{i}$ for all $i\leq t$, hence $h_{i}(\delta_{1:i-1},C_{1:i-1},\{\theta_{j}\}_{j\in\mathcal{I}_{i}})=h_{i}(\bar{\delta}_{1:i-1},\bar{C}_{1:i-1},\{\theta_{j}\}_{j\in\mathcal{I}_{i}})$ for all $i\leq t$.
  2. 2.
    $\bar{\delta}_{t}=\bar{C}_{t}=0$, and hence $\bar{\delta}_{i}\leq\delta_{i}$ for all $i\geq t$, due to monotonicity of the function $h_{i}$.
Hence, on the event $\{p_{t}>\lambda\}$, we have $\delta_{t}=\bar{\delta}_{t}=0$ and $C_{t}=\bar{C}_{t}=0$, and hence also $\delta_{1:T}=\bar{\delta}_{1:T}^{t\to 1}$. Therefore, we obtain
$\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)g(\delta_{1:T})}=\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)g(\bar{\delta}_{1:T})}.$
Since $\bar{\delta}_{1:T}^{t\to 1}$ is independent of $p_{t}$, we may take conditional expectations to obtain:
${\mathbb{E}}\left[\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)g(\delta_{1:T})}\mid\mathcal{J}_{t-1}\right]$
$={\mathbb{E}}\left[\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)g(\bar{\delta}^{t\to 1}_{1:T})}\mid\mathcal{J}_{t-1}\right]$
$\stackrel{{\scriptstyle(i)}}{{\geq}}{\mathbb{E}}\left[\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})}{g(\bar{\delta}^{t\to 1}_{1:T})}\mid\mathcal{J}_{t-1}\right]$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}{\mathbb{E}}\left[\frac{h_{t}(\delta_{1:t-1},C_{1:t-1},\{\theta_{j}\}_{j\in\mathcal{I}_{t}})}{g(\delta_{1:T})}\mid\mathcal{J}_{t-1}\right],$
where inequality (i) follows by taking an expectation only with respect to $p_{t}$ by invoking the conditional super-uniformity property (10); and inequality (ii) follows because $g(\delta_{1:T})\geq g(\bar{\delta}^{t\to 1}_{1:T})$ since $\delta_{i}\geq\bar{\delta}_{i}$ for all $i$ by monotonicity of the online FDR rule. This concludes the proof of Lemma 2. ∎
Lemma 3.
Let $\mathcal{F}$ be a sigma-field and let $\mathcal{E},\mathcal{A}\in\mathcal{F}$ be the events and $\mathcal{E}\perp\mathcal{A}$. Let $Z$ be a random variable and $b\in\mathbb{R}$. Assume that $\Pr(\mathcal{E})\geq 1-\delta^{\prime}$ and on $\mathcal{E}$, $\Pr(Z\leq b\mid\mathcal{F})\geq 1-\delta$. Then
$\Pr(Z\leq b\mid\mathcal{A})\geq 1-\delta-\delta^{\prime}.$
Proof.
Since $\mathcal{E}\in\mathcal{F}$,
$\Pr(Z\leq b\mid\mathcal{A})\ \geq\ \Pr(Z\leq b,\mathcal{E}\mid\mathcal{A})$
$=$ ${\mathbb{E}}\!\left[\mathbb{I}_{\mathcal{E}}\Pr(Z\leq b\mid\mathcal{F})\mid\mathcal{A}\right]\geq(1-\delta)\Pr(\mathcal{E}\mid\mathcal{A})\geq 1-\delta-\delta^{\prime}.$
The last inequality comes from the independence of $\mathcal{E}$ and $\mathcal{A}$ such that $\Pr(\mathcal{E}\mid\mathcal{A})=\Pr(\mathcal{E})\geq 1-\delta^{\prime}$. ∎
F.2 Proof of Theorem 1
Proof.
Proof of Theorem 1 (a). (mFDR control of GAIF procedures.) For the GAIF procedures,
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid{\mathcal{F}}_{j-1}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where equality (i) is derived by conditioning on $\mathcal{F}_{j-1}$ and applying the law of iterated expectations. Inequality (ii) is a consequence of the conditional super-uniformity property stated in (6). Subsequent inequalities (iii) follows from the fact that $1-\theta_{j}\leq 1$ for all $j\in\bar{\mathcal{I}}_{t}$ and (iv) follows from the definition of GAIF procedures. Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for all GAIF procedures.
Proof of Theorem 1 (b). (FDR control of GAIF procedures procedures.) We first prove online FDR control for GAIF procedures. Under the independence and the monotonicity assumptions, we have
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$=$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}}\mid{\mathcal{F}}_{j-1}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[\frac{\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}(1-\theta_{j})\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ ${\mathbb{E}}\left[\frac{\sum_{j\in\mathcal{I}_{t}}(1-\theta_{j})\,\alpha_{j}+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}}{1\vee\sum_{j=1}^{t}\delta_{j}}\right]$ (31)
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\operatorname{GAIF}}(t)\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ $\alpha,$
where the equality (i) follows form the law of iterated expectations by conditioning on $\mathcal{F}_{j-1}$ and inequality (ii) results from Lemma 1 by setting $g(\delta_{1:t})=(1\vee\sum_{j\leq t}\delta_{j})$, and the inequality (iii) holds because $1-\theta_{j}\leq 1$ for all $j\in\bar{\mathcal{I}}_{t}$, the inequality (v) follows from definition of GAIF, which completes the proof of FDR control for monotone GAIF procedures.
F.3 Proof of Theorem 2
Proof.
Proof of Theorem 2 (a). (mFDR control of Adaptive GAIF procedures.)
Now we prove mFDR control for the Adaptive GAIF procedures. The conditional super-uniformity in(10) can be rephrased as:
${\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}>\alpha_{t}\}}{1-\alpha_{t}}\mid\mathcal{J}_{t-1}\right]\geq 1\geq{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq\alpha_{t}\}}{\alpha_{t}}\mid\mathcal{J}_{t-1}\right].$ (32)
Note that for any time $t\in[T]$, we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$ (33)
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid\mathcal{J}_{t-1}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}\right]$
$\stackrel{{\scriptstyle(v)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where the equality (i) follows from the law of iterated expectations by conditioning on $\mathcal{J}_{j-1}$ and inequality (ii) applying the conditional super-uniformity property in (10), and the inequality (iii) also follows by the law of iterated expectations by conditioning on $\mathcal{J}_{j-1}$ and then applying the conditional super-uniformity property in (32), and the inequality (iv) holds since the fact that $1-\theta_{j}\leq 1$ for all $j\in\bar{\mathcal{I}}_{t}$, and the inequality (v) follows from the construction such that $\widehat{\operatorname{FDP}}_{\text{Adaptive-GAIF}}\leq\alpha$. Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for Adaptive GAIF procedures.
Proof of Theorem 2 (b). (FDR control of Adaptive GAIF procedures.)
We next prove the online FDR control for Adaptive GAIF procedures. Under the independence and the monotonicity assumptions, we have
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$=$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}}\mid\mathcal{J}_{j-1}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[\frac{\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ $\sum_{j\leq t,j\in\mathcal{H}_{0}}{\mathbb{E}}\left[\frac{\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\cdot\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}\right]$
$=$ ${\mathbb{E}}\left[\frac{\sum_{j=1}^{t}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ ${\mathbb{E}}\left[\frac{\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\operatorname{Ada-GAIF}}(t)\right]$
$\stackrel{{\scriptstyle(v)}}{{\leq}}$ $\alpha,$
where (i) and (ii) follows form the law of iterated expectations by conditioning on $\mathcal{J}_{j-1}$ and applying Lemma 2 by setting $g(\delta_{1:t})=(1\vee\sum_{j\leq t}\delta_{j})$, the inequality (iii) also follows from the law of iterated expectations by conditioning on $\mathcal{J}_{j-1}$ and applying Lemma 2 by setting $g(\delta_{1:t})=(1\vee\sum_{j\leq t}\delta_{j})$, the inequality (iv) holds since $1-\theta_{j}\leq 1$ for all $j\in\bar{\mathcal{I}}_{t}$, and the inequality (v) follows from the definition of Adaptive GAIF, which completes the proof of online FDR control for monotone Adaptive GAIF.
F.4 Proof of Theorem 3
Proof.
Define locally conditional super-uniformity as follows: if the null hypothesis $H_{t}$ is true, then for all $\alpha_{t}\in[0,1]$,
$\Pr\left(p_{t}\leq\alpha_{t}\mid{\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}\right)\leq\alpha_{t},$
where ${\mathcal{X}}^{t}:=\{t-L_{t},\dots,t-1\}$, and ${\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}:=\sigma(\delta_{1:t-L_{t}-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t-L_{t}-1},1\leq j\leq t-L_{t}-1})$.
This condition is immediately true by local dependence [Zrnic et al., 2021]. Note that for any time $t\in[T]$, we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$ (34)
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid{\mathcal{F}}_{\text{dep}}^{-{\mathcal{X}}^{t}}\right]\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}[(1-\theta_{j})\alpha_{j}]={\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})+\sum_{j\in\bar{\mathcal{I}}_{t}}\alpha_{j}\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\right],\$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where the first inequality follows from the law of iterated expectations by conditioning on $\mathcal{F}_{\text{dep}}^{-{\mathcal{X}}^{t}}$ and then applying the conditional super-uniformity property and by noticing that the measurability of $\alpha_{j}$ with respect to $\mathcal{F}_{\text{dep}}^{-{\mathcal{X}}^{t}}$, and the second inequality follows by the fact that $1-\theta_{t}\leq 1$, and the third inequality follows from the construction such that $\widehat{\operatorname{FDP}}_{\text{GAIF}_{\text{dep}}}\leq\alpha$. Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for $\text{GAIF}_{\text{dep}}$.
Define $\mathcal{J}_{\text{dep}}^{-{\mathcal{X}}^{t}}:=\sigma(\delta_{1:t-L_{t}-1};C_{1:t-L_{t}-1};\{\theta_{j}\}_{j\in\mathcal{I}_{t-L_{t}-1},1\leq j\leq t-L_{t}-1})$. Under local dependence, we have the conditional super-uniformity for null $p$-values:
${\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}>\alpha_{t}\}}{1-\alpha_{t}}\mid\mathcal{J}_{\text{dep}}^{-{\mathcal{X}}_{t}}\right]\geq 1\geq{\mathbb{E}}\left[\frac{\mathbb{I}\{p_{t}\leq\alpha_{t}\}}{\alpha_{t}}\mid\mathcal{J}_{\text{dep}}^{-{\mathcal{X}}_{t}}\right].$ (35)
The mFDR control can be obtained for $\text{Ada-GAIF}_{\text{dep}}$ as follows.
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]=\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$ (36)
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\sum_{j\leq t}{\mathbb{E}}\left[{\mathbb{E}}\left[(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid\mathcal{J}_{\text{dep}}^{-{\mathcal{X}}^{t}}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ $\sum_{j\leq t}{\mathbb{E}}[(1-\theta_{j})\alpha_{j}]={\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\in\mathcal{I}_{t}}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\sum_{j\in\mathcal{I}_{t}}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j<t-L_{t}}\frac{\alpha_{j}\mathbb{I}\{p_{j}>\lambda\}}{1-\lambda}\mathcal{W}_{j}+\sum_{t-L_{t}\leq j\leq t}\frac{\alpha_{j}}{1-\lambda}\mathcal{W}_{j}\right]$
$\stackrel{{\scriptstyle(v)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t,j\notin\{t-L_{t},\dots,t-1\}}\delta_{j}\right]\$
$\stackrel{{\scriptstyle(vi)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right],$
where $\mathcal{W}_{j}=(1-\theta_{j})\mathbb{I}\{j\in\mathcal{I}_{t}\}+\mathbb{I}\{j\in\bar{\mathcal{I}}_{t}\}$ and the equality (i) follows from the law of iterated expectations by conditioning on $\mathcal{J}_{\rm dep}^{-{\mathcal{X}}^{t}}$, and the inequality (ii) applying the conditional super-uniformity, and the inequality (iii) also follows by the law of iterated expectations by conditioning on $\mathcal{J}_{\rm dep}^{-{\mathcal{X}}^{t}}$ and then applying the conditional super-uniformity. The inequality (iv) follows from the fact that $\mathbb{I}\{p_{j}>\lambda\}\leq 1$ for all $j\in\{t-L_{t},\dots,t\}$ and $1-\theta_{t}\leq 1$. The inequality (v) follows from the construction such that $\widehat{\operatorname{FDP}}_{\text{Ada-GAIF}_{\text{dep}}}\leq\alpha$, which concludes the proof.
F.5 Proof of Proposition 4.1
We first restate Proposition 4.1 as follows and finish the proof. The proof of Proposition F.1 essentially follows the argument of Theorem 8.2 in Angelopoulos et al. [2024], with an extension to our online conformal testing setting.
Proposition F.1 (Validity and Mutual Independence of Online Conformal $p$-values under Exchangeability and Symmetric Scores).
Suppose at each time $t$, the score function $V(\cdot;{\mathcal{D}}_{t})$ is constructed through the current data ${\mathcal{D}}_{t}=\big(({\bf X}_{i},Y_{i}):-n\leq i\leq t\big)$. The p-value of each time $t$ is constructed as
$p_{t}=\frac{\sum_{i\in{{\mathcal{C}}}_{0t}}{\mathbb{I}}\{V({\bf X}_{i};{\mathcal{D}}_{t})<V({\bf X}_{t};{\mathcal{D}}_{t})\}+\xi_{t}\cdot(1+\sum_{i\in{\mathcal{C}}_{0t}}\mathbb{I}\{V({\bf X}_{i};{\mathcal{D}}_{t})=V({\bf X}_{t};{\mathcal{D}}_{t})\})}{1+|{\mathcal{C}}_{0t}|}.$ (37)
Suppose Assumption 1 holds and the score function $V(\cdot;{\mathcal{D}}_{t})$ is symmetric to $\{({\bf X}_{i},Y_{i}):-n\leq i\leq t,\theta_{i}=0\}$. Then under the null hypothesis,
  1. 1.
    Each null $p$-value $p_{t}$ with $\theta_{t}=0$ is marginally uniformly distributed on $[0,1]$.
  2. 2.
    The null $p$-values $\{p_{t}:t\in\mathbb{N},\theta_{t}=0\}$ are mutually independent.
Proof.
We adapt the standard argument for the validity of conformal $p$-values to our online setting with exchangeable data. For each time $t$, define
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Phi_{t}=\Biggl(\{({\bf X}_{i},Y_{i}):i\in{\mathcal{C}}_{0t}\cup\{t\}\},\,(\theta_{i}:i<t),\,\bigl(({\bf X}_{i},Y_{i}):i\in{\mathcal{C}}_{1t}\bigr)\Biggr),$
where $\Phi_{t}$ contains the unordered set of conformity scores for indices in ${\mathcal{C}}_{0t}\cup\{t\}$, the true state $\theta_{i}$, and the data corresponding to indices not used in the calibration set (denoted here by ${\mathcal{C}}_{1t}$). To prove the mutual independence, consider any time indices $t\leq T$ and let $x_{t},x_{t+1},\dots,x_{T}\in[0,1]$ be arbitrary. For ease of notation, we first consider the case in which all p-values $\{p_{s}\}_{s=t}^{T}$ are associated with null hypotheses. The argument extends directly to the more general setting in which some non-null hypotheses appear between them.
Since $\{p_{k}\}_{k=t+1}^{T}$ are determined by $\Phi_{t}\cup\{({\bf X}_{i},Y_{i})\}_{i=t+1}^{T}$ and $p_{t}$ is independent of $\{({\bf X}_{i},Y_{i})\}_{i=t+1}^{T}$, it holds that
$\Pr\bigl(p_{t}\leq x_{t},\,p_{t+1}\leq x_{t+1},\,\dots,\,p_{T}\leq x_{T}\bigr)$
$=$ $\;\mathbb{E}\Bigl[\mathbb{I}\{p_{t}\leq x_{t}\}\mathbb{I}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbb{I}\{p_{T}\leq x_{T}\}\Bigr]$
$=$ $\;\mathbb{E}\Biggl[\mathbb{E}\Bigl[\mathbb{I}\{p_{t}\leq x_{t}\}\,\Big|\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Phi_{t}\cup\{({\bf X}_{i},Y_{i})\}_{i=t+1}^{T}}\Bigr]\cdot\mathbb{I}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbb{I}\{p_{T}\leq x_{T}\}\Biggr]$
$=$ $\;\mathbb{E}\Biggl[\mathbb{E}\Bigl[\mathbb{I}\{p_{t}\leq x_{t}\}\,\Big|\,\Phi_{t}\Bigr]\cdot\mathbb{I}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbb{I}\{p_{T}\leq x_{T}\}\Biggr].$
The key observation is that, by exchangeability, the conditional distribution of $p_{t}$ given $\Phi_{t}$ is uniform on $[0,1]$, so that
$\mathbb{E}\Bigl[\mathbb{I}\{p_{t}\leq x_{t}\}\,\Big|\,\Phi_{t}\Bigr]=x_{t}.$
Thus,
$\Pr\bigl(p_{t}\leq x_{t},\,p_{t+1}\leq x_{t+1},\,\dots,\,p_{T}\leq x_{T}\bigr)=x_{t}\,\mathbb{E}\Bigl[\mathbb{I}\{p_{t+1}\leq x_{t+1}\}\cdots\mathbb{I}\{p_{T}\leq x_{T}\}\Bigr].$
Next, we note the following two key facts:
(a) Uniformity of single p-value $p_{t}$:
Fix ant $t\in[T]$. Let $\Omega_{t}$ be the sets of all permutations of ${\mathcal{C}}_{0}\cup[T]$ that fixes indices outside of ${\mathcal{C}}_{0t}\cup\{t\}$. Note that given $\Phi_{t}$, the only randomness for $p_{t}$ is the order of $\{({\bf X}_{i},Y_{i}):i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ and $\xi_{t}$.
For any $\sigma\in\Omega_{t}$ and given $\{{\bf X}_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}=\{{\bf x}_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ as the realizations, define
$p_{t}(({\mathcal{D}}_{T})_{\sigma})$
$=$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{\sigma(i)};({\mathcal{D}}_{t})_{\sigma})<V({\bf x}_{\sigma(t)};({\mathcal{D}}_{t})_{\sigma})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{\sigma(i)};({\mathcal{D}}_{t})_{\sigma})=V({\bf x}_{\sigma(t)};({\mathcal{D}}_{t})_{\sigma})\}}{1+|{\mathcal{C}}_{0t}|}$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{\sigma(i)};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{\sigma(i)};{\mathcal{D}}_{t})=V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}$
$\stackrel{{\scriptstyle(ii)}}{{=}}$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})=V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}.$
Equality (i) holds since $V(\cdot;\sigma(D_{t}))=V(\cdot;D_{t})$ from the symmetry of score function $V$. Equality (ii) is true by the property that $\sigma$ only permutes indices in ${\mathcal{C}}_{0t}\cup\{t\}$ such that $\{V({\bf X}_{i};D_{t}):i\in{\mathcal{C}}_{0t}\}=\{V({\bf X}_{\sigma(i)};D_{t}):i\in{\mathcal{C}}_{0t}\}$.
Denote $Q_{\alpha}(S_{i}:i\in\mathcal{I})$ as the $\alpha$-th quantile of the set $\{S_{i}\}_{i\in{\mathcal{I}}}$. Let $\{X_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}=\{x_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ as a set of realizations and
$q=Q_{\alpha}(V(x_{i};{\mathcal{D}}_{t}):i\in{\mathcal{C}}_{0t}\cup\{t\}).$
For any $\sigma\in\Omega_{t}$ such that $V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})>q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma})\geq\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}{\geq}\frac{N_{=}+N_{-}}{1+|{\mathcal{C}}_{0t}|}\geq\alpha;$
Here
$N_{-}=\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{(V(x_{i};{\mathcal{D}}_{t})<q\},\quad N_{=}=\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{(V(x_{i};{\mathcal{D}}_{t})=q\}.$
The last inequality holds by the property of quantile function such that $N_{=}+N_{-}\geq 1+\alpha(|{\mathcal{C}}_{0t}+1|)$.
For any $\sigma\in\Omega_{t}$ such that $V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})<q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})\leq V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}\leq\frac{N_{-}}{1+|{\mathcal{C}}_{0t}|}\leq\alpha,$
which is from the property that $N_{-}\leq\alpha(|{\mathcal{C}}_{0t}+1|)$.
And for any $\sigma\in\Omega_{t}$ such that $V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma})=$ $\frac{\sum_{i\in{{\mathcal{C}}}_{0t}\cup\{t\}}{\mathbb{I}}\{V({\bf x}_{i};{\mathcal{D}}_{t})<V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}+\xi_{t}\cdot\sum_{i\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})=V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})\}}{1+|{\mathcal{C}}_{0t}|}$
$\leq$ $\frac{N_{-}+\xi_{t}N_{=}}{1+|{\mathcal{C}}_{0t}|}.$
Hence,
$\Pr(p_{t}\leq\alpha\mid\Phi_{t})$ $=\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}\sum_{\sigma\in\Omega_{t}}\Pr(p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq\alpha)$
$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}\sum_{\sigma\in\Omega_{t}}\left(\mathbb{I}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})<q\}+\mathbb{I}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q\}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}\right)$
$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{|{\mathcal{C}}_{0t}|+1}\sum_{i\in{\mathcal{C}}_{0t}\cup\{j\}}\left(\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})<q\}+\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})=q\}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}\right)$
$\stackrel{{\scriptstyle(iii)}}{{=}}\frac{1}{|{\mathcal{C}}_{0t}|+1}\left(N_{-}+N_{=}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}\right)=\alpha,$
where equality (i) holds since $\xi_{t}$ is uniformly distributed and independent of everything else, making
${\mathbb{E}}[\mathbb{I}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q,p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq\alpha\}\mid({\mathcal{D}}_{T})_{\sigma}]=\mathbb{I}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q\}\frac{\alpha(|{\mathcal{C}}_{0t}|+1)-N_{-}}{N_{=}}.$
Equality (ii) is from the fact that $\sum_{\sigma\in\Omega_{t},\sigma(t)=i}\mathbb{I}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})<q\}=|{\mathcal{C}}_{0t}|\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})<q\}$ and $\sum_{\sigma\in\Omega_{t},\sigma(t)=i}\mathbb{I}\{V({\bf x}_{\sigma(t)};{\mathcal{D}}_{t})=q\}=|{\mathcal{C}}_{0t}|\mathbb{I}\{V({\bf x}_{i};{\mathcal{D}}_{t})=q\}$. And equality (iii) is direct by the definition of $N_{-}$ and $N_{=}$. Marginalizing over the $\Phi_{t}$ implies $\Pr(p_{t}\leq\alpha)=\alpha$ for all $t\in[T]$.
(b) Independence of future $p$-values from $\Phi_{t}$:
Define the data set ${\mathcal{D}}_{T}=(({\bf X}_{-n+1},Y_{-n+1}),\dots,({\bf X}_{T},Y_{T}))$. By exchangeability of the data, it holds that ${\mathcal{D}}_{T}\overset{d}{=}({\mathcal{D}}_{T})_{\sigma}$, where $({\mathcal{D}}_{T})_{\sigma}$ is obtained from ${\mathcal{D}}_{T}$ by permuting the data points according to $\sigma$. A key observation is that for any $\sigma\in\Omega_{t}$, defining $\sigma$ as above we have
$p_{t^{\prime}}({\mathcal{D}}_{T})=p_{t^{\prime}}(({\mathcal{D}}_{T})_{\sigma})$
for all $t^{\prime}\in\{t+1,\dots,T\}$. Intuitively, this indicates that permuting the data according to $\sigma$ does not change $p$-values after time $t$.
This is because $p_{t^{\prime}}(({\mathcal{D}}_{T})_{\sigma})$ is
$\frac{\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbb{I}\{V({\bf X}_{\sigma(i)};\sigma({\mathcal{D}}_{t^{\prime}}))<V({\bf X}_{\sigma(t^{\prime})};\sigma({\mathcal{D}}_{t^{\prime}}))+\xi_{t^{\prime}}(1+\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbb{I}\{V({\bf X}_{\sigma(i)};\sigma({\mathcal{D}}_{t^{\prime}}))=V({\bf X}_{\sigma(t^{\prime})};\sigma({\mathcal{D}}_{t^{\prime}}))\})}{1+|{\mathcal{C}}_{0t^{\prime}}|}$
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\frac{\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbb{I}\{V({\bf X}_{\sigma(i)};{\mathcal{D}}_{t^{\prime}})<V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\}+\xi_{t^{\prime}}(1+\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbb{I}\{V({\bf X}_{\sigma(i)};{\mathcal{D}}_{t^{\prime}})=V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\})}{1+|{\mathcal{C}}_{0t^{\prime}}|}\;$
$\stackrel{{\scriptstyle(ii)}}{{=}}$ $\frac{\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbb{I}\{V({\bf X}_{i};{\mathcal{D}}_{t^{\prime}})<V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\}+\xi_{t^{\prime}}(1+\sum_{i\in{\mathcal{C}}_{0t^{\prime}}}\mathbb{I}\{V({\bf X}_{i};{\mathcal{D}}_{t^{\prime}})=V({\bf X}_{t^{\prime}};{\mathcal{D}}_{t^{\prime}})\})}{1+|{\mathcal{C}}_{0t^{\prime}}|}\;$
$=$ $p_{t^{\prime}}({\mathcal{D}}_{T}),$
where equality (i) holds since $\sigma(t^{\prime})=t^{\prime}$ by definition and $V(\cdot;\sigma(D_{t^{\prime}}))=V(\cdot;D_{t^{\prime}})$ from the symmetry of score function $V$. And equality (ii) is true as $\{V({\bf X}_{i};D_{t^{\prime}}):i\in{\mathcal{C}}_{0t^{\prime}}\}=\{V({\bf X}_{\sigma(i)};D_{t^{\prime}}):i\in{\mathcal{C}}_{0t^{\prime}}\}$.
Therefore,
$\Pr(p_{t}\leq x_{t},\cdots,p_{T}\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{T}})$
$=$ $\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}{\mathbb{E}}\left[\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{p_{t}(({\mathcal{D}}_{T})_{\sigma})\leq x_{t},\cdots,p_{T}(({\mathcal{D}}_{T})_{\sigma})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{T}}\}\right]$
$=$ $\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}{\mathbb{E}}\left[\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{p_{t}({\mathcal{D}}_{T})\leq x_{t},\cdots,p_{T}({\mathcal{D}}_{T})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{T}}\}\right]$
$=$ $\frac{1}{(|{\mathcal{C}}_{0t}|+1)!}{\mathbb{E}}\left[\left(\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{p_{t}({\mathcal{D}}_{T})\leq x_{t}\}\right)\mathbb{I}\{p_{t+1}({\mathcal{D}}_{T})\cdots,p_{T}({\mathcal{D}}_{T})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{T}}\}\right]$
$=$ ${\mathbb{E}}\left[\Pr\left(p_{t}\leq x_{t}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Phi_{t}\cup\{({\bf X}_{i},Y_{i})\}_{i=t+1}^{T}}\right)\mathbb{I}\{p_{t+1}({\mathcal{D}}_{T})\cdots,p_{T}({\mathcal{D}}_{T})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{T}}\}\right]$
$=$ $x_{t}\Pr(p_{t+1}({\mathcal{D}}_{T})\cdots,p_{T}({\mathcal{D}}_{T})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{T}}).$
Repeating the above argument recursively for $s=t,t+1,\ldots,T$, and conditioning at step $s$ on $\Phi_{s}\cup\{({\bf X}_{i},Y_{i})\}_{i=s}^{T}$, we obtain
$\Pr\bigl(p_{t}\leq x_{t},\,p_{t+1}\leq x_{t+1},\dots,p_{T}\leq x_{T}\bigr)=x_{t}\,x_{t+1}\,\cdots\,x_{T}.$
Since the joint cumulative distribution function factors as the product of the marginals, it follows that the sequence $\{p_{t},\dots,p_{T}\}$ is mutually independent, with each $p_{t}$ marginally distributed as $\mathrm{Uniform}(0,1)$.
Thus, we have demonstrated that under the exchangeability assumption and the online updating scheme, the online conformal $p$-values defined in (37) are mutually independent. ∎
F.6 Proof of Theorem 4: finite sample mFDR control
Proof.
mFDR control for Algorithm 2. We first prove the results for LFS. Denote $\Psi_{t}=\Big((p_{k}:k\in{\mathcal{C}}_{0t}),(\theta_{k}:k<t),\big((X_{k},Y_{k}):k\in{\mathcal{C}}_{1t}\big)\Big)$. We need to verify two facts, for any time $t\in[T]$,
(i): $\alpha_{t}^{\text{LFS}}$ is fixed given $\Psi_{t}$. Since $\Psi_{t}$ contains all non-null information available up to time $t$, we can regard the p-values $\{p_{i}\}_{i\leq t}$ as being constructed without using any non-null data; the general case then follows directly by conditioning on $\Psi_{t}$ as non-null data are all fixed. Note that $\alpha_{t}^{\text{LFS}}$ is fixed given all past null decisions $(\delta_{i}:1\leq i<t)$ and a null decision $\delta_{i}$ is decided by $((p_{k},\alpha_{k}):\theta_{k}=0,1\leq k\leq i)$. By iterated discussion, $\alpha_{t}$ is decided by null p-values $(p_{k}:k\in{\mathcal{C}}_{0t})$, thereby determined fully by $\Psi_{t}$.
(ii): $p_{t}$ is super-uniform given $\Psi_{t}$ and $\theta_{t}=0$. This is direct as long as $p_{t}$ is independent of past null p-values given non-null data, which is verified by Proposition 4.1.
Then we have
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t}(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$
${=}$ $\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j}){\mathbb{E}}\left[\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid\Psi_{j},\theta_{j}=0\right]\right]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})+\alpha_{t}\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}(1-\theta_{j})\right]$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right]=\alpha\cdot{\mathbb{E}}[R(t)],$
where the inequality (i) follows from the above two facts, thereby $\Pr(p_{j}\leq\alpha_{j}\mid\Psi_{j},\theta_{j}=0)\leq\alpha_{j}$. And the inequality (ii) follows by the fact that $1-\theta_{t}\leq 1$. For inequality (iii) follows from the construction of $\alpha_{t}^{\text{LFS}}$, and the inequality (iv) holds since $1-\theta_{j}\leq 1$ for all $j\leq t$. Therefore, we conclude that $\operatorname{mFDR}(t)\leq\alpha$ for LFS.
Similar results for SFS can also be proved as follows:
${\mathbb{E}}[V(t)]$ $=$ ${\mathbb{E}}\left[\sum_{j\leq t}(1-\theta_{j})\mathbb{I}\{p_{j}\leq\alpha_{j}\}\right]$
${=}$ $\sum_{j\leq t}{\mathbb{E}}\left[(1-\theta_{j}){\mathbb{E}}\left[\mathbb{I}\{p_{j}\leq\alpha_{j}\}\mid\Psi_{j},\theta_{j}=0\right]\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\alpha_{j}(1-\theta_{j})\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}(1-\theta_{t})\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}.\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j=1}^{t-1}\alpha_{j}(1-\theta_{j})\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}+\alpha_{t}\frac{\mathbb{I}\{p_{t}>\lambda\}}{(1-\lambda)}\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}(1-\theta_{j})\right]$
$\leq$ $\alpha{\mathbb{E}}\left[1\vee\sum_{j\leq t}\delta_{j}\right]=\alpha\cdot{\mathbb{E}}[R(t)].$
Therefore, we obtain the conclusion that $\operatorname{mFDR}(t)\leq\alpha$ for the proposed OCTF procedures in Algorithm 2 with $\alpha_{t}=\alpha_{t}^{\text{LFS}}$ or $\alpha_{t}=\alpha_{t}^{\text{SFS}}$. ∎
F.7 Proof of Corollary 4.1
Proof.
It suffices to verify that the optimized score function $V(\cdot;\hat{k}_{t})$ is symmetric with respect to ${\mathcal{C}}_{0t}\cup\{t\}$, so that Proposition F.1 and Theorem 4 can be directly applied to establish the validity of mFDR control at the target level for Algorithm Opt-OCTF.
For any permutation $\sigma\in\Omega_{t}$ that only permutes the indices in ${\mathcal{C}}_{0t}\cup\{t\}$, we have
$\hat{k}_{t}^{\sigma}=\underset{k\in[K]}{\arg\min}\;{\mathcal{M}}^{\mathrm{EWMA}}_{t}(({\mathcal{D}}_{T})_{\sigma};k)=\hat{k}_{t}.$
To see why, the auxiliary p-value for $j\in{\mathcal{C}}_{1t}$ after permutation $\sigma$ is
$(\tilde{p}_{k,j})_{\sigma}$ $=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{\sigma(s)};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|}$
$=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq V(\mathbf{X}_{j};k)\}}{1+|{\mathcal{C}}_{0t}|}=\tilde{p}_{k,j}.$
This means $\{\tilde{p}_{k,j}\}_{j\in{\mathcal{C}}_{1t}}$ is permutation invariant to $\sigma$. Applying this, we have
${\mathcal{M}}^{\mathrm{EWMA}}_{t}(({\mathcal{D}}_{t})_{\sigma};k)=\frac{\sum_{j=1}^{t-1}\rho^{t-1-j}\,(\tilde{p}_{k,j})_{\sigma}\theta_{j}\,}{\sum_{j=1}^{t-1}\rho^{t-1-j}\cdot\theta_{j}}=\frac{\sum_{j=1}^{t-1}\rho^{t-1-j}\,\tilde{p}_{k,j}\theta_{j}\,}{\sum_{j=1}^{t-1}\rho^{t-1-j}\cdot\theta_{j}}={\mathcal{M}}^{\mathrm{EWMA}}_{t}({\mathcal{D}}_{t};k),$
which keeps invariant to the permutation $\sigma$. Combing together, we have $\hat{k}_{t}$ is symmetric to ${\mathcal{C}}_{0t}\cup\{t\}$. And it indicates that $V(\cdot;\hat{k}_{t})$ is symmetric to ${\mathcal{C}}_{0t}\cup\{t\}$. ∎
F.8 Proof of Theorem 6
Proof.
We first prove the results for $e$-GAIF. Since $\delta_{j}=\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}$, by definition, we have
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}(1-\theta_{j})}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\mathbb{I}\{1/e_{j}\leq\alpha_{j}\}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})e_{j}\alpha_{j}}{R({j-1})+1}\right]$
$=$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j}){\mathbb{E}}[e_{j}\mid{\mathcal{F}}_{j-1}]\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{1+R({j-1})}+\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{1+R({j-1})}\right]$
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\text{e-GAIF}}(t)\right]\leq\alpha,$
where the first inequality holds since $R({j-1})+1\leq(R(t)\vee 1)$ for every $j\in\{j\leq t:\delta_{j}=1\}$ by definition, the second inequality holds since $\mathbb{I}\{e_{j}\alpha_{j}\geq 1\}\leq e_{j}\alpha_{j}$, the third inequality uses the law of iterated expectations by conditioning on $\mathcal{F}_{j-1}$ and then applies the property of $e$-values, and the fourth inequality holds since $1-\theta_{t}\leq 1$, and the last inequality follows from the construction of $e$-GAIF, which completes the proof of FDR control for $e$-GAIF.
We now proceed to establish the FDR control guarantee for the $e$-Ada-GAIF procedure. Specifically, we show that:
$\operatorname{FDR}(t)$ $=$ ${\mathbb{E}}\left[\frac{\sum_{j\leq t}\mathbb{I}\{e_{j}\geq 1/\alpha_{j}\}(1-\theta_{j})}{1\vee\sum_{j\leq t}\delta_{j}}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\mathbb{I}\{1/e_{j}\leq\alpha_{j}\}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})e_{j}\alpha_{j}}{R({j-1})+1}\right]$
$=$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j}){\mathbb{E}}[e_{j}\mid\mathcal{J}_{j-1}]\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\leq t}\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\right]$
$\leq$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\cdot\frac{{\mathbb{E}}\left[\mathbb{I}\{e_{j}<1/\lambda\}\mid\mathcal{J}_{j-1}\right]}{1-\lambda}\right]$
$=$ $\sum_{j\leq t}{\mathbb{E}}\left[\frac{(1-\theta_{j})\alpha_{j}}{R({j-1})+1}\cdot\frac{\mathbb{I}\{e_{j}<1/\lambda\}}{1-\lambda}\right]$
$\leq$ ${\mathbb{E}}\left[\sum_{j\in\mathcal{I}_{t}}\frac{\alpha_{j}(1-\theta_{j})}{R({j-1})+1}\cdot\frac{\mathbb{I}\{e_{j}\leq 1/\lambda\}}{1-\lambda}+\sum_{j\in\bar{\mathcal{I}}_{t}}\frac{\alpha_{j}}{R({j-1})+1}\cdot\frac{\mathbb{I}\{e_{j}\leq 1/\lambda\}}{1-\lambda}\right]$
$=$ ${\mathbb{E}}\left[\widehat{\operatorname{FDP}}_{\text{e-Ada-GAIF}}(t)\right]\leq\alpha,$
where the first inequality holds since $R({j-1})+1\leq(R(t)\vee 1)$ for every $j\in\{j\leq t:\delta_{j}=1\}$ by definition, the second inequality holds since $\mathbb{I}\{e_{j}\alpha_{j}\geq 1\}\leq e_{j}\alpha_{j}$, the third inequality uses the law of iterated expectations by conditioning on $\mathcal{F}_{j-1}$ and then applies the property of $e$-values, and the fourth inequality holds since ${\mathbb{E}}\left[\mathbb{I}\{e_{j}<1/\lambda\}\mid\mathcal{J}_{j-1}\right]\geq 1-\lambda$ by the property of $e$-values, the fifth inequality follows from $1-\theta_{t}\leq 1$, and the last inequality follows from the construction of $e$-Ada-GAIF, which completes the proof of FDR control for $e$-Ada-GAIF.
F.9 Proof of Corollary C.1
Proof.
We prove the result in two steps.
Step 1: mFDR control within each sub-stream.
Fix a sub-stream index $j\in\{1,2,\ldots,d+1\}$ and consider the sequence of testing times $\{t_{k}\}_{k\geq 1}$ such that $t_{k}\in S_{j}$. By construction of the sub-streams, the feedback associated with test $t_{k}$ is revealed at time $t_{k+1}$, which belongs to the same sub-stream. Therefore, when conditioning on the sub-stream index $j$, the testing procedure restricted to $\{t_{k}\}_{k\geq 1}$ has fully observed feedback and coincides with the standard OCTF procedure (without delay).
Under Assumption 1, the conformal $p$-values $\{p_{t_{k}}\}_{k\geq 1}$ computed using the calibration sets $\{{\mathcal{C}}_{0t}^{\,j}\}$ are super-uniform under the null, conditional on the past within the same sub-stream. Moreover, the test levels $\{\alpha_{t_{k}}\}_{k\geq 1}$ are updated based only on information available prior to $t_{k}$ within the sub-stream. Hence, by the finite-sample mFDR guarantee of OCTF, the rejection set restricted to sub-stream $S_{j}$, denoted by ${\mathcal{R}}^{(j)}={\mathcal{R}}\cap S_{j}$, satisfies
$\operatorname{mFDR}^{(j)}(t)\leq\alpha\quad\text{for all }t\in S_{j}.$
Step 2: Aggregation across sub-streams.
Let $V(t)$ and $R(t)$ denote the number of false discoveries and total rejections up to time $t$, respectively. Since the sub-streams $\{S_{j}\}_{j=1}^{d+1}$ form a partition of $\mathbb{N}$, we have
$V(t)=\sum_{j=1}^{d+1}V^{(j)}(t),\qquad R(t)=\sum_{j=1}^{d+1}R^{(j)}(t),$
where $V^{(j)}(t)$ and $R^{(j)}(t)$ are the corresponding quantities restricted to sub-stream $S_{j}$.
By definition of mFDR and linearity of expectation,
$\operatorname{mFDR}(t)=\frac{\mathbb{E}[V(t)]}{\mathbb{E}[R(t)\vee 1]}=\frac{\sum_{j=1}^{d+1}\mathbb{E}[V^{(j)}(t)]}{\mathbb{E}[R(t)\vee 1]}.$
Using the fact that $\mathbb{E}[R(t)\vee 1]\geq\sum_{j=1}^{d+1}\mathbb{E}[R^{(j)}(t)]$ and the sub-stream-wise mFDR control established in Step 1, we obtain
$\operatorname{mFDR}(t)\leq\frac{\sum_{j=1}^{d+1}\mathbb{E}[V^{(j)}(t)]}{\sum_{j=1}^{d+1}\mathbb{E}[R^{(j)}(t)\vee 1]}\leq\alpha\frac{\sum_{j=1}^{d+1}\mathbb{E}[R^{(j)}(t)]}{\sum_{j=1}^{d+1}\mathbb{E}[R^{(j)}(t)\vee 1]}\leq\alpha.$
Therefore, the rejection set $\mathcal{R}$} produced by Algorithm 4 satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}$. ∎
F.10 Proof of Theorem 5
Proof.
For simplicity, we define that when $\theta_{t}=1$,
$\hat{M}_{k,t}:={\mathcal{M}}^{\mathrm{EWMA}}_{t}({\mathcal{D}}_{t};k)=\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,\tilde{p}_{k,j}\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}\quad\text{and}\quad M_{k,t}:={\mathbb{E}}[F^{(0)}_{k}(V(\mathbf{X}_{t};k))\mid\theta_{t}=1].$
And we set historical data ${\mathcal{D}}_{\mathcal{C}}$ as empty set for notational convenience. The final guarantee will remain valid by replacing $t$ with $n+t$ when considering historical samples.
We decompose the error by
$\hat{M}_{k,t}-{M}_{k,t}=$ $\underbrace{\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,(\tilde{p}_{k,j}-F^{(0)}_{k}(V({\bf X}_{j};k)))\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}}_{\text{(I) null CDF estimation}}$
$+\underbrace{\frac{\sum_{j<t}\rho^{t-1-j}\theta_{t}\,(F^{(0)}_{k}(V({\bf X}_{j};k))-{\mathbb{E}}[F^{(0)}_{k}(V({\bf X}_{j};k))\mid\theta_{j}=1])\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}}_{\text{(II) gap of EWMA estimation}}$
$+\underbrace{\frac{\sum_{j<t}\rho^{t-1-j}\theta_{j}\,({\mathbb{E}}[F^{(0)}_{k}(V({\bf X}_{j};k))\mid\theta_{j}=1]-{\mathbb{E}}[F^{(0)}_{k}(V({\bf X}_{t};k))\mid\theta_{t}=1])\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}}_{\text{(III) drift bias}}.$ (38)
Then we check the three term one by one.
(I) Null CDF estimation. Firstly, we condition on the past states $(\theta_{s})_{s<t}$ and $\{\theta_{t}=1\}$ to give results. In this case, ${\mathcal{C}}_{0t}$ and ${\mathcal{C}}_{1t}$ are fixed. Denote
$\hat{F}_{k,t}^{(0)}(v)=\frac{\sum_{s\in{\mathcal{C}}_{0t}\cup\{t\}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq v\}}{1+|{\mathcal{C}}_{0t}|}.$
On the event $\{\theta_{t}=1\}$, the score $V({\bf X}_{t};k)$ is drawn from the non-null distribution and thus is not a null sample for estimating $F_{0}^{k}$. Let
$\hat{F}_{k,t}^{(0),\text{null}}(v):=\frac{1}{|{\mathcal{C}}_{0t}|}\sum_{s\in{\mathcal{C}}_{0t}}\mathbb{I}\{V(\mathbf{X}_{s};k)\leq v\}.$
Then we can decompose $\hat{F}_{k,t}^{(0)}$ as
$\hat{F}_{k,t}^{(0)}(v)=\frac{|{\mathcal{C}}_{0t}|}{|{\mathcal{C}}_{0t}|+1}\,\hat{F}_{k,t}^{(0),\text{null}}(v)+\frac{1}{|{\mathcal{C}}_{0t}|+1}\,\mathbb{I}\{V({\bf X}_{t};k)\leq v\}.$
Therefore, for all $v$,
$\big|\hat{F}_{k,t}^{(0)}(v)-F_{k}^{(0)}(v)\big|$ $\leq\frac{|{\mathcal{C}}_{0t}|}{|{\mathcal{C}}_{0t}|+1}\,\big|\hat{F}_{k,t}^{(0),\text{null}}(v)-F_{k}^{(0)}(v)\big|+\frac{1}{|{\mathcal{C}}_{0t}|+1}\,\big|\mathbb{I}\{V({\bf X}_{t};k)\leq v\}-F_{k}^{(0)}(v)\big|$
$\leq\sup_{u}\big|\hat{F}_{k,t}^{(0),\text{null}}(u)-F_{k}^{(0)}(u)\big|+\frac{1}{|{\mathcal{C}}_{0t}|+1},$
since $\big|\mathbb{I}\{V({\bf X}_{t};k)\leq v\}-F_{k}^{(0)}(v)\big|\leq 1$.
Conditional on $(\theta_{s})_{s<t}$, the collection $\{V({\bf X}_{s};k)\}_{s\in{\mathcal{C}}_{0t}}$ are i.i.d. from $F_{k}^{(0)}$. By the DKW inequality, for any $\varepsilon_{1}>0$,
$\Pr\!\left(\sup_{v}\big|\hat{F}_{k,t}^{(0),\text{null}}(v)-F_{k}^{(0)}(v)\big|\leq\varepsilon_{1}\mid(\theta_{s})_{s<t},\theta_{t}=1\right)\geq 1-2\exp\left\{-2|{\mathcal{C}}_{0t}|\varepsilon_{1}^{2}\right\}.$
Combining the two displays yields
$\Pr\!\left(\sup_{v}\big|\hat{F}_{k,t}^{(0)}(v)-F_{k}^{(0)}(v)\big|\leq\varepsilon_{1}+\frac{1}{|{\mathcal{C}}_{0t}|+1}\mid(\theta_{s})_{s<t},\theta_{t}=1\right)\geq 1-2\exp\left\{-2|{\mathcal{C}}_{0t}|\varepsilon_{1}^{2}\right\}.$
As $\tilde{p}_{k,j}=\hat{F}_{k,t}^{(0)}\!\left(V({\bf X}_{j};k)\right)$, we further have
$\Pr\!\left(\sup_{j\in{\mathcal{C}}_{1t}}\big|\tilde{p}_{k,j}-F_{k}^{(0)}(V(\mathbf{X}_{j};k))\big|\leq\varepsilon_{1}+\frac{1}{|{\mathcal{C}}_{0t}|+1}\mid(\theta_{s})_{s<t},\theta_{t}=1\right)\geq 1-2\exp\left\{-2|{\mathcal{C}}_{0t}|\varepsilon_{1}^{2}\right\}.$
Since $|{\mathcal{C}}_{0t}|=\sum_{s=1}^{t-1}\mathbb{I}\{\theta_{s}=0\}$ and $\Pr(\theta_{s}=0)\geq\underline{\pi}$, by Hoeffding’s inequality,
$\Pr\Big(|{\mathcal{C}}_{0t}|\geq(t-1)\underline{\pi}/2\Big)=\Pr\Big(|{\mathcal{C}}_{0t}|-(t-1)\underline{\pi}\geq-(t-1)\underline{\pi}/2\Big)\geq 1-\exp\left\{-(t-1)\underline{\pi}^{2}/2\right\}.$
Therefore, on the event $\mathbb{P}_{t}=\big\{|{\mathcal{C}}_{0t}|\geq(t-1)\underline{\pi}/2\big\}$, we have
$\Pr\!\left(\sup_{j\in{\mathcal{C}}_{1t}}\big|\tilde{p}_{k,j}-F_{k}^{(0)}(V(\mathbf{X}_{j};k))\big|\leq\varepsilon_{1}+\frac{2}{(t-1)\underline{\pi}}\mid\mathbb{P}_{t},\theta_{t}=1\right)\geq 1-2\exp\left\{-\{(t-1)\underline{\pi}\}\varepsilon_{1}^{2}\right\}.$
Then by the independence of $\mathbb{P}_{t}$ and $\{\theta_{t}=1\}$, invoking Lemma 3 leads to
$\Pr\left(\left|\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,(\tilde{p}_{k,j}-F_{k}^{(0)}(V({\bf X}_{j};k)))\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}\right|\leq\varepsilon_{1}+\frac{2}{(t-1)\underline{\pi}}\mid\theta_{t}=1\right)$
$\geq$ $1-2\exp\left\{-\{(t-1)\underline{\pi}\varepsilon_{1}^{2}\right\}-\exp\left\{-(t-1)\underline{\pi}^{2}/2\right\}.$ (39)
(II) Gap of EWMA estimation. Define $\Omega_{t}=\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}=\sum_{j=1}^{t-1}\rho^{t-1-j}\theta_{j}$. It follows that ${\mathbb{E}}[\Omega_{t}]=\sum_{j=1}^{t-1}\rho^{t-1-j}{\mathbb{E}}[\theta_{j}]\geq(1-\overline{\pi})(1-\rho^{t-1})/(1-\rho)$. By Hoeffding inequality, we have
$\Pr\left(\Omega_{t}\geq\frac{1-\overline{\pi}}{2}\frac{1-\rho^{t-1}}{1-\rho}\right)\geq 1-\exp\left\{-\frac{(1-\overline{\pi})^{2}(1-\rho^{2})}{2(1-\rho^{2(t-1)})}\right\}.$ (40)
Then we can apply the Hoeffding inequality to the weighted sum as
$\left\{\theta_{j}\big(F_{k}^{(0)}(V({\bf X}_{j};k))-{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{j};k))\mid\theta_{j}=1]\big)\right\}_{j<t}$
are bounded between $[0,1]$ and independent:
$\Pr\left(\left|\sum_{j<t}\rho^{t-1-j}\theta_{t}\,(F^{(0)}_{k}(V({\bf X}_{j};k))-{\mathbb{E}}[F^{(0)}_{k}(V({\bf X}_{j};k))\mid\theta_{j}=1])\right|\leq\varepsilon_{2}\Omega_{t}\right)$
$\geq$ $\Pr\left(\left|\sum_{j<t}\rho^{t-1-j}\theta_{t}\,(F^{(0)}_{k}(V({\bf X}_{j};k))-{\mathbb{E}}[F^{(0)}_{k}(V({\bf X}_{j};k))\mid\theta_{j}=1])\right|\leq\varepsilon_{2}\Omega_{t},\Omega_{t}\geq\frac{1-\overline{\pi}}{2}\frac{1-\rho^{t-1}}{1-\rho}\right)$
$\geq$ $1-\Pr\left(\left|{\sum_{j<t}\rho^{t-1-j}\theta_{j}\,(F_{k}^{(0)}(V({\bf X}_{j};k))-{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{j};k))\mid\theta_{j}=1])\,}\right|\geq\varepsilon_{2}\left\{\frac{1-\overline{\pi}}{2}\frac{1-\rho^{t-1}}{1-\rho}\right\}\right)$
$-\Pr\left(\Omega_{t}\leq\frac{1-\overline{\pi}}{2}\frac{1-\rho^{t-1}}{1-\rho}\right)$
$\geq$ $1-2\exp\left\{-\frac{2\varepsilon_{2}^{2}(1-\rho^{2})\left\{\frac{1-\overline{\pi}}{2}\frac{1-\rho^{t-1}}{1-\rho}\right\}^{2}}{1-\rho^{2(t-1)}}\right\}-\exp\left\{-\frac{(1-\overline{\pi})^{2}(1-\rho^{2})}{2(1-\rho^{2(t-1)})}\right\}.$ (41)
(III) Drift bias. By Stieltjes integrals,
$\left|{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{j};k))\mid\theta_{j}=1]-{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1]\right|$ $=|\int F_{k}^{(0)}(v)dF_{k,j}^{(1)}(v)-\int F_{k}^{(0)}(v)dF_{k,t}^{(1)}(v)|$
$=|\int(F_{k,j}^{(1)}(v)-F_{k,t}^{(1)}(v))dF_{k}^{(0)}(v)|$
$\leq\|F_{k,j}^{(1)}-F_{k,t}^{(1)}\|_{\infty}.$
So relating it with our slowly varying distribution condition leads to
$\left|\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,({\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{j};k))\mid\theta_{j}=1]-{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1])\,}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}\right|$
$\leq\gamma\frac{\sum_{j=1}^{t-1}\rho^{t-1-j}(t-j)}{\Omega_{t}}\leq\gamma\frac{1}{\Omega_{t}(1-\rho)^{2}}.$
The last inequality comes from the fact that $\sum_{j=1}^{t-1}(t-j)\rho^{t-1-j}=\{1-t\rho^{t-1}+(t-1)\rho^{t}\}/(1-\rho)^{2}\leq 1/(1-\rho)^{2}$.
Since we have already analyzed the high-probability lower bound of $\Omega_{t}$ in (40), it follows that
$\Pr\left(|\bar{\Delta}_{t}|\leq 2\gamma\{(1-\overline{\pi})(1-\rho^{t-1})(1-\rho)\}^{-1}\right)\geq 1-\exp\left\{-\frac{(1-\overline{\pi})^{2}(1-\rho^{2})}{2(1-\rho^{2(t-1)})}\right\},$ (42)
where
$\bar{\Delta}_{t}:=\frac{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}\,\Big({\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{j};k))\mid\theta_{j}=1]-{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1]\Big)}{\sum_{j\in{\mathcal{C}}_{1t}}\rho^{t-1-j}}.$
Combining the results in (F.10), (F.10) and (42) together into (F.10), we have for any $\varepsilon_{1}\in(0,\sqrt{\underline{\pi}/2})$ and $\varepsilon_{2}\in(0,1-\rho)$
$|\hat{M}_{k,t}-{M}_{k,t}|\leq\varepsilon_{1}+\varepsilon_{2}+2\gamma\{(1-\overline{\pi})(1-\rho^{t-1})(1-\rho)\}^{-1}+2\{(t-1)\underline{\pi}\}^{-1}$
with probability
$1-3\exp\left\{-\{(t-1)\underline{\pi}\varepsilon_{1}^{2}\right\}-4\exp\left\{-\frac{\varepsilon_{2}^{2}(1+\rho)({1-\overline{\pi}})^{2}({1-\rho^{t-1}})}{2(1+\rho^{t-1})(1-\rho)}\right\}.$
Suppose $t>1+1/(1-\rho)$ such that $\rho^{t-1}<1/2$, then
$\frac{(1+\rho)({1-\rho^{t-1}})}{2(1+\rho^{t-1})(1-\rho)}\geq\frac{1}{3}\frac{1+\rho}{1-\rho}.$
Then the part (i) is verified.
For part (ii),
$\max_{k\in[K]}|\hat{M}_{t}^{k}-{M}_{t}^{k}|\leq\varepsilon_{1}+\varepsilon_{2}+4\gamma\{(1-\overline{\pi})(1-\rho)\}^{-1}+2\{(t-1)\underline{\pi}\}^{-1}$
holds with probability
$1-3K\exp\left\{-\{(t-1)\underline{\pi}\varepsilon_{1}^{2}\right\}-4K\exp\left\{-\frac{\varepsilon_{2}^{2}({1-\overline{\pi}})^{2}}{6(1-\rho)}\right\}.$
By Assumption 5,
$\Pr(\hat{k}_{t}\neq k^{*}_{t}\mid\theta_{t}=1)\leq\Pr(\max_{k\in[K]}|\hat{M}_{k,t}-{M}_{k,t}|\geq c/2).$
To enforce $\max_{k}|\hat{M}_{k,t}-M_{k,t}|\leq c/2$, it suffices to let
$\varepsilon_{1}=\varepsilon_{2}\leq\frac{c}{8},\;\;\frac{4\gamma}{(1-\overline{\pi})(1-\rho)}\leq\frac{c}{8},\;\;\frac{2}{(t-1)\underline{\pi}}\leq\frac{c}{8}.$
So it is required that
$t\geq 1+\frac{16}{c\underline{\pi}},\quad\frac{\gamma}{1-\rho}\leq\frac{(1-\overline{\pi})c}{32}.$
And we have the explicit bound
$\Pr(\hat{k}_{t}\neq k_{t}^{*})\leq\;$ $3K\exp\!\left\{-(t-1)\underline{\pi}\frac{c^{2}}{64}\right\}+4K\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}.$
And this bound can be directly generalized when the historical data are not empty by replacing $t$ with $n+t$.
Finally, we can jointly guarantee the score selection consistency for all non-null points:
$\Pr(\exists i\leq t:\theta_{i}=1,\ \hat{k}_{i}\neq k_{i}^{*})\leq\sum_{i=1}^{t}\Pr(\theta_{i}=1)\Pr(\hat{k}_{i}\neq k_{i}^{*}\mid\theta_{i}=1)$
$\leq$ $3K(1-\underline{\pi})\sum_{i=1}^{t}\exp\!\left\{-(n+i-1)\underline{\pi}\frac{c^{2}}{64}\right\}+4K(1-\underline{\pi})t\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}$
$\leq$ $\frac{3K(1-\underline{\pi})\exp\!\left\{-n\underline{\pi}{c^{2}}/{64}\right\}}{1-\exp\!\left\{-\underline{\pi}{c^{2}}/{64}\right\}}+4K(1-\underline{\pi})t\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}.$
F.11 Proof of Theorem 7
Proof.
It suffices to prove Proposition F.1 for the permuted $p$-value. Naturally, we can view score selection as part of the score construction such that $V(\cdot;\hat{k}_{t})$ can be defined as $V(\cdot;{\mathcal{D}}_{t})$. Here we do not require $V(\cdot;{\mathcal{D}}_{t})$ is symmetric to past null data and allow it perform non-symmetrically. For convenience, we set $\theta_{t}=\theta_{t+1}=\cdots=\theta_{T}=0$.
For any $\sigma^{\prime}\in\Omega_{t}$ and given $\{{\bf X}_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}=\{{\bf x}_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ as the realizations, define $\bar{\sigma}=\sigma^{\prime}\cdot\sigma$ and
$p_{t}^{\rm orc}(({\mathcal{D}}_{T})_{\sigma^{\prime}})$
$=$ $\frac{1}{|\Omega_{t}|}\sum_{\sigma\in\Omega_{t}}{\mathbb{I}}\{V({\bf x}_{{\sigma^{\prime}(\sigma(i))}};(({\mathcal{D}}_{t})_{\sigma})_{\sigma^{\prime}})<V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})\}$
$+\frac{1}{|\Omega_{t}|}\xi_{t}\cdot\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{V({\bf x}_{\sigma^{\prime}(\sigma(i))};(({\mathcal{D}}_{t})_{\sigma})_{\sigma^{\prime}}))=V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})\}$
$=$ $\frac{\sum_{\bar{\sigma}\in\Omega_{t}}{\mathbb{I}}\{V({\bf x}_{{\bar{\sigma}(i)}};({\mathcal{D}}_{t})_{\bar{\sigma}})<V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})\}+\xi_{t}\cdot\sum_{\bar{\sigma}\in\Omega_{t}}\mathbb{I}\{V({\bf x}_{\bar{\sigma}(i)};({\mathcal{D}}_{t})_{\bar{\sigma}}))=V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})\}}{|\Omega_{t}|}$
$=$ $\frac{\sum_{{\sigma}\in\Omega_{t}}{\mathbb{I}}\{V({\bf x}_{{{\sigma}(i)}};({\mathcal{D}}_{t})_{{\sigma}})<V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})\}+\xi_{t}\cdot\sum_{\sigma\in\Omega_{t}}\mathbb{I}\{V({\bf x}_{{\sigma}(i)};({\mathcal{D}}_{t})_{{\sigma}}))=V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})\}}{|\Omega_{t}|}.$
The last equality comes from the fact that $\sigma^{\prime}\in\Omega_{t}$ is a bijection such that $\{\sigma:\sigma\in\Omega_{t}\}=\{\sigma^{\prime}\cdot\sigma:\sigma\in\Omega_{t}\}$.
Follow the same discussion in the proof of Proposition F.1. Let $\{X_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}=\{x_{i}:i\in{\mathcal{C}}_{0t}\cup\{t\}\}$ as a set of realizations and
$q=Q_{\alpha}(V(x_{\sigma(t)};({\mathcal{D}}_{t})_{\sigma}):\sigma\in\Omega_{t}).$
For any $\sigma^{\prime}\in\Omega_{t}$ such that $V({\bf x}_{\sigma^{\prime}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})>q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma^{\prime}})\geq\frac{N_{=}+N_{-}}{|\Omega_{t}|}\geq\alpha;$
Here
$N_{-}=\sum_{{\sigma^{\prime}}\in\Omega_{t}}\mathbb{I}\{(V(x_{{\sigma^{\prime}}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})<q\},\quad N_{=}=\sum_{{\sigma^{\prime}}\in\Omega_{t}}\mathbb{I}\{V(x_{{\sigma^{\prime}}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})=q\}.$
For any ${\sigma^{\prime}}\in\Omega_{t}$ such that $V({\bf x}_{{\sigma^{\prime}}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})<q$, we have
$p_{t}(({\mathcal{D}}_{T})_{\sigma^{\prime}})\leq\frac{N_{-}}{|\Omega_{t}|}\leq\alpha.$
Hence
$\Pr(p_{t}\leq\alpha\mid\Phi_{t})$ $=\frac{1}{|\Omega_{t}|}\sum_{{\sigma^{\prime}}\in\Omega_{t}}\Pr(p^{\rm opt}_{t}(({\mathcal{D}}_{T})_{\sigma^{\prime}})\leq\alpha)$
${=}\frac{1}{|\Omega_{t}|}\sum_{{\sigma^{\prime}}\in\Omega_{t}}\left(\mathbb{I}\{V({\bf x}_{{\sigma^{\prime}}(t)};({\mathcal{D}}_{t})_{\sigma^{\prime}})<q\}+\mathbb{I}\{V({\bf x}_{{\sigma^{\prime}}(t)};{\mathcal{D}}_{t})=q\}\frac{\alpha|\Omega_{t}|-N_{-}}{N_{=}}\right)$
${=}\frac{1}{|\Omega_{t}|}\left(N_{-}+N_{=}\frac{\alpha|\Omega_{t}|-N_{-}}{N_{=}}\right)=\alpha,$
As for the independence of null $p$-values, by a similar discussion in Proposition F.1, we can still have
$p_{t^{\prime}}({\mathcal{D}}_{T})=p_{t^{\prime}}(({\mathcal{D}}_{T})_{\sigma^{\prime}})$
for all $t^{\prime}\in\{t+1,\dots,T\}$ and any $\sigma^{\prime}\in\Omega_{t}$. Then by the same iterate conditional expectation without relying on additional symmetry discussion, we can provide the independence results.